### v1

# Lagrangian extensions and left symmetric structures on the four-dimensional real Lie superalgebras

Abstract.
Over real numbers, Backhouse classified all four-dimensional Lie superalgebras. From this list, we will investigate those Lie superalgebras that can be obtained as Lagrangian extensions. Moreover, we investigate left-symmetric structures on these Lie superalgebras. Furthermore, except for two of them, they are all Novikov superalgebras.
Key words and phrases:
Quasi-Frobenius Lie (super)algebra, Lagrangian extensions, Left symmetric structures, Novikov structures, Balinsky-Novikov structures.
2020 Mathematics Subject Classification:
17B05; 17B55; 17D25
SB is supported by the grant NYUAD-065.
1. Introduction
1.1. Lagrangian extensions of Lie (super)algebras
The notion of $T^{*}$-extension of Lie algebras was initially introduced by Bordemann [Bor]. A short explanation can be understood as the following. Given a Lie algebra $\mathfrak{h}$ equipped with an $\mathfrak{h}^{*}$-valued bilinear map $\alpha$, a $T^{*}$-extension of $\mathfrak{h}$ is the Lie algebra structure on the vector space $\mathfrak{g}:=\mathfrak{h}\oplus\mathfrak{h}^{*}$ constructed by means of the form $\alpha$ that satisfies a certain cohomological requirement. The Lie algebra $\mathfrak{g}$ naturally admits an invariant non-degenerate symmetric bilinear form. Conversely, Bordemann showed in [Bor] that every finite-dimensional nilpotent Lie algebra $\mathfrak{h}$ with an invariant symmetric non-degenerate bilinear form “is” a suitable $T^{*}$-extension. In particular, if $\mathfrak{h}$ is a real finite-dimensional Lie algebra belonging to a Lie group $G$, then the $T^{*}$-extension of $\mathfrak{h}$ is the Lie algebra of the cotangent bundle $T^{*}G$ of $G$, considered as a Lie group. The terminology and notation are justified by this differential geometric fact.
A superization of this construction was given in [BBB]. In particular, it was shown that every solvable Lie superalgebra with an even symmetric non-degenerate bilinear form is either isomorphic to a $T^{*}$-extension of a certain Lie superalgebra or to an ideal of codimension one of a certain $T^{*}$-extension.
The notion of $T^{*}$-extension was also studied by Baues and Cortés in [BC] in the context of Lie algebras admitting a flat connection, and called Lagrangian extensions since the Lie sub-algebra $\mathfrak{h}^{*}$ is a Lagrangian ideal in $\mathfrak{h}\oplus\mathfrak{h}^{*}$. Unlike the case studied in [Bor], the non-degenerate form in this context is anti-symmetric and defines a 2-cocycle on $\mathfrak{h}$. In addition, the Lie algebra $\mathfrak{h}$ must admit a flat connection so that the Lagrangian extension is possible. Lie algebras with a non-degenerate anti-symmetric form satisfying the 2-cocycle conditions are called quasi-Frobenius, and are studied intensively in the literature, see, e.g., [E, O].
Lie algebras possessing both an invariant non-degenerate symmetric bilinear form and a non-degenerate 2-cocycle of its scalar cohomology, were studied in [BBM] and were called quadratic symplectic Lie algebras. If the ground field is algebraically closed, the authors showed that every quadratic symplectic Lie algebra is a $T^{*}$-extension of a Lie algebra which admits an invertible derivation. Necessary and sufficient conditions on $\mathfrak{h}$ and on the cocycle used in the construction of the $T^{*}$-extension are investigated to ensure that the extended algebra admits a skew-symmetric derivation and, hence, a symplectic structure. As a result, complex symplectic quadratic Lie algebras with dimensions less than or equal to 8 are classified. Some of these Lie algebras correspond to some Lie groups that admit a bi-invariant pseudo-Riemannian metric as well as a left-invariant symplectic form. These Lie groups are nilpotent, and their geometry is very rich, see [BBM, MR] for more details.
The notion of Lagrangian extension was superized in [BM]. The construction is valid for arbitrary Lie superalgebras defined over a field of characteristic not two provided the Lie superalgebra admit a torsion-free flat connection. Moreover, they have observed that there are two ways to perform this extension: they consider either $\mathfrak{g}:=\mathfrak{h}\oplus\mathfrak{h}^{*}$ or $\mathfrak{g}:=\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*})$, where $\Pi$ is the change of parity functor. They call them $T^{*}$-extension and $\Pi T^{*}$-extensions of $\mathfrak{h}$, respectively. In the first case, the non-degenerate form is even, and in the second case, it is odd.
Several examples of Lagrangian extensions are given in [BM], including filiform Lie superalgebras and some Lie superalgebras over reals. As a source of examples, they used the list of 4-dimensional real Lie superalgebras classified by Backhouse in [B].
One of the main objectives of this paper is to complete the list of examples among the 4-dimensional Lie superalgebras and explain how they can be obtained as Lagrangian extensions, see Section 4. According to the Backhouse classification, this investigation will classify all four-dimensional real Lie superalgebras that admit (or don’t admit) Lagrangian extensions.
1.2. Left-symmetric structures
Left-symmetric algebras (also known as Vinberg-Koszul algebras) play an important role in various fields of mathematics and mathematical physics; for instance, rooted tree algebras, operad theory, vertex algebras, convex homogeneous cones, affine manifolds, left-invariant affine structures on Lie groups, to name just a few. For a thorough review, see [B3].
Over the real and complex numbers left-symmetric algebras are closely related to the theory of left-invariant affine structures over Lie groups, see for example [M, GS] (and references therein).
The investigation and classification of left-symmetric structure on Lie (super)algebras is an important subject by its own, as initiated in [S]. The problem, however, seems to be difficult to handle.
For reductive Lie algebras, Baues [Ba] proved that $\mathfrak{gl}(n)$ is the only reductive Lie algebra with one-dimensional center and a simple semisimple ideal which admits left-symmetric algebras over an algebraically closed ground field (see also [B1]). Morover, all left-symmetric algebras on $\mathfrak{gl}(n)$ were classified.
Simple Lie algebras over complex numbers do not admit left-symmetric structures, see [H]. The existence of such structures is not always guaranteed, even for nilpotent Lie algebras. There are filiform nilpotent Lie groups of dimension $10\leq n\leq 13$ which do not admit any left-invariant affine structure, see [B, B3, BG]. Any filiform nilpotent Lie group of dimension $n\leq 9$ admits a left-invariant affine structure. Nevertheless, left-symmetric algebra structures on both finite-dimensional and infinite-dimensional Lie algebras have been studied extensively, see [Bai, B2, YZ] as well as [CL, KCB, TB], respectively. Left symmetric structures on the Virasoro superalgebra were studied in [KB].
In the super setting, it was proved in [DZ] that $\mathfrak{sl}(m+1|m)$ admits a left-symmetric superalgebra, and a full classification is offered in the case of $\mathfrak{sl}(2|1)$. Moreover, it was shown that $\mathfrak{sl}(m|1)$ does not admit left-symmetric superalgebras for $m\geq 3$.
There is also the notion of Novikov algebra. It was introduced in connection with the Poisson brackets of hydrodynamic type, see [BN].
A second aim of this paper is to provide explicitly a left-symmetric structure on the 4-dimensional real Lie superalgebras. Surprisingly enough, all these Lie superalgebras admit such a structure, given explicitly in Section 5. Furthermore, we show that these structures are Novikov, except for the two Lie superalgebras $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$. In addition, all these Lie superalgebras admit a Balinsky Novikov structure. Classifying all left-symmetric structures on each of these Lie superalgebras remains an open problem.
2. Main concepts and definitions
The purpose of this section is to review the main definitions and theorems that will be used in this paper.
Throughout the text $\mathbb{K}$ is a field of characteristic $p\not=2$.
A Lie superalgebra is a $\mathbf{Z}_{2}$-graded space $\mathfrak{g}=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{\bar{1}}$ over a field of characteristic $p\not=2$ together with a bilinear binary operator $[\cdot,\cdot]:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$, called the Lie superbracket that has the following properties: (for any homogeneous $x,y,z\in\mathfrak{g}$)
  1. (i)
    Super anti-commutativity: $[x,y]=-(-1)^{|x||y|}[y,x]$;
  2. (ii)
    Super Jacobi identity: $[x,[y,z]]=[[x,y],z]+(-1)^{|x||y|}[y,[x,z]]$ where $|x|={\bar{0}}$ if $x\in\mathfrak{g}_{\bar{0}}$ and $|x|={\bar{1}}$ if $x\in\mathfrak{g}_{\bar{1}}$;
  3. (iii)
    If $p=3$, one must add $[x,[x,x]]=0$ if $x\in\mathfrak{g}_{\bar{1}}$.
There is also the notion of Lie superalgebra in characteristic $2$ but we are not studying them here.
Let us introduce the notion of bilinear forms over a superspace. For more details, see [LSoS].
Let $V$ and $W$ be two superspaces defined over an arbitrary field ${\mathbb{K}}$. Let $\omega\in\text{Bil}(V,W)$ be a homogeneous bilinear form. The Gram matrix $B=(B_{ij})$ associated to $\omega$ is given by the formula,
(1) $B_{ij}=(-1)^{|\omega||v_{i}|}\omega(v_{i},v_{j})\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for the basis vectors $v_{i}\in V$.}$
This definition can be extended by linearity to non-homogeneous forms. Moreover, it allows us to identify a bilinear form $B(V,W)$ with an element of $\mathrm{Hom}(V,W^{*})$. Consider the upsetting of bilinear forms $u\colon\mathrm{Bil}(V,W)\rightarrow\mathrm{Bil}(W,V)$ given by the formula
(2) $u(\omega)(u,v)=(-1)^{|v||u|}\omega(v,u)\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for any homogeneous $v\in V$ and $u\in W$.}$
In terms of the Gram matrix $B$ of $\omega$: the form $\omega$ is symmetric if and only if
(3) $u(B)=B,\;\text{ where $u(B)=\left(\begin{array}[]{cc}R^{t}&(-1)^{|\omega|}T^{t}\\[5.69054pt] (-1)^{|\omega|}S^{t}&-U^{t}\end{array}\right),$ for $B=\left(\begin{array}[]{cc}R&S\\[5.69054pt] T&U\end{array}\right)$.}$
Similarly, anti-symmetry of $\omega$ means that $u(B)=-B$.
An even (resp. odd) non-degenerate bilinear form is called orthosymplectic (resp. periplectic).
Proposition 2.1 (Superdimension constraints).
If $V=V_{\bar{0}}\oplus V_{\bar{1}}$ is a finite-dimensional superspace such that $V_{\bar{1}}\not={0}$, equipped with an anti-symmetric non-degenerate bilinear form $\omega$, then:
  1. (i)
    If $p(\omega)=\bar{0}$, then $\mathrm{dim}(V_{\bar{0}})$ is even.
  2. (ii)
    If $p(\omega)=\bar{1}$, then $\mathrm{dim}(V_{\bar{0}})=\mathrm{dim}(V_{\bar{1}})$.
Proof.
See, [BM]
We will denote by $\Pi$ the functor that assigns to $V$ the superspace $\Pi(V)$ defined as the other copy of $V$ with the opposite parity of its homogeneous components. We have:
$(\Pi(V))_{\bar{0}}=V_{\bar{1}},\quad(\Pi(V))_{\bar{1}}=V_{\bar{0}}.$
The space $\Pi(V)$ consists of the linear combinations of elements that we denote by $\Pi(f)$ for every homogeneous $f\in V$. For every superspace $V$, denote the canonical odd homomorphism induced by the functor $\Pi$ by
$\Pi:V\rightarrow\Pi(V)\quad f\mapsto\Pi(f).$
We will identify $V$ and $\Pi(\Pi(V))$, naturally.
Let $\omega\in\mathrm{Bil}(\mathfrak{g},\mathfrak{g})$ be an anti-symmetric bilinear form on a Lie superalgebra $\mathfrak{g}$. The form $\omega$ is called closed if it satisfies:
$(-1)^{|x||z|}\omega(x,[y,z])+(-1)^{|z||y|}\omega(z,[x,y])+(-1)^{|y||x|}\omega(y,[z,x])=0,$
for all $x,y,z\in\mathfrak{g}_{\bar{i}}$. This means that $\omega$ is 2-cocyle with scalar values.
A Lie superalgebra $\mathfrak{g}$ is called quasi-Frobenius if it is equipped with a closed anti-symmetric non-degenerate bilinear form $\omega$. If $\omega$ is even (resp. odd) we call it orthosymplectic (resp. periplectic). We denote such a Lie superalgebra by $(\mathfrak{g},\omega)$. Recently, these Lie superalgebras were studied in [BE, BEM, BM] in the context of Lagrangian and double extensions.
A Lie superalgebra is called Frobenius if $\omega$ is exact, which means that there is $f\in\mathfrak{g}^{*}$ such that $\omega(x,y)=f([x,y])$ for all $x,y\in\mathfrak{g}.$
Let $(\mathfrak{g},\omega)$ be a quasi-Frobenius Lie superalgebra. An ideal $I\subseteq\mathfrak{g}$ is called Lagrangian if and only if $I^{\perp}=I$.
2.1. Connections on Lie (super)algebras
To write down a Lagrangian extension of a Lie (super)algebra, we will need the notion of a connection on a Lie (super)algebra. In our context, connections are defined as non-associative products on a superalgebra with purely algebraic definitions. See [BM, MKSV] for the non-super case.
An even bilinear map $\nabla:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$, written as $(x,y)\rightarrow\nabla_{x}y$ is called a connection on $\mathfrak{g}.$
We will introduce the notion of torsion and curvature of a connection.
Let $\nabla:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ be a connection on $\mathfrak{g}$. For all $x,y,z\in\mathfrak{g}_{\bar{i}}$, the torsion $T^{\nabla}$ of $\nabla$ is defined as
(4) $T(x,y):=\nabla_{x}y-(-1)^{|x||y|}\nabla_{y}x-[x,y].$
The curvature $R^{\nabla}$ of $\nabla$ is defined as
(5) $R^{\nabla}(x,y)h:=\nabla_{x}\nabla_{y}z-(-1)^{|x||y|)}\nabla_{y}\nabla_{x}z-\nabla_{[x,y]}z.$
By bilinearity, the definition of $T$ and $R^{\nabla}$ can be extended to the whole $\mathfrak{g}$.
If $T=0$, then the connection $\nabla$ is called torsion-free. The curvature $R^{\nabla}$ vanishes if and only if the map $\sigma^{\nabla}:\mathfrak{g}\rightarrow\mathrm{End}(\mathfrak{g}),\;x\mapsto\nabla_{x}$ is a representation of $\mathfrak{g}$ on itself. In this case, the connection $\nabla$ will be called flat.
2.2. Left-Symmetric Lie (super)algebras
Let us review a few notions regarding left-symmetric structures.
A superalgebra $L=L_{\bar{0}}\oplus L_{\bar{1}}$ over a field $\mathbb{K}$ equipped with an even bilinear operation $\cdot$ is called a left-symmetric superalgebra (or an LSSA for short) if the associator
$(x,y,z):=(x\cdot y)\cdot z-x\cdot(y\cdot z)$
is supersymmetric in $x$ and $y$, i.e., $(x,y,z)=(-1)^{|x||y|}(y,x,z)$; or, equivalently,
$(x\cdot y)\cdot z-x\cdot(y\cdot z)=(-1)^{|x||y|}((y\cdot x)\cdot z-y\cdot(x\cdot z)),\text{ for all }x,y,z\in L_{\bar{i}}.$
These superalgebras are Lie-admissible. The supercommutator $[x,y]:=x\cdot y-(-1)^{|x||y|}y\cdot x$ defines then a Lie superalgebra structure on $L$ and we denote it by ${\mathfrak{g}}(L)$. The resulting Lie superalgebra is called the associated Lie superalgebra of the LSSA. Note that if $L$ admits an LSSA then the even part $L_{\bar{0}}$ also admits one. Associative superalgebras are all LSSAs.
In [H], Helmstetter showed that it is necessary to have $[L,L]\subsetneq L$ if $L$ is a left symmetric algebra defined over the field of complex numbers. Over a field of characteristic $p$, this is not true anymore. Burde showed in [B2] that classical Lie algebras over a field $p>3$ admit an LSS only in case $p$ divides $\mathrm{dim}(L)$. In the super setting and over a field of prime characteristic, there has not been much research done so far.
The existence of a flat connection $\nabla$ on a Lie superalgebra $\mathfrak{g}$ will induce a left-symmetric structure given by
$x\cdot y:=\nabla_{x}(y)\text{ for all }x,y\in{\mathfrak{g}}.$
Additionally, every quasi-Frobenius Lie superalgebra $({\mathfrak{g}},\omega)$ admits a left symmetric structure given by
$\omega(x\cdot y,z)=(-1)^{|x||y|}\omega(y,[x,z])\text{ for all }x,y\in{\mathfrak{g}}_{\bar{i}},\text{ and }z\in{\mathfrak{g}}.$
The converse is not always true. There will be several examples that illustrate that in Section 5.
Some subclasses of LSSAs are important as well. An LSSA with supercommutative right multiplication is called a Novikov superalgebra (cf. [BN, Xu]). Explicitly, a Novikov superalgebra $(L,\cdot)$ is an LSSA with an additional condition:
$(z\cdot x)\cdot y=(-1)^{|x||y|}(z\cdot y)\cdot x,\;\text{ for all }x,y\in L_{\bar{i}},\text{ and }z\in L.$
It turns out that there are two superizations of the notion of Novikov algebras. The second superization is due to [Bal] (see also [PB]) and will be referred to as Balinsky Novikov superalgebras, or BN superalgebras for short. They are defined as follows:
  1. (1)
    Left symmetry and commutativity:
    $(x\cdot y)\cdot z-x\cdot(y\cdot z)=(y\cdot x)\cdot z-y\cdot(x\cdot z),\qquad(x\cdot y)\cdot z=(x\cdot z)\cdot y,$
    for any homogeneous elements
    $x,y,z\in L_{\bar{0}};\quad x,y\in L_{\bar{0}},z\in L_{\bar{1}};\quad x,z\in L_{\bar{0}},y\in L_{\bar{1}};\quad y,z\in L_{\bar{0}},x\in L_{\bar{1}};$
  2. (2)
    Commutativity: $x\cdot y=y\cdot x$ for all $x,y\in L_{\bar{1}}$;
  3. (3)
    Compatibility conditions:
    $\begin{array}[]{ll}(x\cdot y)\cdot z=(x\cdot z)\cdot y&\text{ for all }x,y\in L_{\bar{1}}\text{ and }z\in L_{\bar{0}};\\[5.69054pt] x\cdot(y\cdot z)=(x\cdot y)\cdot z+y\cdot(x\cdot z)&\text{ for all }x\in L_{\bar{0}}\text{ and }y,z\in L_{\bar{1}};\\[5.69054pt] x\cdot(y\cdot z)=(x\cdot y)\cdot z+(x\cdot z)\cdot y&\text{ for all }x,y,z\in L_{\bar{1}}.\\[5.69054pt] \end{array}$
Balinsky Novikov superalgebras are also admissible Lie superalgebras. The Lie structure is given as follows:
$\begin{array}[]{ll}[x,y]:=x\cdot y-y\cdot x,&\text{ for all }x,y\in L_{\bar{0}};\\[5.69054pt] [x,y]:=x\cdot y-\frac{1}{2}y\cdot x=-[y,x],&\text{ for all }x\in L_{\bar{0}},\text{ and }y\in L_{\bar{1}}\\[5.69054pt] [x,y]:=x\cdot y&\text{ for all }x,y\in L_{\bar{1}}.\end{array}$
2.3. The 4-dimensional real Lie superalgebras
The ground field $\mathbb{K}$ is assumed to be the field of real numbers $\mathbb{R}$. The classification of four-dimensional real Lie superalgebras has been carried out by Backhouse, see [B]. The Lie superalgebras with anti-symmetric bilinear forms were examined in [BM].
As in the [BM] paper, the tables in the Appendix represent the classification of 4-dimensional real Lie superalgebras as follows:
  1. (1)
    Table 1 having the trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=2|2$;
  2. (2)
    Table 2 having the non-trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=2|2$;
  3. (3)
    Table 3 having the trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=3|1$;
  4. (4)
    Table 4 having the trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=1|3$;
  5. (5)
    Table 5 having the non-trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=3|1$;
  6. (6)
    Table 6 having the non-trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=1|3$.
In this paper, we present only the forms as in Table 1-6 which are suitable for Lagrangian extensions, as opposed to [BM], where the forms are given in the form of parametric families. The classification of these structures up to a symplectomorphism is still an open problem. We also correct a claim in [BM] stating that only non-homogenous forms exist for the Lie superalgebras $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$. It turns out that these two superalgebras both admit an even and an odd non-degenerate closed form, and both of them can be obtained as Lagrangian extensions.
3. The notion of $T^{*}$-extensions and $\Pi T^{*}$- (Lagrangian extensions)
We will introduce the notion of $T^{*}$-extension and $\Pi T^{*}$-extension of Lie superalgebras (Lagrangian extensions). Originally, the notion of $T^{*}$-extension for Lie algebras was first introduced by Bordemann in [Bor]. Recently, it was studied in [BC] in the context of quasi-Frobenius Lie algebras, and then in [BM] in the context of quasi-Frobenius Lie superalgebras.
3.1. Flat connections and representations
Let $({\mathfrak{h}},\nabla)$ be a Lie superalgebra endowed with a flat connection $\nabla$. Since the connection $\nabla$ is flat, the map
$\mathfrak{h}\rightarrow\mathrm{End}(\mathfrak{h})\quad u\mapsto\nabla_{u},$
defines a representation. We define a map $\rho:\mathfrak{h}\rightarrow\mathrm{End}(\mathfrak{h}^{*})$, the dual representation, as follows (for all $u\in\mathfrak{h}$ and $\xi\in\mathfrak{h}^{*}$):
(6) $\rho:\mathfrak{h}\rightarrow\mathrm{End}(\mathfrak{h}^{*})\quad u\mapsto\rho(u),\text{ where }\rho(u)\cdot\xi=-(-1)^{|u||\xi|}\xi\circ\nabla_{u}.$
Similarly, we have a representation
(7) $\chi:\mathfrak{h}\rightarrow\mathrm{End}(\Pi(\mathfrak{h}^{*}))\quad u\mapsto\chi(u):=(-1)^{|u|}\,\Pi\circ\rho(u)\circ\Pi.$
Lemma 3.1 (Dual representation).
The maps (6) and (7) are indeed representations.
Proof.
See [BM]. ∎
3.2. Polarization and Lagrangian extensions
Following [BC] (for the supercase, see [BM]), a polarization for a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is a choice of a Lagrangian subalgebra $\mathfrak{l}$ of $(\mathfrak{g},\omega)$. A strong polarization of a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is a pair $(\mathfrak{a},N)$ consisting of a Lagrangian ideal $\mathfrak{a}\subset\mathfrak{g}$ and a complementary Lagrangian subspace $N\subset\mathfrak{g}$. The quadruple $(\mathfrak{g},\omega,\mathfrak{a},N)$ is referred to as a strongly polarized quasi-Frobenius Lie superalgebra.
By means of a 2-cocycle $\alpha\in Z^{2}(\mathfrak{h},\mathfrak{h}^{*})$ (resp. $\beta\in Z^{2}(\mathfrak{h},\Pi(\mathfrak{h}^{*}))$ we will construct a Lie superalgebra structure on $\mathfrak{g}:=\mathfrak{h}\oplus\mathfrak{h}^{*}$ (resp. $\mathfrak{g}:=\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*})$), called $T^{*}$-extension (resp. $\Pi T^{*}$-extension) (Lagrangian extension). By construction, we will see that these Lie superalgebras are strongly polarized quasi-Frobenius.
Recall that the space $\mathfrak{h}^{*}$ and $\Pi(\mathfrak{h}^{*})$ are $\mathfrak{h}$-modules by means of the representations (6) and (7).
On the space $\mathfrak{g}:=\mathfrak{h}\oplus\mathfrak{h}^{*}$: The brackets are defined as follows: (for all $u,v\in\mathfrak{h}$ and $\xi\in\mathfrak{h}^{*}$)
$\begin{array}[]{l}[u,v]_{\mathfrak{g}}:=[u,v]_{\mathfrak{h}}+\alpha(u,v),\quad[u,\xi]_{\mathfrak{g}}:=\rho(u)\cdot\xi.\end{array}$
An even form $\omega$ is defined on $\mathfrak{g}$ as follows: (for all $u+\xi,v+\zeta\in\mathfrak{h}\oplus\mathfrak{h}^{*}$)
(8) $\omega(u+\xi,v+\zeta)=\xi(v)-(-1)^{|\zeta||u|}\zeta(u).$
This construction will be referred to as the $T^{*}$-extension.
On the space $\mathfrak{k}:=\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*})$: The brackets are defined as follows: (for any $u,v\in\mathfrak{a}$)
$\begin{array}[]{l}[u,v]_{\mathfrak{g}}:=[u,v]_{\mathfrak{h}}+\beta(u,v),\quad[u,\Pi(\xi)]_{\mathfrak{k}}:=\chi(u)\cdot\Pi(\xi).\end{array}$
We define an odd form $\kappa$ on $\mathfrak{k}$ as follows: (for all $u+\Pi(\xi),v+\Pi(\zeta)\in\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*}$))
(9) $\kappa(u+\Pi(\xi),v+\Pi(\zeta))=\xi(v)-(-1)^{(|\zeta|+1)|u|}\zeta(u).$
This construction will be referred to as the $\Pi T^{*}$-extension.
Lemma 3.2 (Conditions on $\alpha$ and $\beta$).
(i) The form $\omega$ is closed on $\mathfrak{g}$ if and only if
(10) $(-1)^{|u||w|}\alpha(u,v)(w)+\circlearrowleft(u,v,w)=0\text{ for all $u,v,w\in\mathfrak{h}$}.$
(ii) The form $\kappa$ is closed on $\mathfrak{g}$ if and only if
(11) $(-1)^{|u||w|}\Pi\circ\beta(u,v)(w)+\circlearrowleft(u,v,w)=0\text{ for all $u,v,w\in\mathfrak{h}$}.$
Proof.
See, [BM]. ∎
We arrive to the following theorem proved in [BM].
Theorem 3.3 (Lagrangian or $T^{*}$- and $\Pi T^{*}$-extensions).
Let $(\mathfrak{h},\nabla)$ be a flat Lie superalgebra. To every 2-cocycle $\alpha\in Z^{2}(\mathfrak{h},\mathfrak{h}^{*})$ (resp. $\beta\in Z^{2}(\mathfrak{h},\Pi(\mathfrak{h}^{*}))$) which satisfies (10) (resp. (11)) one can canonically associate a strongly polarized quasi-Frobenius Lie superalgebra $\textup{(}\mathfrak{g},\omega,\mathfrak{h}^{*},\mathfrak{h})$ (resp. $\textup{(}\mathfrak{k},\kappa,\Pi(\mathfrak{h}^{*}),\mathfrak{h}$), where $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{h}^{*}$ (resp. $\mathfrak{k}=\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*})$) and $\omega$ (resp. $\kappa$) is given as in (8) (resp. (9)).
The converse of this theorem says that every orthosymplectic or periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ that has a Lagrangian ideal $\mathfrak{a}$ can be obtained as a $T^{*}$-extension or $\Pi T^{*}$-extension of a flat Lie superalgebra.
Let $(\mathfrak{g},\omega,\mathfrak{a},N)$ be a strongly polarized quasi-Frobenius Lie superalgebra. Here the form $\omega$ is either orthosymplectic or periplectic. Consider the quotient space $\mathfrak{h}=\mathfrak{g}/\mathfrak{a}$ (recall that $\mathfrak{a}^{\perp}=\mathfrak{a}$). We will construct a flat connection $\nabla$ on $\mathfrak{h}$ (see [BM]), superizing the construction of [BC]. For every $u,v\in\mathfrak{h}$, we denote by $\tilde{u}$ and $\tilde{v}$ their lift to $\mathfrak{g}$. We then write
(12) $\omega_{\mathfrak{h}}(\nabla_{u}v,a)=-(-1)^{|u||v|}\omega_{\mathfrak{g}}(\tilde{u},[\tilde{v},a]).$
The pair $(\mathfrak{h},\nabla)$ is called the quotient flat Lie superalgebra associated with the Lagrangian ideal $\mathfrak{a}$ of $(\mathfrak{g},\omega)$.
Theorem 3.4 (Converse of Theorem 3.3).
Let $(\mathfrak{g},\omega,\mathfrak{a},N)$ be a strongly polarized ortho-symplectic or periplectic quasi-Frobenius Lie superalgebra and $(\mathfrak{h},\nabla)$ be its associated quotient flat Lie superalgebra.
(i) If the form $\omega$ is orthosymplectic, then there exists $\alpha\in Z^{2}(\mathfrak{h},\mathfrak{h}^{*})$ satisfying Eq. (10), such that $(\mathfrak{g},\omega,\mathfrak{a},N)$ is isomorphic to the $T^{*}$-extension of $(\mathfrak{h},\nabla)$ by means of $\alpha$.
(ii) If the form $\omega$ is periplectic, then there exists $\beta\in Z^{2}(\mathfrak{h},\Pi(\mathfrak{h}^{*}))$ satisfying Eq. (11), such that $(\mathfrak{g},\omega,\mathfrak{a},N)$ is isomorphic to the $\Pi T^{*}$-extension of $(\mathfrak{h},\nabla)$ by means of $\beta$.
Proof.
See [BM]. ∎
4. The 4-dimensional Lie superalgebras as Lagrangian extensions
Using Theorem 3.3 and Theorem 3.4 above, we will determine whether the 4-dimensional real Lie superalgebras classified by Backhouse (see the Appendix) are isomorphic to a $T^{*}$-extension or to a $\Pi T^{*}$-extension of a smaller Lie superalgebra.
We will outline the proof and computation for the Lie superalgebra $D^{6}$ (row 2 in the table); however, we will only exhibit the connection $\nabla$ for the other Lie superalgebras.
  1. (1)
    The Lie superalgebra $D^{5}$: We correct here the description of the Lie superalgebra $D_{5}$ in [BM]. It is $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=e_{4}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ on $\mathfrak{h}$ is given by:
    $\nabla_{e_{4}}(e_{1})=-e_{4},\quad\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=0,\quad\nabla_{e_{4}}(e_{4})=0.$
  2. (2)
    The Lie superalgebra $D^{6}$: We will show first that this Lie superalgebra does have a Lagrangian ideal. Indeed, consider $\mathfrak{a}=\mathrm{Span}\{e_{3},e_{4}\}$. This is obviously an ideal because of the multiplication table. Let us show it is Lagrangian. We have
    $\mathfrak{a}^{\bot}=\{v\in D^{6}\;|\;\omega(v,x)=0,\forall x\in\mathfrak{a}\},$
    where the form we are considering is given by $\omega=e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}$. Let $v=l_{1}e_{1}+l_{2}e_{2}+l_{3}e_{3}+l_{4}e_{4}$. We have
    $0=\omega(v,e_{3})=-l_{1}\implies l_{1}=0,$
    $00=\omega(v,e_{4})=-l_{2}\implies l_{2}=$ 0 .
    This implies that $\mathfrak{a}^{\bot}=\mathfrak{a}$. The Lie superalgebra $\mathfrak{h}$ is given by $\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=0$. The connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{2})=-e_{2},\quad\nabla_{e_{2}}(e_{1})=-e_{2},\quad\nabla_{e_{2}}(e_{2})=e_{1}.$
    Now, we compute the torsion and curvature of $\nabla$, using Eqns. (4, 5). For the torsion, we have:
    $T(e_{1},e_{2}):=\nabla_{e_{1}}(e_{2})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}(e_{1})-[e_{1},e_{2}]_{\mathfrak{h}}=-e_{2}-(-e_{2})-0=0.$
    For the curvature, we have
    $\begin{array}[]{lcl}R^{\nabla}(e_{1},e_{2})(e_{1})&:=&\nabla_{e_{1}}\nabla_{e_{2}}(e_{1})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}\nabla_{e_{1}}(e_{1})-\nabla_{[e_{1},e_{2}]_{\mathfrak{h}}}(e_{1})\\[5.69054pt] &=&\nabla_{e_{1}}(-e_{2})-\nabla_{e_{2}}(-e_{1})=e_{2}-e_{2}=0.\end{array}$
    Similarly,
    $\begin{array}[]{lcl}R^{\nabla}(e_{1},e_{2})(e_{2})&:=&\nabla_{e_{1}}\nabla_{e_{2}}(e_{2})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}\nabla_{e_{1}}(e_{2})-\nabla_{[e_{1},e_{2}]_{\mathfrak{h}}}(e_{2})\\[5.69054pt] &=&\nabla_{e_{1}}(e_{1})-\nabla_{e_{2}}(-e_{2})=-e_{1}+e_{1}=0.\end{array}$
    Therefore, the connection is torsion-free and flat.
    Thus, The Lie superalgebra $D^{6}$ is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}$, where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{3}=\Pi(e_{1}^{*})|e_{4}=\Pi(e_{2}^{*})$. The 2-form $\beta\equiv 0$.
  3. (3)
    The Lie superalgebra $D^{7}_{pq}$: This Lie superalgebra is not quasi-Frobenius.
  4. (4)
    The Lie superalgebra $D^{7}_{-1q}$, with $q\leq 1$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{4}=\Pi(e_{1}^{*}),e_{3}=\Pi(e_{2}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-qe_{1},\quad\nabla_{e_{1}}(e_{2})=e_{2},\quad\nabla_{e_{2}}(e_{1})=0,\quad\nabla_{e_{2}}(e_{2})=0.$
  5. (5)
    The Lie superalgebra $D^{7}_{pq}$, with $p=-q,q\neq 0,-1$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=-pe_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-pe_{4},\quad\nabla_{e_{4}}(e_{1})=0,\quad\nabla_{e_{4}}(e_{4})=0.$
  6. (6)
    The Lie superalgebra $D^{7}_{pq}$, with $p=-q,q=-1$: In the case where the form is odd, then this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{3}=\Pi(e_{1}^{*}),e_{4}=\Pi(e_{2}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{2})=e_{2},\quad\nabla_{e_{2}}(e_{1})=0,\quad\nabla_{e_{2}}(e_{2})=0.$
    In the case where the form is even, then this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=-e_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=0,\quad\nabla_{e_{4}}(e_{4})=0.$
  7. (7)
    The Lie superalgebra $D^{8}_{p}$, with $p\neq 0$: This Lie superalgebra is not quasi-Frobenius.
  8. (8)
    The Lie superalgebra $D^{8}_{-1}$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{3}=\Pi(e_{1}^{*}),e_{4}=\Pi(e_{2}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=e_{1}-e_{2},\quad\nabla_{e_{1}}(e_{2})=e_{2},\quad\nabla_{e_{2}}(e_{1})=0,\quad\nabla_{e_{2}}(e_{2})=0.$
  9. (9)
    The Lie superalgebra $D^{9}_{pq}$, with $q>0$: This Lie superalgebra is not quasi-Frobenius.
  10. (10)
    The Lie superalgebra $D^{10}_{q}$ with $q\not=-1$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=qe_{4}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{2}=\Pi(e_{4}^{*})|e_{3}=-(q+1)\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-(q+1)e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=-(q+1)e_{4},\quad\nabla_{e_{4}}(e_{4})=0.$
  11. (11)
[…]
    $\begin{array}[]{ll}[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3}+\dfrac{1}{2}x_{4}e_{4},&[e_{2},X]=-x_{1}e_{2},\\[5.69054pt] [e_{3},X]=-\dfrac{1}{2}x_{1}e_{3}+x_{3}e_{2},&[e_{4},X]=-\dfrac{1}{2}x_{1}e_{4}+x_{4}e_{2}.\end{array}$
    If $x_{1}\neq 0$, from the second bracket we get that $e_{2}\in\mathfrak{a}$, followed by $e_{3}\in\mathfrak{a}$ and $e_{4}\in\mathfrak{a}$, which implies that the ideal is not of dimension 2. We are in the case where $x_{1}=0$ (and similarly $y_{1}=0$).
    $\begin{array}[]{ll}[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3}+\dfrac{1}{2}x_{4}e_{4},&[e_{2},X]=0,\\[5.69054pt] [e_{3},X]=x_{3}e_{2},&[e_{4},X]=x_{4}e_{2}.\end{array}$
    We consider the following cases:
    1. (a)
      If $x_{3}=x_{4}=0$. In this case, $e_{2}\in\mathfrak{a}$ and we can write $X=e_{2}$ and $Y=y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$. We can simplify and write $Y=z_{3}e_{3}+z_{4}e_{4}$. Since [^1] As in [BM, BEM, BE], we adopt the following convention: $\langle e_{i}^{*},e_{j}\rangle=\delta_{ij}$ and $\langle e_{i}^{*}\otimes e_{j}^{*},e_{k}\otimes e_{l}\rangle=(-1)^{|e_{k}||e_{j}|}\langle e_{i}^{*},e_{k}\rangle\langle e_{j}^{*},e_{l}\rangle$. $\omega(Y,Y)=z_{3}^{2}+z_{4}^{2}\not=0,$ it follows that $\mathfrak{a}$ is not Lagrangian.
    2. (b)
      If $x_{3}=0,$ and $x_{4}\neq 0$. Looking at brackets again we get $[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{4}e_{4},$ and $[e_{4},X]=x_{4}e_{2}$, which implies $e_{2},e_{4}\in\mathfrak{a}$, so $\mathfrak{a}=\mathrm{Span}\{e_{2},e_{4}\}$. We have to check that $\mathfrak{a}$ is not Lagrangian. Indeed, $\omega(e_{4},e_{4})=1$.
    3. (c)
      If $x_{3}\neq 0,$ and $x_{4}=0$. This case will be similar to the previous one since the equations are symmetrical in $x_{3}$ and $x_{4}$, which implies that no Lagrangian ideal exists.
    4. (d)
      If $x_{3}\neq 0,$ and $x_{4}\neq 0$. Since all the previous cases showed that we can’t find a Lagrangian ideal, and since the proof is symmetrical in X and Y without loss of generality we can state that $y_{3}\neq 0,y_{4}\neq 0$. This implies that $e_{2}\in\mathfrak{a}$ and that $x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3}+\dfrac{1}{2}x_{4}e_{4}\in\mathfrak{a}$. We can write $X=e_{2}$ and $Y=z_{3}e_{3}+z_{4}e_{4}$. Once again, $\omega(Y,Y)=z_{3}^{2}+z_{4}^{2}\not=0$ which implies that $\mathfrak{a}$ is not Lagrangian.
  12. (12)
[…]
    $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(x)=-\frac{1}{2}x,\quad\nabla_{x}(e_{1})=-x,\quad\nabla_{x}(x)=0.$
  13. (13)
[…]
  14. (14)
    The Lie superalgebra $(D^{7}_{1-p,p})$, with $p\leq\dfrac{1}{2}$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=(1-p)e_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-pe_{4},\quad\nabla_{e_{4}}(e_{1})=-e_{4},\quad\nabla_{e_{4}}(e_{4})=0.$
  15. (15)
    The Lie superalgebra $(D^{8}_{1/2})$: This Lie superalgebra is not quasi-Frobenius.
  16. (16)
    The Lie superalgebra $(D^{9}_{1/2,p})$, with $p>0$: This Lie superalgebra does not admit a 2-dimensional Lagrangian ideal. Indeed, let us suppose there exists one, denoted by $\mathfrak{a}=\mathrm{Span}\{X,Y\}$. Let us write $X=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+x_{4}e_{4}$ and $Y=y_{1}e_{1}+y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$. We compute the brackets $[e_{i},X]$ for all $i\in\{1,2,3,4\}$. We get
    $\begin{array}[]{l}[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3}-x_{3}pe_{4}+x_{4}pe_{3}+\dfrac{1}{2}x_{4}e_{4},\\[5.69054pt] [e_{2},X]=-x_{1}e_{2},\\[5.69054pt] [e_{3},X]=-\dfrac{1}{2}x_{1}e_{3}+px_{1}e_{4}+x_{3}e_{2},\\[8.53581pt] [e_{4},X]=-\dfrac{1}{2}x_{1}e_{4}-px_{1}e_{3}+x_{4}e_{2}.\end{array}$
    We will analyze a few cases. If $x_{1}\neq 0$, given the second bracket we get that $e_{2}\in\mathfrak{a}$ and according to the last brackets $e_{3}\in\mathfrak{a}$ and $e_{4}\in\mathfrak{a}$. This case is not possible since $\mathfrak{a}$ has dimension 2. This implies that $x_{1}=0$. We get then:
    $\begin{array}[]{ll}[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3}-x_{3}pe_{4}+x_{4}pe_{3}+\dfrac{1}{2}x_{4}e_{4},&[e_{2},X]=0,\\[5.69054pt] [e_{3},X]=x_{3}e_{2},&[e_{4},X]=x_{4}e_{2}.\end{array}$
    For all the cases $x_{3},x_{4}=0$ or $\neq 0$ we get that $e_{2}\in\mathfrak{a}$. We can write $X=e_{2}$ and $Y=z_{3}e_{3}+z_{4}e_{4}$. As $\omega(Y,Y)\not=0$, it follows that $\mathfrak{a}$ is not Lagrangian.
    Thus, this Lie superalgebra does not admit a 2-dimensional Lagrangian ideal.
  17. (17)
    The Lie superalgebra $(D^{10}_{0})^{1}$: In the case where the form is odd, this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=-e_{4},\quad\nabla_{e_{4}}(e_{4})=\frac{1}{2}e_{1}.$
    In the case where the form is even, this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=2e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=-e_{4},\quad\nabla_{e_{4}}(e_{4})=\frac{1}{2}e_{1}.$
  18. (18)
    The Lie superalgebra $(D^{10}_{0})^{2}$: In the case where the form is odd, this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=-e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=-e_{4},\quad\nabla_{e_{4}}(e_{4})=-\frac{1}{2}e_{1}.$
    In the case where the form is even, this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=-e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=2e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=-e_{4},\quad\nabla_{e_{4}}(e_{4})=-\frac{1}{2}e_{1}.$
  19. (19)
    The Lie superalgebra $(2A_{1,1}+2A)^{1}$: This Lie superalgebra is not quasi-Frobenius.
  20. (20)
    The Lie superalgebra $(2A_{1,1}+2A)^{2}$: This Lie superalgebra is not quasi-Frobenius.
  21. (21)
    The Lie superalgebra $(2A_{1,1}+2A)^{3}_{p}$, for $p=\frac{1}{2}$: This Lie superalgebra is a $\Pi T^{*}$-extension of the abelian Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{3},e_{4}\}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{1}=-\Pi(e_{4}^{*}),\;e_{2}=\Pi(e_{3}^{*})$. The connection $\nabla$ is trivial, and the 2-form $\beta$ is given by
    $\beta=\frac{1}{2}\Pi(e_{4}^{*})\otimes e_{3}^{*}\wedge e_{3}^{*}-\frac{1}{2}\Pi(e_{3}^{*})\otimes e_{4}^{*}\wedge e_{4}^{*}-\frac{1}{2}(\Pi(e_{3}^{*})-\Pi(e_{4}^{*}))\otimes e_{3}^{*}\wedge e_{4}^{*}.$
  22. (22)
    The Lie superalgebra $(2A_{1,1}+2A)^{4}_{p}$: This Lie superalgebra is not quasi-Frobenius.
  23. (23)
    The Lie superalgebra $(C^{1}_{1}+A)$: This Lie superalgebra is a $T^{*}$-extension of the abelian Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|x:=e_{3}-2e_{4}\}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=\frac{1}{2}e_{1}^{*}|e_{3}=x^{*}$. The 2-form $\alpha$ is given by:
    $\alpha=e^{*}_{1}\otimes(x^{*}\wedge x^{*})+x^{*}\otimes(e^{*}_{1}\wedge x^{*})$
    The connection $\nabla$ is given by:
    $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(x)=-x,\quad\nabla_{x}(e_{1})=-x,\quad\nabla_{x}(x)=0.$
  24. (24)
    The Lie superalgebra $(C^{1}_{1/2}+A)$: This Lie superalgebra does not admit a 2-dimensional Lagrangian ideal. Indeed, let us suppose there is one, denoted by $\mathfrak{a}=\mathrm{Span}\{X,Y\}$. Let us write $X=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+x_{4}e_{4}$ and $Y=y_{1}e_{1}+y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$. We compute the brackets $[e_{i},X]$ for all $i\in\{1,2,3,4\}$. We get:
    $\begin{array}[]{l}[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3},\quad[e_{2},X]=-x_{1}e_{2},\quad[e_{3},X]=-\dfrac{1}{2}x_{1}e_{3}+x_{3}e_{2},\quad[e_{4},X]=0.\end{array}$
    We will analyze a few cases.
    If $x_{1}\neq 0$, given the second bracket implies that $e_{2}\in\mathfrak{a}$ and according to the third bracket $e_{3}\in\mathfrak{a}$, which implies that $\mathfrak{a}=\mathrm{Span}\{e_{2},e_{3}\}$. Since $\omega(e_{3},e_{3})\not=0$, it follows that $\mathfrak{a}$ is not Lagrangian.
    If $x_{1}=0$, and $x_{3}\neq 0$, we get that $e_{2}\in\mathfrak{a}$, followed by $e_{3}\in\mathfrak{a}$, so $\mathfrak{a}=\mathrm{Span}\{e_{2},e_{3}\}$, which we showed is not possible. It follows that $x_{3}=0$, and using the same reasoning as above, we also conclude that for $Y$, the scalars $y_{1}=y_{3}=0$. Looking at the brackets, we either have $x_{2}=0$ or $e_{2}\in\mathfrak{a}$. If $x_{2}=0$, then $X=x_{4}e_{4}$, and $Y=y_{2}e_{2}+y_{4}e_{4}$. We can simplify and write $X=e_{4}$, and $Y=e_{2}$. Since $\omega(e_{4},e_{4})\not=0$, it follows that $\mathfrak{a}$ is not Lagrangian. On the other hand, if $x_{2}\neq 0$, then $X=x_{2}e_{2}+x_{4}e_{4}$, and $Y=y_{2}e_{2}+y_{4}e_{4}$. We can simplify and write $X=e_{2}+z_{4}e_{4}$, and $Y=e_{4}$. Again, $\mathfrak{a}$ is not Lagrangian.
    Thus, this Lie superalgebra does not admit a 2-dimensional Lagrangian ideal.
  25. (25)
    The Lie superalgebra $(C^{2}_{-1}+A)$: This Lie superalgebra is not quasi-Frobenuis.
  26. (26)
    The Lie superalgebra $(C^{3}+A)$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{2}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=e_{2}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{1}=2e_{2}^{*}\,|\,e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
    $\nabla_{e_{2}}(e_{2})=0,\quad\nabla_{e_{2}}(e_{4})=0,\quad\nabla_{e_{4}}(e_{2})=0,\quad\nabla_{e_{4}}(e_{4})=\frac{1}{2}e_{2}.$
  27. (27)
    The Lie superalgebra $(C^{5}_{0}+A)$: This Lie superalgebra is not quasi-Frobenuis.
  28. (28)
    The Lie superalgebra $D^{1}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
  29. (29)
    The Lie superalgebra $D^{2}_{q},\text{ with }q\not=-1,0$: This Lie superalgebra is not quasi-Frobenuis.
  30. (30)
    The Lie superalgebra $D^{2}_{-1}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
  31. (31)
    The Lie superalgebra $D^{3}_{pq}$: This Lie superalgebra is not quasi-Frobenuis.
  32. (32)
    The Lie superalgebra $D^{11}_{pq}\text{ with }0<|p|\leq|q|\leq 1$: This Lie superalgebra is not quasi-Frobenuis.
  33. (33)
    The Lie superalgebra $D^{11}_{pq}\text{ with }(p,q)=(-1,-1)$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
  34. (34)
    The Lie superalgebra $D^{11}_{pq}\text{ with }0<|p|\leq 1,p=-q$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
  35. (35)
    The Lie superalgebra $D^{11}_{pq}\text{ with }0<|p|<1,q=-1$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
  36. (36)
    The Lie superalgebra $D^{12}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
  37. (37)
    The Lie superalgebra $D^{13}_{p},\text{ with }p\not=0$: This Lie superalgebra is not quasi-Frobenuis.
  38. (38)
    The Lie superalgebra $D^{13}_{-1}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
  39. (39)
    The Lie superalgebra $D^{14}_{pq},\text{ with }p\not=0,q\geq 0$: This Lie superalgebra is not quasi-Frobenuis.
  40. (40)
    The Lie superalgebra $D^{15}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
  41. (41)
    The Lie superalgebra $D^{16}$: This Lie superalgebra is in fact not quasi-Frobenuis.
  42. (42)
    The Lie superalgebra $(A_{3,1}+A)$: This Lie superalgebra is not quasi-Frobenuis.
  43. (43)
    The Lie superalgebra $(D^{3}_{p,-1/2}),\text{ with }p\not=0$: This Lie superalgebra is not quasi-Frobenuis.
  44. (44)
    The Lie superalgebra $(D^{3}_{-1/2,-1/2})$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
  45. (45)
    The Lie superalgebra $(D^{2}_{-1/2})^{1}$: This Lie superalgebra is not quasi-Frobenuis.
  46. (46)
    The Lie superalgebra $(D^{2}_{-1/2})^{2}$: This Lie superalgebra is not quasi-Frobenuis.
  47. (47)
    The Lie superalgebra $(A_{1,1}+3A)^{1}$: This Lie superalgebra is not quasi-Frobenuis.
  48. (48)
    The Lie superalgebra $(A_{1,1}+3A)^{2}$: This Lie superalgebra is not quasi-Frobenuis.
5. Left-symmetric structures on the 4-dimensional real Lie superalgebras
In the list of 4-dimensional real Lie superalgebras, we will list left-symmetric structures that are compatible with the Lie structure. With the exception of $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$, all of these structures are Novikov. Moreover, we show that each of these Lie superalgebras admit a Balinsky-Novikov structure. In line with standard practice, only non-zero products are displayed.
  1. (1)
    The Lie superalgebra $D^{5}$: A Novikov structure is given by:
    $e_{1}\cdot e_{3}=e_{3},\quad e_{1}\cdot e_{4}=e_{4},\quad e_{2}\cdot e_{4}=e_{3}.$
    Additionally, this structure is also Balinsky-Novikov. As a matter of fact, the non-zero associators are given by:
    $(e_{1},e_{1},e_{4})=-e_{4},\quad(e_{2},e_{1},e_{4})=-e_{3},\quad(e_{1},e_{1},e_{3})=-e_{3},\quad(e_{1},e_{2},e_{4})=-e_{3}.$
  2. (2)
    The Lie superalgebra $D^{6}$: A Novikov structure is given by:
    $e_{1}\cdot e_{3}=e_{3},\quad e_{1}\cdot e_{4}=e_{4},\quad e_{2}\cdot e_{3}=-e_{4},\quad e_{2}\cdot e_{4}=e_{3}.$
    Additionally, this structure is also Balinsky-Novikov.
  3. (3)
    The Lie superalgebra $D^{7}_{pq}$ with ($p\geq q,\;pq\not=0$): A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
    $e_{1}\cdot e_{1}=\gamma e_{2},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{1}\cdot e_{3}=pe_{3},\quad e_{1}\cdot e_{4}=qe_{4}.$
    Additionally, this structure is also Balinsky-Novikov.
  4. (4)
    The Lie superalgebra $D^{8}_{p}$ with $p\not=0$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
    $e_{1}\cdot e_{1}=\gamma e_{2},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{1}\cdot e_{3}=pe_{3},\quad e_{1}\cdot e_{4}=e_{3}+pe_{4}.$
    Additionally, this structure is also Balinsky-Novikov.
  5. (5)
    The Lie superalgebra $D^{9}_{pq}$ with $q>0$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=(p+\gamma)e_{3}-qe_{4},&e_{1}\cdot e_{4}=qe_{3}+(p+\gamma)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{4}\cdot e_{1}=\gamma e_{4}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R}$):
    $\begin{array}[]{lllll}e_{1}\cdot e_{1}=\gamma e_{2},&e_{1}\cdot e_{2}=e_{2},&e_{1}\cdot e_{3}=pe_{3}-qe_{4},&e_{1}\cdot e_{4}=qe_{3}+pe_{4}.\end{array}$
  6. (6)
    The Lie superalgebra $D^{10}_{q}:$ A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=-qe_{1}+\gamma(q-1)e_{2},&e_{1}\cdot e_{2}=(1-q)e_{2},&e_{1}\cdot e_{3}=e_{3},&e_{1}\cdot e_{4}=\gamma e_{3},\\[5.69054pt] e_{2}\cdot e_{1}=-qe_{2},&e_{3}\cdot e_{1}=-qe_{3},&e_{4}\cdot e_{1}=\gamma e_{3}-qe_{4},&e_{4}\cdot e_{2}=-e_{3}.\end{array}$
    A Balinsky-Novikov structure is given by
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=-e_{1},&e_{1}\cdot e_{3}=\frac{1}{2}(1+2q)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(2q-1)e_{4},&e_{2}\cdot e_{1}=-e_{2},\\[5.69054pt] e_{2}\cdot e_{4}=e_{3},&e_{3}\cdot e_{1}=-e_{3},&e_{4}\cdot e_{1}=-e_{4}.\end{array}$
  7. (7)
[…]
    $\begin{array}[]{lllll}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\frac{1}{2}e_{2},&e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},&e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},&e_{3}\cdot e_{3}=\frac{1}{2}e_{2}\\[5.69054pt] e_{3}\cdot e_{4}=\gamma e_{2},&e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},&e_{4}\cdot e_{3}=-\gamma e_{2},&e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3},&e_{1}\cdot e_{4}=\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=(\gamma-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{4}=e_{2}.\end{array}$
  8. (8)
[…]
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\frac{1}{2}e_{2},&e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},&e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\\[5.69054pt] e_{3}\cdot e_{3}=\frac{1}{2}e_{2},&e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=-\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(1+\gamma)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{4}=-e_{2}.\end{array}$
  9. (9)
[…]
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\frac{1}{2}e_{2},&e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},&e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\\[5.69054pt] e_{3}\cdot e_{3}=\frac{1}{2}e_{2},&e_{4}\cdot e_{1}=-\frac{1}{2}e_{4}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(1+\gamma)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4}.\end{array}$
  10. (10)
    The Lie superalgebra $(D^{7}_{1-p,p})$ with $p\leq\frac{1}{2}:$ A Novikov structure is given by (where $\lambda\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=-pe_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1-p)e_{2},&e_{1}\cdot e_{4}=(1-2p)e_{4},&e_{2}\cdot e_{1}=-pe_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-pe_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=-pe_{4}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(2p+\gamma-1)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(1-2p+\gamma)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{4}\cdot e_{1}=(\gamma-1)e_{4}\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
  11. (11)
    The Lie superalgebra $(D^{8}_{1/2}):$ A Novikov structure is given by (where $\lambda\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\frac{1}{2}e_{2},&e_{1}\cdot e_{4}=e_{3},&e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},&e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=e_{3}+\frac{1}{2}(1+\gamma)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{4}\cdot e_{1}=\gamma e_{4},&e_{4}\cdot e_{4}=e_{2}.\\[5.69054pt] \end{array}$
  12. (12)
    The Lie superalgebra $(D^{9}_{1/2,p})$ with $p>0:$ A Novikov structure is given by (where $\lambda\in\mathbb{R}$):
    $\begin{array}[]{lllll}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\frac{1}{2}e_{2},&e_{1}\cdot e_{3}=-pe_{4},&e_{1}\cdot e_{4}=pe_{3},&e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},&e_{3}\cdot e_{3}=\frac{1}{2}e_{2},&e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3}-pe_{4},&e_{1}\cdot e_{4}=pe_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=(\gamma-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{4}=e_{2}.\end{array}$
  13. (13)
    The Lie superalgebra $(D^{10}_{0})^{1}:$ An LSSA structure is given by (where $\gamma\in\mathbb{R}$):
    $\begin{array}[]{lllll}e_{1}\cdot e_{1}=-e_{1},&e_{1}\cdot e_{4}=-e_{4},&e_{2}\cdot e_{1}=-e_{2},&e_{2}\cdot e_{4}=(1+\gamma)e_{3},&e_{3}\cdot e_{1}=-e_{3},\\[5.69054pt] e_{3}\cdot e_{4}=-\frac{1}{2}e_{2},&e_{4}\cdot e_{1}=-e_{4},&e_{4}\cdot e_{2}=\gamma e_{3},&e_{4}\cdot e_{4}=\frac{1}{2}e_{1}.\end{array}$
    Claim. There is no Novikov structure on $(D^{10}_{0})^{1}$ compatible with the Lie structure.
    Proof.
    Let us write:
    $\begin{array}[]{lll}e_{i}\cdot e_{j}=\sum_{i,j}\lambda^{1}_{ij}e_{1}+\sum_{i,j}\lambda^{2}_{ij}e_{2}&\text{ if }&|e_{i}|+|e_{j}|={\bar{0}},\text{ and }\\[5.69054pt] e_{i}\cdot e_{j}=\sum_{i,j}\lambda^{3}_{ij}e_{3}+\sum_{i,j}\lambda^{4}_{ij}e_{4}&\text{ if }&|e_{i}|+|e_{j}|={\bar{1}}.\end{array}$
    The compatibility condition implies that
    $\begin{array}[]{l}\lambda^{3}_{13}=\lambda_{31}^{3}+1,\quad\lambda^{4}_{13}=\lambda_{31}^{4},\quad\lambda^{2}_{12}=\lambda_{21}^{2}+1,\quad\lambda^{1}_{21}=\lambda_{12}^{1},\quad\lambda^{3}_{41}=\lambda_{14}^{3},\quad\lambda^{4}_{41}=\lambda_{14}^{4},\\[8.53581pt] \lambda^{3}_{24}=\lambda_{42}^{3}+1,\quad\lambda^{4}_{42}=\lambda_{24}^{4},\quad\lambda^{3}_{32}=\lambda_{23}^{3},\quad\lambda^{4}_{32}=\lambda_{23}^{4},\quad\lambda^{1}_{33}=0,\quad\lambda^{2}_{33}=0,\quad\lambda^{1}_{44}=\frac{1}{2},\\[8.53581pt] \lambda^{2}_{44}=0\quad\lambda^{2}_{34}=-\lambda^{2}_{43}-\frac{1}{2},\quad\lambda^{1}_{43}=-\lambda^{1}_{34}.\end{array}$
    Let us put
    $\begin{array}[]{lcl}N(x,y,z):&=&(z\cdot x)\cdot y-(-1)^{|x||y|}(z\cdot y)\cdot x,\\[8.53581pt] T(x,y,z):&=&(x,y,z)-(-1)^{|x||y|}(y,x,z).\end{array}$
    A direct calculation gives
    $N(e_{4},e_{4},e_{4})=0\implies\lambda^{3}_{14}=\lambda_{14}^{4}=0,\text{ also }N(e_{4},e_{1},e_{4})=0\implies\lambda^{1}_{11}=\lambda_{11}^{2}=0.$
    Similarly,
    $N(e_{1},e_{3},e_{1})=0\implies\lambda^{3}_{31}=\lambda_{13}^{4}=0,\text{ also }T(e_{2},e_{1},e_{2})=0\implies\lambda^{2}_{21}=\lambda_{12}^{1}=\lambda^{1}_{22}=\lambda_{22}^{2}=0.$
    Additionally,
    $N(e_{4},e_{2},e_{1})=0\implies\lambda^{4}_{24}=\lambda_{42}^{3}=-1\text{ also }N(e_{4},e_{2},e_{4})=0\implies\lambda^{1}_{34}=\lambda_{43}^{2}=0.$
    Now, $N(e_{3},e_{4},e_{4})=\frac{1}{2}e_{3}$ which is a not zero. Therefore, it cannot be Novikov. ∎
    However, there is a Balinsky-Novikov structure given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(2+\gamma)e_{3},&e_{1}\cdot e_{4}=-\lambda e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{2}\cdot e_{4}=-\gamma e_{3},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{4}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{4}\cdot e_{1}=-2\lambda e_{3}+\gamma e_{4},&e_{4}\cdot e_{2}=-2(1+\gamma)e_{3}&e_{4}\cdot e_{3}=-\frac{1}{2}e_{2},&e_{4}\cdot e_{4}=e_{1}.\end{array}$
  14. (14)
    The Lie superalgebra $(D^{10}_{0})^{2}:$ AAn LSSA structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{lllll}e_{1}\cdot e_{1}=-e_{1}-2\lambda\gamma e_{2},&e_{1}\cdot e_{4}=\gamma e_{3}-e_{4},&e_{2}\cdot e_{1}=-e_{2},&e_{2}\cdot e_{4}=e_{3},&e_{3}\cdot e_{1}=-e_{3},\\[5.69054pt] e_{3}\cdot e_{4}=\frac{1}{2}(1-2\lambda)e_{2},&e_{4}\cdot e_{1}=\gamma e_{3}-e_{4},&e_{4}\cdot e_{3}=\lambda e_{2},&e_{4}\cdot e_{4}=-\frac{1}{2}e_{1}.\end{array}$
    Claim. There is no Novikov structure on $(D^{10}_{0})^{2}$ compatible with the Lie structure.
    However, there is a Balinsky-Novikov structure given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=-\lambda e_{3}+\frac{1}{2}(\gamma-1)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{2}\cdot e_{4}=(1-\gamma)e_{3},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=\frac{1}{2}e_{2},\\[5.69054pt] e_{4}\cdot e_{1}=-2\lambda e_{3}+(\gamma-1)e_{4},&e_{4}\cdot e_{2}=-2\gamma e_{3}&e_{4}\cdot e_{3}=\frac{1}{2}e_{2},&e_{4}\cdot e_{4}=-e_{1}.\end{array}$
  15. (15)
    The Lie superalgebra $(2A_{1,1}+2A)^{1}:$ A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R})$:
    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\lambda e_{1}+\gamma e_{2},\quad e_{4}\cdot e_{3}=-\lambda e_{1}-\gamma e_{2},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1},&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3},&e_{2}\cdot e_{2}=2\lambda e_{2},&e_{2}\cdot e_{4}=\lambda e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{1},&e_{4}\cdot e_{2}=2\lambda e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
  16. (16)
    The Lie superalgebra $(2A_{1,1}+2A)^{2}:$ A Novikov structure is given by (where $\gamma\in{\mathbb{R}})$::
    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(1-\gamma_{)}e_{1},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=2\gamma e_{1},&e_{1}\cdot e_{2}=2\gamma e_{1},&e_{1}\cdot e_{3}=\gamma e_{3},&e_{1}\cdot e_{4}=\gamma e_{3},\\[5.69054pt] e_{2}\cdot e_{1}=2\gamma e_{1},&e_{2}\cdot e_{2}=-2(\lambda-\gamma)e_{1}+2\lambda e_{2},&e_{2}\cdot e_{3}=\gamma e_{3},&e_{2}\cdot e_{4}=(\gamma-\lambda)e_{3}+\lambda e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=2\gamma e_{3}&e_{3}\cdot e_{2}=2\gamma e_{3},&e_{3}\cdot e_{3}=e_{1}&e_{3}\cdot e_{4}=e_{1},\\[5.69054pt] e_{4}\cdot e_{1}=2\gamma e_{3}&e_{4}\cdot e_{2}=-2(\lambda-\gamma)e_{3}+2\lambda e_{4}&e_{4}\cdot e_{3}=e_{1}&e_{4}\cdot e_{4}=e_{2}.\end{array}$
  17. (17)
    The Lie superalgebra $(2A_{1,1}+2A)^{3}_{p}:$ A Novikov structure is given by (where $\gamma\in{\mathbb{R}})$:
    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(p-\gamma_{)}e_{1}+pe_{2},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{lll}e_{1}\cdot e_{1}=\frac{2}{p}(p^{2}\gamma+\lambda(2p^{2}-1))e_{1}+2p\lambda e_{2},&e_{1}\cdot e_{2}=-2p\lambda e_{1}+2p\gamma e_{2},\\[5.69054pt] e_{1}\cdot e_{3}=\frac{1}{p}(p^{2}\gamma+\lambda(p^{2}-1))e_{3}+\lambda e_{4},&e_{1}\cdot e_{4}=-\lambda e_{3}+p(\gamma+\lambda)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=-2p\lambda e_{1}+2p\gamma e_{2},&e_{2}\cdot e_{2}=-2p\gamma e_{1}-\frac{2}{p}(p^{2}\lambda+(2p^{2}-1)\gamma)e_{2},\\[5.69054pt] e_{2}\cdot e_{3}=-p(\lambda+\gamma)e_{3}+\gamma e_{4},&e_{2}\cdot e_{4}=-\gamma e_{3}+\frac{1}{p}((1-p^{2})\gamma-p^{2}\lambda)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\frac{2}{p}(p^{2}\gamma+(p^{2}-1)\lambda)e_{3}+2\lambda e_{4},&e_{3}\cdot e_{2}=-2p(\gamma+\lambda)e_{3}+2\gamma e_{4},\\[5.69054pt] e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=pe_{1}+pe_{2},\\[5.69054pt] e_{4}\cdot e_{1}=-2\lambda e_{3}+2p(\gamma+\lambda)e_{4}&e_{4}\cdot e_{2}=-2\gamma e_{3}-\frac{2}{p}((p^{2}-1)\gamma+p^{2}\lambda))e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=pe_{1}+pe_{2},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
  18. (18)
    The Lie superalgebra $(2A_{1,1}+2A)^{4}_{p}:$ A Novikov structure is given by (where $\gamma\in{\mathbb{R}})$:
    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(p-\gamma)e_{1}-pe_{2},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{lll}e_{1}\cdot e_{1}=-\frac{2}{p}(p^{2}\gamma-\lambda(2p^{2}+1))e_{1}-2p\lambda e_{2},&e_{1}\cdot e_{2}=2p\lambda e_{1}-2p\gamma e_{2},\\[5.69054pt] e_{1}\cdot e_{3}=\frac{1}{p}(-p^{2}\gamma+\lambda(p^{2}+1))e_{3}+\lambda e_{4},&e_{1}\cdot e_{4}=\lambda e_{3}+p(-\gamma+\lambda)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2p\lambda e_{1}-2p\gamma e_{2},&e_{2}\cdot e_{2}=2p\gamma e_{1}-\frac{2}{p}(-p^{2}\lambda+(2p^{2}+1)\gamma)e_{2},\\[5.69054pt] e_{2}\cdot e_{3}=p(-\gamma+\lambda)e_{3}+\gamma e_{4},&e_{2}\cdot e_{4}=\gamma e_{3}+\frac{1}{p}((-1-p^{2})\gamma+p^{2}\lambda)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\frac{2}{p}(-p^{2}\gamma+(p^{2}+1)\lambda)e_{3}+2\lambda e_{4},&e_{3}\cdot e_{2}=-2p(\gamma-\lambda)e_{3}+2\gamma e_{4},\\[5.69054pt] e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=pe_{1}-pe_{2},\\[5.69054pt] e_{4}\cdot e_{1}=2\lambda e_{3}-2p(\gamma-\lambda)e_{4}&e_{4}\cdot e_{2}=2\gamma e_{3}-\frac{2}{p}((p^{2}+1)\gamma-p^{2}\lambda))e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=pe_{1}-pe_{2},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
  19. (19)
    The Lie superalgebra $(C_{1}^{1}+A):$ A Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R}):$
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\lambda e_{1}+\gamma e_{2},&e_{1}\cdot e_{2}=(1+\lambda)e_{2},&e_{1}\cdot e_{3}=(1+\lambda)e_{3},&e_{1}\cdot e_{4}=\lambda e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\lambda e_{2},&e_{3}\cdot e_{1}=\lambda e_{3},&e_{3}\cdot e_{4}=-\lambda e_{2},&e_{4}\cdot e_{1}=\lambda e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=(1+\lambda)e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(\gamma-1)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=(\gamma-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
  20. (20)
    The Lie superalgebra $(C_{1/2}^{1}+A):$ A Novikov structure is given by:
    $\begin{array}[]{l}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\gamma e_{2},\quad e_{1}\cdot e_{2}=\frac{1}{2}e_{2},\quad e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\quad e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\quad e_{3}\cdot e_{3}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2}&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4}.\end{array}$
  21. (21)
    The Lie superalgebra $(C_{-1}^{2}+A):$ A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R})$:
    $\begin{array}[]{l}e_{1}\cdot e_{1}=(2\lambda-1)e_{1}+\gamma e_{2},\quad e_{1}\cdot e_{2}=(2\lambda-1)e_{2},\quad e_{1}\cdot e_{3}=2\lambda e_{3},\quad e_{1}\cdot e_{4}=2(\lambda-1)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(2\lambda-1)e_{2},\quad e_{3}\cdot e_{1}=(2\lambda-1)e_{3},\quad e_{3}\cdot e_{4}=(1-\lambda)e_{2},\quad e_{4}\cdot e_{1}=(2\lambda-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=\lambda e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{1}{2}(2+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(\gamma-2)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
  22. (22)
    The Lie superalgebra $(C^{3}+A):$ A Novikov structure is given by:
    $\begin{array}[]{c}e_{4}\cdot e_{1}=-e_{3},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3},&e_{1}\cdot e_{4}=e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{4}\cdot e_{1}=\gamma e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
  23. (23)
    The Lie superalgebra $(C_{0}^{5}+A):$ A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R})$:
    $\begin{array}[]{l}e_{1}\cdot e_{1}=2\lambda e_{1}+\gamma e_{2},\quad e_{1}\cdot e_{2}=2\lambda e_{2},\quad e_{1}\cdot e_{3}=2\lambda e_{3}-e_{4},\quad e_{1}\cdot e_{4}=e_{3}+2\lambda e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2\lambda e_{2},\quad e_{3}\cdot e_{1}=2\lambda e_{3},\quad e_{3}\cdot e_{3}=\frac{1}{2}e_{2},\quad e_{3}\cdot e_{4}=-\lambda e_{2},\quad e_{4}\cdot e_{1}=2\lambda e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=\lambda e_{2},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3}-e_{4},&e_{1}\cdot e_{4}=e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{4}=e_{2}.\end{array}$
  24. (24)
    The Lie superalgebra $D^{1}$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in\mathbb{R})$:
    $\begin{array}[]{l}e_{2}\cdot e_{2}=\lambda e_{1}+\gamma e_{3},\quad e_{2}\cdot e_{3}=(1+\mu)e_{1},\quad e_{2}\cdot e_{4}=e_{4},\quad e_{3}\cdot e_{2}=\mu e_{1}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{2}\cdot e_{2}=\gamma e_{1}+\lambda e_{3},&e_{2}\cdot e_{4}=e_{4},&e_{3}\cdot e_{2}=-e_{1}.\end{array}$
  25. (25)
    The Lie superalgebra $D^{2}_{q}$ with $q\not=0$: A Novikov structure is given by:
[…]
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{3}\cdot e_{1}=-e_{1},&e_{3}\cdot e_{2}=-e_{1}-e_{2},&e_{3}\cdot e_{3}=\lambda e_{1}+\gamma e_{2}&e_{3}\cdot e_{4}=qe_{4}.\end{array}$
  26. (26)
    The Lie superalgebra $D^{3}_{pq}$ with $pq\not=0$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in{\mathbb{R}}$:
    $\begin{array}[]{l}e_{1}\cdot e_{3}=(p+\gamma)e_{1},\quad e_{2}\cdot e_{3}=(p+\gamma)e_{2},\quad e_{3}\cdot e_{1}=\gamma e_{1}+e_{2},\quad e_{3}\cdot e_{2}=-e_{1}+\gamma e_{2},\\[5.69054pt] e_{3}\cdot e_{3}=\lambda e_{1}+\mu e_{2}+(p+\gamma)e_{3},\quad e_{3}\cdot e_{4}=(p+q+\gamma)e_{4},\quad e_{4}\cdot e_{3}=(p+\gamma)e_{4}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{3}\cdot e_{1}=-pe_{1}+e_{2},&e_{3}\cdot e_{2}=-e_{1}-pe_{2},&e_{3}\cdot e_{3}=\lambda e_{1}+\gamma e_{2}&e_{3}\cdot e_{4}=qe_{4}.\end{array}$
  27. (27)
    The Lie superalgebra $D^{11}_{pq}$ with $0<|p|\leq|q|\leq 1$: A Novikov structure is given by:
    $\begin{array}[]{l}e_{1}\cdot e_{2}=e_{2},\quad e_{1}\cdot e_{3}=pe_{3},\quad e_{1}\cdot e_{4}=qe_{4}.\end{array}$
    A Balinsky-Novikov structure is given by:
    $\begin{array}[]{llll}e_{1}\cdot e_{2}=e_{2},&e_{1}\cdot e_{3}=pe_{3},&e_{1}\cdot e_{4}=qe_{4}.\end{array}$
  28. (28)
    The Lie superalgebra $D^{12}$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
    $\begin{array}[]{l}e_{1}\cdot e_{2}=e_{2},\quad e_{1}\cdot e_{4}=(1+\gamma)e_{3},\quad e_{4}\cdot e_{1}=\gamma e_{3}.\end{array}$
    A Balinsky-Novikov structure is given by:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=-2e_{1},&e_{1}\cdot e_{3}=-e_{3},&e_{1}\cdot e_{4}=e_{3}-e_{4},&e_{2}\cdot e_{1}=-2e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-2e_{3},&e_{4}\cdot e_{1}=-2e_{4}.\end{array}$
  29. (29)
    The Lie superalgebra $D^{13}_{p}$ with $p\not=0$: A Novikov structure is given by:
    $\begin{array}[]{l}e_{1}\cdot e_{2}=pe_{2},\quad e_{1}\cdot e_{3}=e_{3},\quad e_{1}\cdot e_{4}=e_{3}+e_{4}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=2(\gamma-1)e_{1},&e_{1}\cdot e_{2}=(p-1+\gamma)e_{2},&e_{1}\cdot e_{3}=\gamma e_{3},&e_{1}\cdot e_{4}=e_{3}+\gamma e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2(\gamma-1)e_{2},&e_{3}\cdot e_{1}=2(\gamma-1)e_{3},&e_{4}\cdot e_{1}=2(\gamma-1)e_{4}.\end{array}$
  30. (30)
    The Lie superalgebra $D^{14}_{pq}$ with $p\not=0,q\geq 0$: A Novikov structure is given by:
    $\begin{array}[]{l}e_{1}\cdot e_{2}=pe_{2},\quad e_{1}\cdot e_{3}=qe_{3}-e_{4},\quad e_{1}\cdot e_{4}=e_{3}+qe_{4}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=-2qe_{1},&e_{1}\cdot e_{2}=(p-q)e_{2},&e_{1}\cdot e_{3}=-e_{4},&e_{1}\cdot e_{4}=e_{3},\\[5.69054pt] e_{2}\cdot e_{1}=-2qe_{2},&e_{3}\cdot e_{1}=-2qe_{3},&e_{4}\cdot e_{1}=-2qe_{4}.\end{array}$
  31. (31)
    The Lie superalgebra $D^{15}$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
    $\begin{array}[]{l}e_{1}\cdot e_{3}=e_{2},\quad e_{1}\cdot e_{4}=\gamma e_{2}+e_{3},\quad e_{4}\cdot e_{1}=\gamma e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=2\gamma e_{1},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=e_{2}+\gamma e_{3},&e_{1}\cdot e_{4}=e_{3}+\gamma e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2\gamma e_{2},&e_{3}\cdot e_{1}=2\gamma e_{3},&e_{4}\cdot e_{1}=2\gamma e_{4}.\end{array}$
  32. (32)
    The Lie superalgebra $D^{16}$: A Novikov structure is given by:
    $\begin{array}[]{l}e_{1}\cdot e_{2}=e_{2},\quad e_{1}\cdot e_{3}=e_{2}+e_{3},\quad e_{1}\cdot e_{4}=e_{3}+e_{4}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=2(\gamma-1)e_{1},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=e_{2}+\gamma e_{3},&e_{1}\cdot e_{4}=e_{3}+\gamma e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2(\gamma-1)e_{2},&e_{3}\cdot e_{1}=2(\gamma-1)e_{3},&e_{4}\cdot e_{1}=2(\gamma-1)e_{4}.\end{array}$
  33. (33)
    The Lie superalgebra $(A_{3,1}+A)$: A Novikov structure is given by (where $\lambda,\gamma,\mu\in\mathbb{R}$):
    $\begin{array}[]{l}e_{2}\cdot e_{2}=\gamma e_{1},\quad e_{2}\cdot e_{3}=(1+\lambda)e_{1},\quad e_{3}\cdot e_{2}=\lambda e_{1},\quad e_{3}\cdot e_{3}=\mu e_{1},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{1}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda,\mu\in\mathbb{R})$:
    $\begin{array}[]{lllll}e_{2}\cdot e_{2}=\gamma e_{1},&e_{2}\cdot e_{3}=\lambda e_{1},&e_{3}\cdot e_{2}=(\lambda-1)e_{1},&e_{3}\cdot e_{3}=\mu e_{1}&e_{4}\cdot e_{4}=e_{1}.\end{array}$
  34. (34)
    The Lie superalgebra $(D^{3}_{p,-1/2})$ with $p\not=0$: A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{l}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2}+\gamma e_{3},\quad e_{1}\cdot e_{2}=\frac{1}{2}e_{2},\quad e_{1}\cdot e_{3}=\frac{1}{2}(2p-1)e_{3},\quad e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\quad e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{lllll}e_{1}\cdot e_{1}=\gamma e_{2}+\lambda e_{3},&e_{1}\cdot e_{2}=e_{2},&e_{1}\cdot e_{3}=pe_{3},&e_{1}\cdot e_{4}=\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
  35. (35)
    The Lie superalgebra $(D^{2}_{-1/2})^{1}$: A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{l}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2}+\gamma e_{3},\quad e_{1}\cdot e_{2}=\frac{1}{2}e_{2},\quad e_{1}\cdot e_{3}=e_{2}+\frac{1}{2}e_{3},\quad e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\quad e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{lllll}e_{1}\cdot e_{1}=\gamma e_{2}+\lambda e_{3},&e_{1}\cdot e_{2}=e_{2},&e_{1}\cdot e_{3}=e_{2}+e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
  36. (36)
    The Lie superalgebra $(D^{2}_{-1/2})^{2}$: A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{l}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2}+\gamma e_{3},\quad e_{1}\cdot e_{2}=\frac{1}{2}e_{2},\quad e_{1}\cdot e_{3}=-e_{2}+\frac{1}{2}e_{3},\quad e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\quad e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{lllll}e_{1}\cdot e_{1}=\gamma e_{2}+\lambda e_{3},&e_{1}\cdot e_{2}=e_{2},&e_{1}\cdot e_{3}=-e_{2}+e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
  37. (37)
    The Lie superalgebra $(A_{1,1}+3A)^{1}$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in\mathbb{R}$):
    $\begin{array}[]{l}e_{2}\cdot e_{2}=\frac{1}{2}e_{1},\quad e_{2}\cdot e_{3}=\lambda e_{1},\quad e_{2}\cdot e_{4}=\gamma e_{1},\quad e_{3}\cdot e_{2}=-\lambda e_{1},\quad e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\\[5.69054pt] e_{3}\cdot e_{4}=\mu e_{1},\quad e_{4}\cdot e_{2}=-\gamma e_{1},\quad e_{4}\cdot e_{3}=-\mu e_{1},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{1}.\end{array}$
    A Balinsky-Novikov structure is given by:
    $\begin{array}[]{llll}e_{2}\cdot e_{2}=e_{1},&e_{3}\cdot e_{3}=e_{1},&e_{4}\cdot e_{4}=e_{1}.\end{array}$
  38. (38)
    The Lie superalgebra $(A_{1,1}+3A)^{2}$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in\mathbb{R}$):
    $\begin{array}[]{l}e_{2}\cdot e_{2}=\frac{1}{2}e_{1},\quad e_{2}\cdot e_{3}=\lambda e_{1},\quad e_{2}\cdot e_{4}=\gamma e_{1},\quad e_{3}\cdot e_{2}=-\lambda e_{1},\quad e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\\[5.69054pt] e_{3}\cdot e_{4}=\mu e_{1},\quad e_{4}\cdot e_{2}=-\gamma e_{1},\quad e_{4}\cdot e_{3}=-\mu e_{1},\quad e_{4}\cdot e_{4}=-\frac{1}{2}e_{1}.\end{array}$
    A Balinsky-Novikov structure is given by:
    $\begin{array}[]{llll}e_{2}\cdot e_{2}=e_{1},&e_{3}\cdot e_{3}=e_{1},&e_{4}\cdot e_{4}=-e_{1}.\end{array}$
6. The list of four-dimensional real Lie superalgebras
The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3},e_{4}$ Symplectic structure $\omega$ $p(\omega)$
$D^{5}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{4},\\[2.84526pt] [e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
$D^{6}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{4},\\[2.84526pt] [e_{2},e_{3}]&=&-e_{4},\\[2.84526pt] [e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
$\begin{array}[]{c}D^{7}_{pq},\\[2.84526pt] pq\not=0,\\[2.84526pt] p\geq q\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ None
$\begin{array}[]{c}D^{7}_{-1q},\\[2.84526pt] q\leq-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&-e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ $\begin{array}[]{l}e_{4}^{*}\wedge e_{1}^{*}+e_{3}^{*}\wedge e_{2}^{*}\end{array}$ ${\bar{1}}$
$\begin{array}[]{c}D^{7}_{pp},\\[5.69054pt] p=-q\\[5.69054pt] q\not=0,-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&pe_{3}\\[2.84526pt] [e_{1},e_{4}]&=&-pe_{4}\end{array}$ $\begin{array}[]{l}e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\end{array}$ $\bar{0}$
$\begin{array}[]{c}D^{7}_{pq},\\[2.84526pt] p=-q,\\[2.84526pt] q=-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&-e_{4}\end{array}$ $\begin{array}[]{l}e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\begin{array}[]{l}\bar{0}\\[2.84526pt] \bar{1}\end{array}$
$\begin{array}[]{c}D^{8}_{p},\\ p\not=0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+pe_{4}\end{array}$ None
$D^{8}_{-1}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&-e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}-e_{4}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
$\begin{array}[]{c}D^{9}_{pq}\\[2.84526pt] q>0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&pe_{3}-qe_{4},\\[2.84526pt] [e_{1},e_{2}]&=&e_{2},\\[5.69054pt] [e_{1},e_{4}]&=&qe_{3}+pe_{4}\end{array}$ None
$D^{10}_{q}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&(q+1)e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4},\\[2.84526pt] [e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}(1+q)e_{1}^{*}\wedge e_{3}^{*}+e_{2}^{*}\wedge e_{4}^{*},\\[5.69054pt] \text{where }q\not=-1\end{array}$ $\bar{1}$
Table 1. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=2|2$
The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3},e_{4}$ Symplectic structure $p(\omega)$
[…]
$\begin{array}[]{c}(D^{7}_{1-p,p})\\[2.84526pt] p\leq\frac{1}{2}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3}\\[2.84526pt] [e_{1},e_{4}]&=&(1-p)e_{4},\\[2.84526pt] [e_{3},e_{4}]&=&e_{2}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}-e_{3}^{*}\wedge e_{4}^{*}$ $\bar{0}$
$(D^{8}_{1/2})$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+\frac{1}{2}e_{4},\\[2.84526pt] [e_{4},e_{4}]&=&e_{2}\end{array}$ None --
$\begin{array}[]{c}(D^{9}_{1/2,p})\\[2.84526pt] p>0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{3},e_{3}]&=&e_{2}\\[2.84526pt] [e_{1},e_{4}]&=&pe_{3}+\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{1},e_{3}]&=&\frac{1}{2}e_{3}-pe_{4}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}-\frac{1}{2}e_{3}^{*}\wedge e_{3}^{*}-\frac{1}{2}e_{4}^{*}\wedge e_{4}^{*}$ ${\bar{0}}$
$(D^{10}_{0})^{1}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{3}\\[2.84526pt] [e_{2},e_{4}]&=&e_{3},&[e_{4},e_{4}]&=&e_{1}\\[2.84526pt] [e_{3},e_{4}]&=&-\frac{1}{2}e_{2}\end{array}$ $\begin{array}[]{l}2e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] e_{1}^{*}\wedge e_{3}^{*}+e_{2}^{*}\wedge e_{4}^{*}\end{array}$ $\begin{array}[]{l}{\bar{0}}\\[2.84526pt] {\bar{1}}\end{array}$
$(D^{10}_{0})^{2}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{3}\\[2.84526pt] [e_{2},e_{4}]&=&e_{3},&[e_{4},e_{4}]&=&-e_{1}\\[2.84526pt] [e_{3},e_{4}]&=&\frac{1}{2}e_{2}\end{array}$ $\begin{array}[]{l}2e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] e_{1}^{*}\wedge e_{3}^{*}+e_{2}^{*}\wedge e_{4}^{*}\end{array}$ $\begin{array}[]{l}{\bar{0}}\\[2.84526pt] {\bar{1}}\end{array}$
$(2A_{1,1}+2A)^{1}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None --
$(2A_{1,1}+2A)^{2}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{3},e_{4}]&=&e_{1}\\ \end{array}$ None --
$\begin{array}[]{c}(2A_{1,1}+2A)^{3}_{p}\\ p>0\end{array}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{3},e_{4}]&=&p(e_{1}+e_{2})\\ \end{array}$ For p12: NoneFor $\begin{array}[]{l}\text{For $p\not=\frac{1}{2}$: None}\\[2.84526pt] \text{For $p=\frac{1}{2}:\,e_{2}^{*}\wedge e_{3}^{*}-e_{1}^{*}\wedge e_{4}^{*}$}\end{array}$ $\ \bar{1}\end{array}$
$\begin{array}[]{c}(2A_{1,1}+2A)^{4}_{p}\\ p>0\end{array}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{3},e_{4}]&=&p(e_{1}-e_{2})\\ \end{array}$ None --
$(C^{1}_{1}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{3}\\[2.84526pt] [e_{3},e_{4}]&=&e_{2}\\ \end{array}$ $\frac{1}{2}e_{2}^{*}\wedge e_{1}^{*}+\frac{1}{2}e_{3}^{*}\wedge e_{4}^{*}+\frac{1}{4}e_{4}^{*}\wedge e_{4}^{*}$ $\bar{0}$
$(C^{1}_{1/2}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{3},e_{3}]&=&e_{2}\\ \end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}-\frac{\lambda}{2}e_{3}^{*}\wedge e_{3}^{*}-\frac{\mu}{2}e_{4}^{*}\wedge e_{4}^{*},\\ \text{where }\lambda\mu\not=0\end{array}$ $\bar{0}$
$(C^{2}_{-1}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&-e_{4}\\[2.84526pt] [e_{3},e_{4}]&=&e_{2}\\ \end{array}$ None --
$(C^{3}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{4}]&=&e_{3},&[e_{4},e_{4}]&=&e_{2}\end{array}$ $2e_{1}^{*}\wedge e_{2}^{*}-e_{3}^{*}\wedge e_{4}^{*}$ $\bar{0}$
$(C^{5}_{0}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&-e_{4},&[e_{1},e_{4}]&=&e_{3}\\[2.84526pt] [e_{3},e_{3}]&=&e_{2}&[e_{4},e_{4}]&=&e_{2}\\ \end{array}$ None --
Table 2. Non-trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=2|2$
The LSA Relations in the basis: $e_{1},e_{2},e_{3}\,|\,e_{4}$ Symplectic structure $p(\omega)$
$D^{1}$ $\begin{array}[]{lcrlcl}[e_{2},e_{3}]&=&e_{1},&[e_{2},e_{4}]&=&e_{4}\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{3}^{*}+\mu e_{2}\wedge e_{4}^{*}\\[2.84526pt] +\nu e_{1}^{*}\wedge e_{2}^{*}+\gamma e_{2}\wedge e_{3}^{*}\\[2.84526pt] \text{where }\lambda\mu\not=0\end{array}$ NH
$\begin{array}[]{l}D^{2}_{q}\\[2.84526pt] q\not=-1,0\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{1},&[e_{2},e_{3}]&=&e_{1}+e_{2},\\[2.84526pt] [e_{3},e_{4}]&=&qe_{4}\end{array}$ None --
$\begin{array}[]{l}D^{2}_{-1}\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{1},&[e_{2},e_{3}]&=&e_{1}+e_{2},\\[2.84526pt] [e_{3},e_{4}]&=&-e_{4}\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{3}^{*}+\mu e_{2}^{*}\wedge e_{3}^{*}+\\[5.69054pt] \nu e_{2}^{*}\wedge e_{4}^{*}+\gamma e_{3}^{*}\wedge e_{4}^{*}\\[5.69054pt] \text{where }\lambda\nu\not=0\end{array}$ NH
$\begin{array}[]{l}D^{3}_{pq}\\[2.84526pt] pq\not=0\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&pe_{1}-e_{2},&[e_{2},e_{3}]&=&e_{1}+pe_{2},\\[2.84526pt] [e_{3},e_{4}]&=&qe_{4}\end{array}$ None --
Table 3. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=3|1$
The LSA Relations in the basis: $e_{1}\,|\,e_{2},e_{3},e_{4}$ Symplectic structure $p(\omega)$
$\begin{array}[]{c}D^{11}_{pq}\\ 0<|p|\leq|q|\leq 1\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ None --
$\begin{array}[]{c}D^{11}_{pq}\\ (p,q)=(-1,-1)\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}+\nu e_{1}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\gamma e_{2}^{*}\wedge e_{3}^{*}+\delta e_{2}^{*}\wedge e_{4}^{*},\\[2.84526pt] \text{where }\nu\gamma-\mu\delta\not=0\end{array}$ NH
$\begin{array}[]{c}D^{11}_{pq}\\ 0<|p|\leq 1\\ p=-q\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}+\nu e_{1}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\gamma e_{3}^{*}\wedge e_{4}^{*},\text{ where }\lambda\gamma\not=0\end{array}$ NH
$\begin{array}[]{c}D^{11}_{pq}\\ 0<|p|<1\\ q=-1\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}+\nu e_{1}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\gamma e_{2}^{*}\wedge e_{4}^{*},\text{ where }\mu\gamma\not=0\end{array}$ NH
$D^{12}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{4}]&=&e_{3},\end{array}$ None --
$\begin{array}[]{c}D^{13}_{p}\\[2.84526pt] p\not=0\text{ (generic)}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&pe_{2},&[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ None --
$\begin{array}[]{c}D^{13}_{-1}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&-e_{2},&[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}+\gamma e_{1}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\nu e_{2}^{*}\wedge e_{4}^{*},\text{ where }\mu\nu\not=0\end{array}$ NH
$\begin{array}[]{c}D^{14}_{pq}\\[2.84526pt] p\not=0,q\geq 0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&pe_{2},\\[2.84526pt] [e_{1},e_{3}]&=&qe_{3}-e_{4},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}+e_{3}\end{array}$ None --
$D^{15}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{2},&[e_{1},e_{4}]&=&e_{3},\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}+\nu e_{1}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\delta e_{2}^{*}\wedge e_{4}^{*}-\frac{1}{2}\delta e_{3}^{*}\wedge e_{3}^{*},\\[2.84526pt] \text{where }\delta^{2}(\mu^{2}-2\lambda\nu)\not=0.\end{array}$ NH
$D^{16}$ $\begin{array}[]{lcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ None --
Table 4. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=1|3$
The LSA Relations in the basis: $e_{1},e_{2},e_{3}\,|\,e_{4}$ Symplectic structure $p(\omega)$
$(A_{3,1}+A)$ $\begin{array}[]{lcrlcl}[e_{2},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{1}\end{array}$ None --
$\begin{array}[]{l}(D^{3}_{p,-1/2})\\[2.84526pt] p\not=0\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] \end{array}$ None --
$(D^{3}_{-1/2,-1/2})$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-\frac{1}{2}e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] \end{array}$ $\begin{array}[]{l}\frac{1}{2}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}\\[2.84526pt] +\nu e_{1}^{*}\wedge e_{4}^{*}+\gamma e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\frac{1}{2}\lambda e_{4}^{*}\wedge e_{4}^{*},\text{ where }\gamma\lambda\not=0\end{array}$ NH
$(D^{2}_{-1/2})^{1}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None --
$(D^{2}_{-1/2})^{2}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None --
Table 5. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=3|1$
The LSA Relations in the basis: $e_{1}\,|\,e_{2},e_{3},e_{4}$ Symplectic structure $p(\omega)$
$(A_{1,1}+3A)^{1}$ $\begin{array}[]{lcrlcl}[e_{2},e_{2}]&=&e_{1},&[e_{3},e_{3}]&=&e_{1},\\[2.84526pt] [e_{4},e_{4}]&=&e_{1}\end{array}$ None --
$(A_{1,1}+3A)^{2}$ $\begin{array}[]{lcrlcl}[e_{2},e_{2}]&=&e_{1},&[e_{3},e_{3}]&=&e_{1},\\[2.84526pt] [e_{4},e_{4}]&=&-e_{1}\end{array}$ None --
Table 6. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{h}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=1|3$
Acknowledgment. We would like to acknowledge Chengming Bai for pointing out that Novikov algebras have two superizations. Also, we would like to thank Saïd Benayadi, Isabel Cunha, and Quentin Ehret for reading the manuscript and making pertinent comments.

### v2

@@
-Lagrangian extensions and left symmetric structures on the four-dimensional real Lie superalgebras
+Lagrangian extensions and left-symmetric structures on the four-dimensional real Lie superalgebras
 Abstract.
@@
 Moreover, we investigate left-symmetric structures on these Lie superalgebras.
-Furthermore, except for two of them, they are all Novikov superalgebras.
+Furthermore, except for two of them, we show that they are all Novikov superalgebras.
 Key words and phrases:
-Quasi-Frobenius Lie (super)algebra, Lagrangian extensions, Left symmetric structures, Novikov structures, Balinsky-Novikov structures.
+Quasi-Frobenius Lie (super)algebra, Lagrangian extensions, Left-symmetric structures, Novikov structures, Balinsky-Novikov structures.
 2020 Mathematics Subject Classification:
 17B05; 17B55; 17D25
-SB is supported by the grant NYUAD-065.
 1. Introduction
@@
 The notion of $T^{*}$-extension of Lie algebras was initially introduced by Bordemann [Bor].
-A short explanation can be understood as the following.
+A short explanation is as follows.
 Given a Lie algebra $\mathfrak{h}$ equipped with an $\mathfrak{h}^{*}$-valued bilinear map $\alpha$, a $T^{*}$-extension of $\mathfrak{h}$ is the Lie algebra structure on the vector space $\mathfrak{g}:=\mathfrak{h}\oplus\mathfrak{h}^{*}$ constructed by means of the form $\alpha$ that satisfies a certain cohomological requirement.
-The Lie algebra $\mathfrak{g}$ naturally admits an invariant non-degenerate symmetric bilinear form.
+The Lie algebra $\mathfrak{g}$ naturally admits a non-degenerate invariant symmetric bilinear form.
 Conversely, Bordemann showed in [Bor] that every finite-dimensional nilpotent Lie algebra $\mathfrak{h}$ with an invariant symmetric non-degenerate bilinear form “is” a suitable $T^{*}$-extension.
@@
 A superization of this construction was given in [BBB].
-In particular, it was shown that every solvable Lie superalgebra with an even symmetric non-degenerate bilinear form is either isomorphic to a $T^{*}$-extension of a certain Lie superalgebra or to an ideal of codimension one of a certain $T^{*}$-extension.
+In particular, it was shown that every solvable Lie superalgebra with an even non-degenerate symmetric bilinear form is either isomorphic to a $T^{*}$-extension of a certain Lie superalgebra or to an ideal of codimension one of a certain $T^{*}$-extension.
 The notion of $T^{*}$-extension was also studied by Baues and Cortés in [BC] in the context of Lie algebras admitting a flat connection, and called Lagrangian extensions since the Lie sub-algebra $\mathfrak{h}^{*}$ is a Lagrangian ideal in $\mathfrak{h}\oplus\mathfrak{h}^{*}$.
@@
 Lie algebras with a non-degenerate anti-symmetric form satisfying the 2-cocycle conditions are called quasi-Frobenius, and are studied intensively in the literature, see, e.g., [E, O].
-Lie algebras possessing both an invariant non-degenerate symmetric bilinear form and a non-degenerate 2-cocycle of its scalar cohomology, were studied in [BBM] and were called quadratic symplectic Lie algebras.
+Lie algebras possessing both a non-degenerate invariant symmetric bilinear form and a non-degenerate 2-cocycle of its scalar cohomology, were studied in [BBM] and were called quadratic symplectic Lie algebras.
 If the ground field is algebraically closed, the authors showed that every quadratic symplectic Lie algebra is a $T^{*}$-extension of a Lie algebra which admits an invertible derivation.
@@
 These Lie groups are nilpotent, and their geometry is very rich, see [BBM, MR] for more details.
-The notion of Lagrangian extension was superized in [BM].
-The construction is valid for arbitrary Lie superalgebras defined over a field of characteristic not two provided the Lie superalgebra admit a torsion-free flat connection.
+The notion of Lagrangian extension was superized in [BM] (see also [BBE] where the ground field is of characteristic 2).
+The construction is valid for arbitrary Lie superalgebras defined over a field of arbitrary characteristic provided the Lie superalgebra admit a torsion-free flat connection.
 Moreover, they have observed that there are two ways to perform this extension: they consider either $\mathfrak{g}:=\mathfrak{h}\oplus\mathfrak{h}^{*}$ or $\mathfrak{g}:=\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*})$, where $\Pi$ is the change of parity functor.
@@
 One of the main objectives of this paper is to complete the list of examples among the 4-dimensional Lie superalgebras and explain how they can be obtained as Lagrangian extensions, see Section 4.
-According to the Backhouse classification, this investigation will classify all four-dimensional real Lie superalgebras that admit (or don’t admit) Lagrangian extensions.
+According to the Backhouse classification, this investigation will classify all four-dimensional real Lie superalgebras that admit (or do not admit) Lagrangian extensions.
 1.2. Left-symmetric structures
-Left-symmetric algebras (also known as Vinberg-Koszul algebras) play an important role in various fields of mathematics and mathematical physics; for instance, rooted tree algebras, operad theory, vertex algebras, convex homogeneous cones, affine manifolds, left-invariant affine structures on Lie groups, to name just a few.
+Left-symmetric algebras (also known in the literature as Vinberg-Koszul algebras) play an important role in various fields of mathematics and mathematical physics; for instance, rooted tree algebras, operad theory, vertex algebras, convex homogeneous cones, affine manifolds, left-invariant affine structures on Lie groups, to name just a few.
 For a thorough review, see [B3].
-Over the real and complex numbers left-symmetric algebras are closely related to the theory of left-invariant affine structures over Lie groups, see for example [M, GS] (and references therein).
-The investigation and classification of left-symmetric structure on Lie (super)algebras is an important subject by its own, as initiated in [S]. The problem, however, seems to be difficult to handle.
+Over the real and complex numbers left-symmetric algebras are closely related to the theory of left-invariant affine structures over Lie groups; see, for example, [M, GS] (and references therein).
+The investigation and classification of left-symmetric structure on Lie (super)algebras is an important subject on its own, as initiated in [S]. The problem, however, seems to be difficult to handle.
 For reductive Lie algebras, Baues [Ba] proved that $\mathfrak{gl}(n)$ is the only reductive Lie algebra with one-dimensional center and a simple semisimple ideal which admits left- symmetric algebras over an algebraically closed ground field (see also [B1]).
-Morover, all left-symmetric algebras on $\mathfrak{gl}(n)$ were classified.
+Moreover, all left-symmetric algebras on $\mathfrak{gl}(n)$ were classified.
 Simple Lie algebras over complex numbers do not admit left-symmetric structures, see [H].
 The existence of such structures is not always guaranteed, even for nilpotent Lie algebras.
-There are filiform nilpotent Lie groups of dimension $10\leq n\leq 13$ which do not admit any left-invariant affine structure, see [B, B3, BG].
+There are filiform nilpotent Lie groups of dimension $10\leq n\leq 13$ that do not admit any left-invariant affine structure, see [B, B3, BG].
 Any filiform nilpotent Lie group of dimension $n\leq 9$ admits a left-invariant affine structure.
 Nevertheless, left-symmetric algebra structures on both finite-dimensional and infinite-dimensional Lie algebras have been studied extensively, see [Bai, B2, YZ] as well as [CL, KCB, TB], respectively.
-Left symmetric structures on the Virasoro superalgebra were studied in [KB].
+Left-symmetric structures on the Virasoro superalgebra were studied in [KB].
 In the super setting, it was proved in [DZ] that $\mathfrak{sl}(m+1|m)$ admits a left-symmetric superalgebra, and a full classification is offered in the case of $\mathfrak{sl}(2|1)$.
-Moreover, it was shown that $\mathfrak{sl}(m|1)$ does not admit left-symmetric superalgebras for $m\geq 3$.
+Moreover, it was shown that $\mathfrak{sl}(m|1)$ does not admit a left-symmetric structure for $m\geq 3$.
+Over a field of characteristic 2, all left-symmetric structures on a 2-dimensional superspace have been classified in [BBE].
 There is also the notion of Novikov algebra. It was introduced in connection with the Poisson brackets of hydrodynamic type, see [BN].
@@
 Furthermore, we show that these structures are Novikov, except for the two Lie superalgebras $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$.
-In addition, all these Lie superalgebras admit a Balinsky Novikov structure.
+In addition, all these Lie superalgebras admit a Balinsky-Novikov structure.
 Classifying all left-symmetric structures on each of these Lie superalgebras remains an open problem.
@@
 The purpose of this section is to review the main definitions and theorems that will be used in this paper.
-Throughout the text $\mathbb{K}$ is a field of characteristic $p\not=2$.
-A Lie superalgebra is a $\mathbf{Z}_{2}$-graded space $\mathfrak{g}=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{\bar{1}}$ over a field of characteristic $p\not=2$ together with a bilinear binary operator $[\cdot,\cdot]:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$, called the Lie superbracket that has the following properties: (for any homogeneous $x,y,z\in\mathfrak{g}$)
-  1. (i)
-    Super anti-commutativity: $[x,y]=-(-1)^{|x||y|}[y,x]$;
-  2. (ii)
-    Super Jacobi identity: $[x,[y,z]]=[[x,y],z]+(-1)^{|x||y|}[y,[x,z]]$ where $|x|={\bar{0}}$ if $x\in\mathfrak{g}_{\bar{0}}$ and $|x|={\bar{1}}$ if $x\in\mathfrak{g}_{\bar{1}}$;
-  3. (iii)
-    If $p=3$, one must add $[x,[x,x]]=0$ if $x\in\mathfrak{g}_{\bar{1}}$.
-There is also the notion of Lie superalgebra in characteristic $2$ but we are not studying them here.
-Let us introduce the notion of bilinear forms over a superspace. For more details, see [LSoS].
-Let $V$ and $W$ be two superspaces defined over an arbitrary field ${\mathbb{K}}$.
-Let $\omega\in\text{Bil}(V,W)$ be a homogeneous bilinear form.
-The Gram matrix $B=(B_{ij})$ associated to $\omega$ is given by the formula,
-(1) $B_{ij}=(-1)^{|\omega||v_{i}|}\omega(v_{i},v_{j})\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for the basis vectors $v_{i}\in V$.}$
-This definition can be extended by linearity to non-homogeneous forms.
-Moreover, it allows us to identify a bilinear form $B(V,W)$ with an element of $\mathrm{Hom}(V,W^{*})$.
-Consider the upsetting of bilinear forms $u\colon\mathrm{Bil}(V,W)\rightarrow\mathrm{Bil}(W,V)$ given by the formula
-(2) $u(\omega)(u,v)=(-1)^{|v||u|}\omega(v,u)\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for any homogeneous $v\in V$ and $u\in W$.}$
-In terms of the Gram matrix $B$ of $\omega$: the form $\omega$ is symmetric if and only if
-(3) $u(B)=B,\;\text{ where $u(B)=\left(\begin{array}[]{cc}R^{t}&(-1)^{|\omega|}T^{t}\\[5.69054pt] (-1)^{|\omega|}S^{t}&-U^{t}\end{array}\right),$ for $B=\left(\begin{array}[]{cc}R&S\\[5.69054pt] T&U\end{array}\right)$.}$
-Similarly, anti-symmetry of $\omega$ means that $u(B)=-B$.
-An even (resp. odd) non-degenerate bilinear form is called orthosymplectic (resp. periplectic).
-Proposition 2.1 (Superdimension constraints).
-If $V=V_{\bar{0}}\oplus V_{\bar{1}}$ is a finite-dimensional superspace such that $V_{\bar{1}}\not={0}$, equipped with an anti-symmetric non-degenerate bilinear form $\omega$, then:
-  1. (i)
-    If $p(\omega)=\bar{0}$, then $\mathrm{dim}(V_{\bar{0}})$ is even.
-  2. (ii)
-    If $p(\omega)=\bar{1}$, then $\mathrm{dim}(V_{\bar{0}})=\mathrm{dim}(V_{\bar{1}})$.
-Proof.
-See, [BM]
-We will denote by $\Pi$ the functor that assigns to $V$ the superspace $\Pi(V)$ defined as the other copy of $V$ with the opposite parity of its homogeneous components. We have:
+Throughout the text, $\mathbb{K}$ is a field of characteristic $p\not=2$.
+Let $V=V_{\bar{0}}\oplus V_{\bar{1}}$ be a superspace defined over $\mathbb{K}$}.
+The parity of a homogeneous element $v\in V_{\bar{i}}$ is denoted by $|v|:=\bar{i}$.
+The element $v$ is called even if $v\in V_{\bar{0}}$ and odd if $v\in V_{\bar{1}}$.
+The superdimension of $V$ is $\mathrm{sdim}(V)=a+b\epsilon$, where $\epsilon^{2}=1$, and $a=\mathrm{dim}(V_{\bar{0}})$, $b=\mathrm{dim}(V_{\bar{1}})$.
+Usually, $\mathrm{sdim}(V)$ is shorthanded as $a\mid b$.
+We will denote by $\Pi$ the functor that assigns to $V$ the superspace $\Pi(V)$ defined as the other copy of $V$ with the opposite parity of its homogeneous components. We have
 $(\Pi(V))_{\bar{0}}=V_{\bar{1}},\quad(\Pi(V))_{\bar{1}}=V_{\bar{0}}.$
@@
 We will identify $V$ and $\Pi(\Pi(V))$, naturally.
-Let $\omega\in\mathrm{Bil}(\mathfrak{g},\mathfrak{g})$ be an anti-symmetric bilinear form on a Lie superalgebra $\mathfrak{g}$. The form $\omega$ is called closed if it satisfies:
+Let us introduce the notion of bilinear forms over a superspace. For more details, see [LSoS].
+Let $V$ and $W$ be two superspaces defined over an arbitrary field $\mathbb{K}$}.
+Let $\mathcal{B}\in\text{Bil}(V,W)$ be a homogeneous bilinear form.
+The Gram matrix $B=(B_{ij})$ associated to $\mathcal{B}$} is given by the formula
+$B_{ij}=(-1)^{|{\mathcal{B}}||v_{i}|}{\mathcal{B}}(v_{i},w_{j})\text{\penalty 10000\ \penalty 10000\ for the basis vectors $v_{i}\in V$ and $w_{j}\in W$.}$
+This definition can be extended by linearity to non-homogeneous forms.
+Moreover, it allows us to identify a bilinear form $B(V,W)$ with an element of $\mathrm{Hom}(V,W^{*})$.
+Consider the upsetting of bilinear forms $u\colon\mathrm{Bil}(V,W)\rightarrow\mathrm{Bil}(W,V)$ given by the formula
+[…]
+From now on, let us assume that $V=W$.
+In terms of the Gram matrix $B$ of $\mathcal{B}$}: the form $\mathcal{B}$} is symmetric if and only if
+$u(B)=B,\;\text{ where $u(B)=\left(\begin{array}[]{cc}R^{t}&(-1)^{|\omega|}T^{t}\\[5.69054pt] (-1)^{|\omega|}S^{t}&-U^{t}\end{array}\right),$ for $B=\left(\begin{array}[]{cc}R&S\\[5.69054pt] T&U\end{array}\right)$.}$
+Similarly, anti-symmetry of $\mathcal{B}$} means that u(B)=Bu(B)=-B.
+Throughout the text, we denote by $\omega$ an anti-symmetric bilinear form.
+An even (resp. odd) non-degenerate anti-symmetric bilinear form is called orthosymplectic (resp. periplectic).
+Proposition 2.1 (Superdimension constraints).
+If $V=V_{\bar{0}}\oplus V_{\bar{1}}$ is a finite-dimensional superspace such that $V_{\bar{1}}\not={0}$, equipped with a non-degenerate anti-symmetric bilinear form $\omega$, then:
+  1. (i)
+    If $|\omega|=\bar{0}$, then $\mathrm{dim}(V_{\bar{0}})$ is even.
+  2. (ii)
+    If $|\omega|=\bar{1}$, then $\mathrm{dim}(V_{\bar{0}})=\mathrm{dim}(V_{\bar{1}})$.
+Proof.
+See, [BM]
+A Lie superalgebra is a superspace $\mathfrak{g}=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{\bar{1}}$ over a field of characteristic $p\not=2$ together with a bilinear binary operator $\cdot,\cdot]:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$, called the Lie bracket that has the following properties: (for all homogeneous $x,y,z\in\mathfrak{g}$)
+  1. (i)
+    Super anti-commutativity: $x,y]=-(-1)^{|x||y|}[y,x]$;
+  2. (ii)
+    Super Jacobi identity: $x,[y,z]]=[[x,y],z]+(-1)^{|x||y|}[y,[x,z]]$ where $|x|={\bar{0}}$ if $x\in\mathfrak{g}_{\bar{0}}$ and $|x|={\bar{1}}$ if $x\in\mathfrak{g}_{\bar{1}}$;
+  3. (iii)
+    If $p=3$, one must add $[x,[x,x]]=0$ if $x\in\mathfrak{g}_{\bar{1}}$.
+There is also the notion of Lie superalgebra in characteristic $2$ but we are not studying them here.
+Let $\omega\in\mathrm{Bil}(\mathfrak{g},\mathfrak{g})$ be an anti-symmetric bilinear form on a Lie superalgebra $\mathfrak{g}$. The form $\omega$ is called closed if it satisfies
 $(-1)^{|x||z|}\omega(x,[y,z])+(-1)^{|z||y|}\omega(z,[x,y])+(-1)^{|y||x|}\omega(y,[z,x])=0,$
-for all $x,y,z\in\mathfrak{g}_{\bar{i}}$. This means that $\omega$ is 2-cocyle with scalar values.
-A Lie superalgebra $\mathfrak{g}$ is called quasi-Frobenius if it is equipped with a closed anti-symmetric non-degenerate bilinear form $\omega$.
+for all homogeneous $x,y,z\in\mathfrak{g}_{\bar{i}}$. This means that $\omega$ is 2-cocycle with scalar values.
+A Lie superalgebra $\mathfrak{g}$ is called quasi-Frobenius if it is equipped with a closed non-degenerate anti-symmetric bilinear form $\omega$.
 If $\omega$ is even (resp.
@@
 We denote such a Lie superalgebra by $(\mathfrak{g},\omega)$.
-Recently, these Lie superalgebras were studied in [BE, BEM, BM] in the context of Lagrangian and double extensions.
+Recently, these Lie superalgebras were studied in [BBE, BE, BEM, BM] in the context of Lagrangian and double extensions.
 A Lie superalgebra is called Frobenius if $\omega$ is exact, which means that there is $f\in\mathfrak{g}^{*}$ such that $\omega(x,y)=f([x,y])$ for all $x,y\in\mathfrak{g}.$
@@
 We will introduce the notion of torsion and curvature of a connection.
-Let $\nabla:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ be a connection on $\mathfrak{g}$. For all $x,y,z\in\mathfrak{g}_{\bar{i}}$, the torsion $T^{\nabla}$ of $\nabla$ is defined as
+Let $\nabla:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ be a connection on $\mathfrak{g}$.
+For any homogeneous $x,y,z\in\mathfrak{g}_{\bar{i}}$, the torsion $T^{\nabla}$ of $\nabla$ is defined as
 (1) $T(x,y):=\nabla_{x}y-(-1)^{|x||y|}\nabla_{y}x-[x,y].$
@@
 In this case, the connection $\nabla$ will be called flat.
-2.2. Left-Symmetric Lie (super)algebras
+2.2. Left-symmetric (super)algebras
 Let us review a few notions regarding left- symmetric structures.
@@
 These superalgebras are Lie-admissible.
-The supercommutator $[x,y]:=x\cdot y-(-1)^{|x||y|}y\cdot x$ defines then a Lie superalgebra structure on $L$ and we denote it by ${\mathfrak{g}}(L)$.
+The supercommutator $x,y]:=x\cdot y-(-1)^{|x||y|}y\cdot x$ defines then a Lie superalgebra structure on $L$ and we denote it by ${\mathfrak{g}}(L)$.
 The resulting Lie superalgebra is called the associated Lie superalgebra of the LSSA.
@@
 Associative superalgebras are all LSSAs.
-In [H], Helmstetter showed that it is necessary to have $[L,L]\subsetneq L$ if $L$ is a left symmetric algebra defined over the field of complex numbers.
+In [H], Helmstetter showed that it is necessary to have $L,L]\subsetneq L$ if $L$ is a left-symmetric algebra defined over the field of complex numbers.
 Over a field of characteristic $p$, this is not true anymore.
-Burde showed in [B2] that classical Lie algebras over a field $p>3$ admit an LSS only in case $p$ divides $\mathrm{dim}(L)$.
+Burde showed in [B2] that classical Lie algebras over a field $p>3$ admit an LSA only in case $p$ divides $\mathrm{dim}(L)$.
 In the super setting and over a field of prime characteristic, there has not been much research done so far.
@@
 $x\cdot y:=\nabla_{x}(y)\text{ for all }x,y\in{\mathfrak{g}}.$
-Additionally, every quasi-Frobenius Lie superalgebra $({\mathfrak{g}},\omega)$ admits a left symmetric structure given by
+Additionally, every quasi-Frobenius Lie superalgebra $({\mathfrak{g}},\omega)$ admits a left-symmetric structure given by
 $\omega(x\cdot y,z)=(-1)^{|x||y|}\omega(y,[x,z])\text{ for all }x,y\in{\mathfrak{g}}_{\bar{i}},\text{ and }z\in{\mathfrak{g}}.$
-The converse is not always true. There will be several examples that illustrate that in Section 5.
-Some subclasses of LSSAs are important as well.
+The converse is not always true. Section 5 provides several examples illustrating this.
+Certain subclasses of LSSAs are also of particular interest.
 An LSSA with supercommutative right multiplication is called a Novikov superalgebra (cf.
 [BN, Xu]).
-Explicitly, a Novikov superalgebra $(L,\cdot)$ is an LSSA with an additional condition:
+Explicitly, a Novikov superalgebra $(L,\cdot)$ is an LSSA with an additional condition
 $(z\cdot x)\cdot y=(-1)^{|x||y|}(z\cdot y)\cdot x,\;\text{ for all }x,y\in L_{\bar{i}},\text{ and }z\in L.$
@@
   1. (1)
-    Left symmetry and commutativity:
+    Left-symmetry and commutativity:
     $(x\cdot y)\cdot z-x\cdot(y\cdot z)=(y\cdot x)\cdot z-y\cdot(x\cdot z),\qquad(x\cdot y)\cdot z=(x\cdot z)\cdot y,$
-    for any homogeneous elements
+    for all homogeneous elements
     $x,y,z\in L_{\bar{0}};\quad x,y\in L_{\bar{0}},z\in L_{\bar{1}};\quad x,z\in L_{\bar{0}},y\in L_{\bar{1}};\quad y,z\in L_{\bar{0}},x\in L_{\bar{1}};$
@@
     Compatibility conditions:
-    $\begin{array}[]{ll}(x\cdot y)\cdot z=(x\cdot z)\cdot y&\text{ for all }x,y\in L_{\bar{1}}\text{ and }z\in L_{\bar{0}};\\[5.69054pt] x\cdot(y\cdot z)=(x\cdot y)\cdot z+y\cdot(x\cdot z)&\text{ for all }x\in L_{\bar{0}}\text{ and }y,z\in L_{\bar{1}};\\[5.69054pt] x\cdot(y\cdot z)=(x\cdot y)\cdot z+(x\cdot z)\cdot y&\text{ for all }x,y,z\in L_{\bar{1}}.\\[5.69054pt] \end{array}$
+[…]
 Balinsky Novikov superalgebras are also admissible Lie superalgebras. The Lie structure is given as follows:
-$\begin{array}[]{ll}[x,y]:=x\cdot y-y\cdot x,&\text{ for all }x,y\in L_{\bar{0}};\\[5.69054pt] [x,y]:=x\cdot y-\frac{1}{2}y\cdot x=-[y,x],&\text{ for all }x\in L_{\bar{0}},\text{ and }y\in L_{\bar{1}}\\[5.69054pt] [x,y]:=x\cdot y&\text{ for all }x,y\in L_{\bar{1}}.\end{array}$
+[…]
 2.3. The 4-dimensional real Lie superalgebras
@@
   1. (1)
-    Table 1 having the trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=2|2$;
+    Table 1 having the trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=2|2$;
   2. (2)
-    Table 2 having the non-trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=2|2$;
+    Table 2 having the non-trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=2|2$;
   3. (3)
-    Table 3 having the trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=3|1$;
+    Table 3 having the trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=3|1$;
   4. (4)
-    Table 4 having the trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=1|3$;
+    Table 4 having the trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=1|3$;
   5. (5)
-    Table 5 having the non-trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=3|1$;
+    Table 5 having the non-trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=3|1$;
   6. (6)
-    Table 6 having the non-trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=1|3$.
+    Table 6 having the non-trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=1|3$.
 In this paper, we present only the forms as in Table 1-6 which are suitable for Lagrangian extensions, as opposed to [BM], where the forms are given in the form of parametric families.
 The classification of these structures up to a symplectomorphism is still an open problem.
-We also correct a claim in [BM] stating that only non-homogenous forms exist for the Lie superalgebras $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$.
+We also correct a claim in [BM] stating that only non-homogeneous forms exist for the Lie superalgebras $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$.
 It turns out that these two superalgebras both admit an even and an odd non-degenerate closed form, and both of them can be obtained as Lagrangian extensions.
-3. The notion of $T^{*}$-extensions and $\Pi T^{*}$- (Lagrangian extensions)
+We only consider indecomposable Lie superalgebras.
+3. The notion of $T^{*}$- extensions and $\Pi T^{*}$- extensions
 We will introduce the notion of $T^{*}$-extension and $\Pi T^{*}$-extension of Lie superalgebras (Lagrangian extensions).
-Originally, the notion of $T^{*}$-extension for Lie algebras was first introduced by Bordemann in [Bor].
-Recently, it was studied in [BC] in the context of quasi-Frobenius Lie algebras, and then in [BM] in the context of quasi-Frobenius Lie superalgebras.
+The notion of a $T^{*}$-extension for Lie algebras was originally introduced by Bordemann in [Bor].
+Recently, it was studied in [BC] in the context of quasi-Frobenius Lie algebras, and then in [BBE, BM] in the context of quasi-Frobenius Lie superalgebras.
 3.1. Flat connections and representations
-Let $({\mathfrak{h}},\nabla)$ be a Lie superalgebra endowed with a flat connection $\nabla$. Since the connection $\nabla$ is flat, the map
+Let $({\mathfrak{h}},\nabla)$ be a Lie superalgebra endowed with a flat connection $\nabla$. Since the connection $\nabla$ is flat, it follows that the map
 $\mathfrak{h}\rightarrow\mathrm{End}(\mathfrak{h})\quad u\mapsto\nabla_{u},$
-defines a representation. We define a map $\rho:\mathfrak{h}\rightarrow\mathrm{End}(\mathfrak{h}^{*})$, the dual representation, as follows (for all $u\in\mathfrak{h}$ and $\xi\in\mathfrak{h}^{*}$):
+defines a representation.
+We define a map $\rho:\mathfrak{h}\rightarrow\mathrm{End}(\mathfrak{h}^{*})$, the dual representation, as follows: (for all homogeneous $u\in\mathfrak{h}$ and $\xi\in\mathfrak{h}^{*}$)
 (3) $\rho:\mathfrak{h}\rightarrow\mathrm{End}(\mathfrak{h}^{*})\quad u\mapsto\rho(u),\text{ where }\rho(u)\cdot\xi=-(-1)^{|u||\xi|}\xi\circ\nabla_{u}.$
@@
 See [BM]. ∎
+These two representations will be used to build Lagrangian extensions in the next subsection.
 3.2. Polarization and Lagrangian extensions
-Following [BC] (for the supercase, see [BM]), a polarization for a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is a choice of a Lagrangian subalgebra $\mathfrak{l}$ of $(\mathfrak{g},\omega)$.
-A strong polarization of a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is a pair $(\mathfrak{a},N)$ consisting of a Lagrangian ideal $\mathfrak{a}\subset\mathfrak{g}$ and a complementary Lagrangian subspace $N\subset\mathfrak{g}$.
-The quadruple $(\mathfrak{g},\omega,\mathfrak{a},N)$ is referred to as a strongly polarized quasi-Frobenius Lie superalgebra.
+Following [BC] (for the supercase, see [BBE, BM]), a polarization for a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is a choice of a homogeneous Lagrangian subalgebra $\mathfrak{l}$ of $(\mathfrak{g},\omega)$ (namely, $\mathfrak{l}=\mathfrak{l}^{\perp}$, with $\mathfrak{l}^{\perp}$ the orthogonal with respect to $\omega$).
+A strong polarization of a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is a pair $(\mathfrak{a},N)$ consisting of a homogeneous Lagrangian ideal $\mathfrak{a}\subset\mathfrak{g}$ and a complementary Lagrangian subspace $N\subset\mathfrak{g}$.
+Following [BC], the quadruple $(\mathfrak{g},\omega,\mathfrak{a},N)$ is referred to as a strongly polarized quasi-Frobenius Lie superalgebra.
+The following construction is borrowed from [BM], as a generalization to [BC].
 By means of a 2-cocycle $\alpha\in Z^{2}(\mathfrak{h},\mathfrak{h}^{*})$ (resp.
@@
 By construction, we will see that these Lie superalgebras are strongly polarized quasi-Frobenius.
-Recall that the space $\mathfrak{h}^{*}$ and $\Pi(\mathfrak{h}^{*})$ are $\mathfrak{h}$-modules by means of the representations (6) and (7).
+Recall that the two spaces $\mathfrak{h}^{*}$ and $\Pi(\mathfrak{h}^{*})$ are $\mathfrak{h}$-modules by means of the representations (3) and (4).
 On the space $\mathfrak{g}:=\mathfrak{h}\oplus\mathfrak{h}^{*}$: The brackets are defined as follows: (for all $u,v\in\mathfrak{h}$ and $\xi\in\mathfrak{h}^{*}$)
 $\begin{array}[]{l}[u,v]_{\mathfrak{g}}:=[u,v]_{\mathfrak{h}}+\alpha(u,v),\quad[u,\xi]_{\mathfrak{g}}:=\rho(u)\cdot\xi.\end{array}$
-An even form $\omega$ is defined on $\mathfrak{g}$ as follows: (for all $u+\xi,v+\zeta\in\mathfrak{h}\oplus\mathfrak{h}^{*}$)
-(8) $\omega(u+\xi,v+\zeta)=\xi(v)-(-1)^{|\zeta||u|}\zeta(u).$
+An even form $\omega$ is defined on $\mathfrak{g}$ as follows
+(5) $\omega(u+\xi,v+\zeta):=\xi(v)-(-1)^{|\zeta||u|}\zeta(u),$
+for all homogeneous $u+\xi,v+\zeta\in\mathfrak{h}\oplus\mathfrak{h}^{*}$.
 This construction will be referred to as the $T^{*}$-extension.
@@
 $\begin{array}[]{l}[u,v]_{\mathfrak{g}}:=[u,v]_{\mathfrak{h}}+\beta(u,v),\quad[u,\Pi(\xi)]_{\mathfrak{k}}:=\chi(u)\cdot\Pi(\xi).\end{array}$
-We define an odd form $\kappa$ on $\mathfrak{k}$ as follows: (for all $u+\Pi(\xi),v+\Pi(\zeta)\in\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*}$))
-(9) $\kappa(u+\Pi(\xi),v+\Pi(\zeta))=\xi(v)-(-1)^{(|\zeta|+1)|u|}\zeta(u).$
+We define an odd form $\kappa$ on $\mathfrak{k}$ as follows
+(6) $\kappa(u+\Pi(\xi),v+\Pi(\zeta)):=\xi(v)-(-1)^{(|\zeta|+1)|u|}\zeta(u),$
+for all homogeneous $u+\Pi(\xi),v+\Pi(\zeta)\in\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*})$.
 This construction will be referred to as the $\Pi T^{*}$-extension.
@@
 (i) The form $\omega$ is closed on $\mathfrak{g}$ if and only if
-(10) $(-1)^{|u||w|}\alpha(u,v)(w)+\circlearrowleft(u,v,w)=0\text{ for all $u,v,w\in\mathfrak{h}$}.$
+(7) $(-1)^{|u||w|}\alpha(u,v)(w)+\circlearrowleft(u,v,w)=0\text{ for all homogeneous $u,v,w\in\mathfrak{h}$},$
+where here and in what follows, the symbol $\circlearrowleft(u,v,w)$ denotes the sum over cyclic permutations of the variables $u,v,w$.
 (ii) The form $\kappa$ is closed on $\mathfrak{g}$ if and only if
-(11) $(-1)^{|u||w|}\Pi\circ\beta(u,v)(w)+\circlearrowleft(u,v,w)=0\text{ for all $u,v,w\in\mathfrak{h}$}.$
+(8) $(-1)^{|u||w|}\Pi\circ\beta(u,v)(w)+\circlearrowleft(u,v,w)=0\text{ for all homogeneous $u,v,w\in\mathfrak{h}$}.$
 Proof.
 See, [BM]. ∎
-We arrive to the following theorem proved in [BM].
+We arrive to the following theorem proved in [BM] (see also [BC] in the non-super case).
 Theorem 3.3 (Lagrangian or $T^{*}$- and $\Pi T^{*}$-extensions).
@@
 (6)).
-The converse of this theorem says that every orthosymplectic or periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ that has a Lagrangian ideal $\mathfrak{a}$ can be obtained as a $T^{*}$-extension or $\Pi T^{*}$-extension of a flat Lie superalgebra.
+The converse of this theorem says that every orthosymplectic or periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ that has a homogeneous Lagrangian ideal $\mathfrak{a}$ can be obtained as a $T^{*}$-extension or $\Pi T^{*}$-extension of a flat Lie superalgebra.
 Let $(\mathfrak{g},\omega,\mathfrak{a},N)$ be a strongly polarized quasi-Frobenius Lie superalgebra.
@@
 We will construct a flat connection $\nabla$ on $\mathfrak{h}$ (see [BM]), superizing the construction of [BC].
-For every $u,v\in\mathfrak{h}$, we denote by $\tilde{u}$ and $\tilde{v}$ their lift to $\mathfrak{g}$.
+For every homogeneous $u,v\in\mathfrak{h}$, we denote by $\tilde{u}$ and $\tilde{v}$ their lift to $\mathfrak{g}$.
 We then write
@@
 The pair $(\mathfrak{h},\nabla)$ is called the quotient flat Lie superalgebra associated with the Lagrangian ideal $\mathfrak{a}$ of $(\mathfrak{g},\omega)$.
+The following theorem is needed to determine whether the Lie superalgebras in Backhouse’s list arise as Lagrangian extensions of smaller Lie superalgebras. We refer to [BM] for its proof.
 Theorem 3.4 (Converse of Theorem 3.3).
@@
 (8), such that $(\mathfrak{g},\omega,\mathfrak{a},N)$ is isomorphic to the $\Pi T^{*}$-extension of $(\mathfrak{h},\nabla)$ by means of $\beta$.
-Proof.
-See [BM]. ∎
 4. The 4-dimensional Lie superalgebras as Lagrangian extensions
@@
     The Lie superalgebra $D^{5}$: We correct here the description of the Lie superalgebra $D_{5}$ in [BM].
-It is $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=e_{4}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$.
+It is $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{1},e_{4}]_{\mathfrak{h}}=e_{4}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$.
 The 2-form $\beta\equiv 0$ and the connection $\nabla$ on $\mathfrak{h}$ is given by:
@@
     $0=\omega(v,e_{3})=-l_{1}\implies l_{1}=0,$
-    $00=\omega(v,e_{4})=-l_{2}\implies l_{2}=$ 0 .
+    $0=\omega(v,e_{4})=-l_{2}\implies l_{2}=0.$
     This implies that $\mathfrak{a}^{\bot}=\mathfrak{a}$.
-The Lie superalgebra $\mathfrak{h}$ is given by $\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=0$.
-The connection $\nabla$ is given by:
+The Lie superalgebra $\mathfrak{h}$ is given by $\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $e_{1},e_{2}]_{\mathfrak{h}}=0$.
+The connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{2})=-e_{2},\quad\nabla_{e_{2}}(e_{1})=-e_{2},\quad\nabla_{e_{2}}(e_{2})=e_{1}.$
-    Now, we compute the torsion and curvature of $\nabla$, using Eqns. (4, 5). For the torsion, we have:
-    $T(e_{1},e_{2}):=\nabla_{e_{1}}(e_{2})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}(e_{1})-[e_{1},e_{2}]_{\mathfrak{h}}=-e_{2}-(-e_{2})-0=0.$
+    Now, we compute the torsion and curvature of $\nabla$, using Eqns. (1, 2). For the torsion, we have
+    $T(e_{1},e_{2})=\nabla_{e_{1}}(e_{2})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}(e_{1})-[e_{1},e_{2}]_{\mathfrak{h}}=-e_{2}-(-e_{2})-0=0.$
     For the curvature, we have
-    $\begin{array}[]{lcl}R^{\nabla}(e_{1},e_{2})(e_{1})&:=&\nabla_{e_{1}}\nabla_{e_{2}}(e_{1})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}\nabla_{e_{1}}(e_{1})-\nabla_{[e_{1},e_{2}]_{\mathfrak{h}}}(e_{1})\\[5.69054pt] &=&\nabla_{e_{1}}(-e_{2})-\nabla_{e_{2}}(-e_{1})=e_{2}-e_{2}=0.\end{array}$
+    $\begin{array}[]{lcl}R^{\nabla}(e_{1},e_{2})(e_{1})&=&\nabla_{e_{1}}\nabla_{e_{2}}(e_{1})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}\nabla_{e_{1}}(e_{1})-\nabla_{[e_{1},e_{2}]_{\mathfrak{h}}}(e_{1})\\[5.69054pt] &=&\nabla_{e_{1}}(-e_{2})-\nabla_{e_{2}}(-e_{1})=e_{2}-e_{2}=0.\end{array}$
     Similarly,
-    $\begin{array}[]{lcl}R^{\nabla}(e_{1},e_{2})(e_{2})&:=&\nabla_{e_{1}}\nabla_{e_{2}}(e_{2})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}\nabla_{e_{1}}(e_{2})-\nabla_{[e_{1},e_{2}]_{\mathfrak{h}}}(e_{2})\\[5.69054pt] &=&\nabla_{e_{1}}(e_{1})-\nabla_{e_{2}}(-e_{2})=-e_{1}+e_{1}=0.\end{array}$
+    $\begin{array}[]{lcl}R^{\nabla}(e_{1},e_{2})(e_{2})&=&\nabla_{e_{1}}\nabla_{e_{2}}(e_{2})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}\nabla_{e_{1}}(e_{2})-\nabla_{[e_{1},e_{2}]_{\mathfrak{h}}}(e_{2})\\[5.69054pt] &=&\nabla_{e_{1}}(e_{1})-\nabla_{e_{2}}(-e_{2})=-e_{1}+e_{1}=0.\end{array}$
     Therefore, the connection is torsion-free and flat.
@@
   4. (4)
-    The Lie superalgebra $D^{7}_{-1q}$, with $q\leq 1$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{4}=\Pi(e_{1}^{*}),e_{3}=\Pi(e_{2}^{*})$.
-The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
+    The Lie superalgebra $D^{7}_{-1q}$, with $q\leq 1$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{4}=\Pi(e_{1}^{*}),e_{3}=\Pi(e_{2}^{*})$.
+The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-qe_{1},\quad\nabla_{e_{1}}(e_{2})=e_{2},\quad\nabla_{e_{2}}(e_{1})=0,\quad\nabla_{e_{2}}(e_{2})=0.$
   5. (5)
-    The Lie superalgebra $D^{7}_{pq}$, with $p=-q,q\neq 0,-1$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=-pe_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}=e_{4}^{*}$.
-The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
+    The Lie superalgebra $D^{7}_{pq}$, with $p=-q,q\neq 0,-1$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{1},e_{4}]_{\mathfrak{h}}=-pe_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}=e_{4}^{*}$.
+The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-pe_{4},\quad\nabla_{e_{4}}(e_{1})=0,\quad\nabla_{e_{4}}(e_{4})=0.$
   6. (6)
-    The Lie superalgebra $D^{7}_{pq}$, with $p=-q,q=-1$: In the case where the form is odd, then this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{3}=\Pi(e_{1}^{*}),e_{4}=\Pi(e_{2}^{*})$.
-The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
+    The Lie superalgebra $D^{7}_{pq}$, with p=q,q=1p=-q,q=-1: In the case where the form is odd, then this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{3}=\Pi(e_{1}^{*}),e_{4}=\Pi(e_{2}^{*})$.
+The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{2})=e_{2},\quad\nabla_{e_{2}}(e_{1})=0,\quad\nabla_{e_{2}}(e_{2})=0.$
-    In the case where the form is even, then this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=-e_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}=e_{4}^{*}$.
-The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
+    In the case where the form is even, then this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{1},e_{4}]_{\mathfrak{h}}=-e_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}=e_{4}^{*}$.
+The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=0,\quad\nabla_{e_{4}}(e_{4})=0.$
@@
   8. (8)
-    The Lie superalgebra $D^{8}_{-1}$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{3}=\Pi(e_{1}^{*}),e_{4}=\Pi(e_{2}^{*})$.
-The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
+    The Lie superalgebra $D^{8}_{-1}$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{3}=\Pi(e_{1}^{*}),e_{4}=\Pi(e_{2}^{*})$.
+The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=e_{1}-e_{2},\quad\nabla_{e_{1}}(e_{2})=e_{2},\quad\nabla_{e_{2}}(e_{1})=0,\quad\nabla_{e_{2}}(e_{2})=0.$
@@
   10. (10)
-    The Lie superalgebra $D^{10}_{q}$ with $q\not=-1$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=qe_{4}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{2}=\Pi(e_{4}^{*})|e_{3}=-(q+1)\Pi(e_{1}^{*})$.
-The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
+    The Lie superalgebra $D^{10}_{q}$ with $q\not=-1$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{1},e_{4}]_{\mathfrak{h}}=qe_{4}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{2}=\Pi(e_{4}^{*})|e_{3}=-(q+1)\Pi(e_{1}^{*})$.
+The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-(q+1)e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=-(q+1)e_{4},\quad\nabla_{e_{4}}(e_{4})=0.$
   11. (11)
-[…]
+    The Lie superalgebra $(D^{7}_{1/2\;1/2})^{1}$: This Lie superalgebra does not admit a 2-dimensional Lagrangian ideal.
+Indeed, let us suppose the contrary, and let us denote by $\mathfrak{a}=\mathrm{Span}\{X,Y\}$ such an ideal.
+Let us write $X=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+x_{4}e_{4}$ and $Y=y_{1}e_{1}+y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$.
+We compute the brackets $e_{i},X]$, for all $i\in\{1,2,3,4\}$.
+we obtain
     $\begin{array}[]{ll}[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3}+\dfrac{1}{2}x_{4}e_{4},&[e_{2},X]=-x_{1}e_{2},\\[5.69054pt] [e_{3},X]=-\dfrac{1}{2}x_{1}e_{3}+x_{3}e_{2},&[e_{4},X]=-\dfrac{1}{2}x_{1}e_{4}+x_{4}e_{2}.\end{array}$
@@
       If $x_{3}=0,$ and $x_{4}\neq 0$.
-Looking at brackets again we get $[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{4}e_{4},$ and $[e_{4},X]=x_{4}e_{2}$, which implies $e_{2},e_{4}\in\mathfrak{a}$, so $\mathfrak{a}=\mathrm{Span}\{e_{2},e_{4}\}$.
+Looking at brackets again we get $e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{4}e_{4},$ and $e_{4},X]=x_{4}e_{2}$, which implies $e_{2},e_{4}\in\mathfrak{a}$, so $\mathfrak{a}=\mathrm{Span}\{e_{2},e_{4}\}$.
 We have to check that $\mathfrak{a}$ is not Lagrangian.
@@
       If $x_{3}\neq 0,$ and $x_{4}\neq 0$.
-Since all the previous cases showed that we can’t find a Lagrangian ideal, and since the proof is symmetrical in X and Y without loss of generality we can state that $y_{3}\neq 0,y_{4}\neq 0$.
+Since all the previous cases showed that we cannot find a Lagrangian ideal, and since the proof is symmetrical in X and Y without loss of generality we can state that $y_{3}\neq 0,y_{4}\neq 0$.
 This implies that $e_{2}\in\mathfrak{a}$ and that $x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3}+\dfrac{1}{2}x_{4}e_{4}\in\mathfrak{a}$.
@@
   12. (12)
-[…]
+    The Lie superalgebra $(D^{7}_{1/2\;1/2})^{2}$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|x:=e_{3}-e_{4}\}$ with the bracket $e_{1},x]_{\mathfrak{h}}=\dfrac{1}{2}x$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}+e_{4}=-2x^{*}$.
+The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(x)=-\frac{1}{2}x,\quad\nabla_{x}(e_{1})=-x,\quad\nabla_{x}(x)=0.$
   13. (13)
-[…]
+    The Lie superalgebra $(D^{7}_{1/2\;1/2})^{3}$: This Lie superalgebra is not quasi-Frobenius.
   14. (14)
-    The Lie superalgebra $(D^{7}_{1-p,p})$, with $p\leq\dfrac{1}{2}$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=(1-p)e_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-e_{1}^{*}|e_{3}=e_{4}^{*}$.
-The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
+    The Lie superalgebra $(D^{7}_{1-p,p})$, with $p\leq\dfrac{1}{2}$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{1},e_{4}]_{\mathfrak{h}}=(1-p)e_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-e_{1}^{*}|e_{3}=e_{4}^{*}$.
+The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-pe_{4},\quad\nabla_{e_{4}}(e_{1})=-e_{4},\quad\nabla_{e_{4}}(e_{4})=0.$
@@
 Let us write $X=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+x_{4}e_{4}$ and $Y=y_{1}e_{1}+y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$.
-We compute the brackets $[e_{i},X]$ for all $i\in\{1,2,3,4\}$.
-We get
+We compute the brackets $e_{i},X]$, for all $i\in\{1,2,3,4\}$.
+We obtain
     $\begin{array}[]{l}[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3}-x_{3}pe_{4}+x_{4}pe_{3}+\dfrac{1}{2}x_{4}e_{4},\\[5.69054pt] [e_{2},X]=-x_{1}e_{2},\\[5.69054pt] [e_{3},X]=-\dfrac{1}{2}x_{1}e_{3}+px_{1}e_{4}+x_{3}e_{2},\\[8.53581pt] [e_{4},X]=-\dfrac{1}{2}x_{1}e_{4}-px_{1}e_{3}+x_{4}e_{2}.\end{array}$
@@
 This implies that $x_{1}=0$.
-We get then:
+We get then
     $\begin{array}[]{ll}[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3}-x_{3}pe_{4}+x_{4}pe_{3}+\dfrac{1}{2}x_{4}e_{4},&[e_{2},X]=0,\\[5.69054pt] [e_{3},X]=x_{3}e_{2},&[e_{4},X]=x_{4}e_{2}.\end{array}$
@@
   17. (17)
-    The Lie superalgebra $(D^{10}_{0})^{1}$: In the case where the form is odd, this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$.
-The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
+    The Lie superalgebra $(D^{10}_{0})^{1}$: In the case where the form is odd, this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{4},e_{4}]_{\mathfrak{h}}=e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$.
+The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=-e_{4},\quad\nabla_{e_{4}}(e_{4})=\frac{1}{2}e_{1}.$
-    In the case where the form is even, this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=2e_{1}^{*}|e_{3}=e_{4}^{*}$.
-The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
+    In the case where the form is even, this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{4},e_{4}]_{\mathfrak{h}}=e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=2e_{1}^{*}|e_{3}=e_{4}^{*}$.
+The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=-e_{4},\quad\nabla_{e_{4}}(e_{4})=\frac{1}{2}e_{1}.$
   18. (18)
-    The Lie superalgebra $(D^{10}_{0})^{2}$: In the case where the form is odd, this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=-e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$.
-The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
+    The Lie superalgebra $(D^{10}_{0})^{2}$: In the case where the form is odd, this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{4},e_{4}]_{\mathfrak{h}}=-e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$.
+The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=-e_{4},\quad\nabla_{e_{4}}(e_{4})=-\frac{1}{2}e_{1}.$
-    In the case where the form is even, this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=-e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=2e_{1}^{*}|e_{3}=e_{4}^{*}$.
-The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
+    In the case where the form is even, this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{4},e_{4}]_{\mathfrak{h}}=-e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=2e_{1}^{*}|e_{3}=e_{4}^{*}$.
+The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(e_{4})=-e_{4},\quad\nabla_{e_{4}}(e_{1})=-e_{4},\quad\nabla_{e_{4}}(e_{4})=-\frac{1}{2}e_{1}.$
@@
     $\alpha=e^{*}_{1}\otimes(x^{*}\wedge x^{*})+x^{*}\otimes(e^{*}_{1}\wedge x^{*})$
-    The connection $\nabla$ is given by:
+    The connection $\nabla$ is given by
     $\nabla_{e_{1}}(e_{1})=-e_{1},\quad\nabla_{e_{1}}(x)=-x,\quad\nabla_{x}(e_{1})=-x,\quad\nabla_{x}(x)=0.$
   24. (24)
-    The Lie superalgebra $(C^{1}_{1/2}+A)$: This Lie superalgebra does not admit a 2-dimensional Lagrangian ideal.
+    The Lie superalgebra $(C^{1}_{1/2}+A)$: This Lie superalgebra does not admit a Lagrangian ideal of dimension 2.
 Indeed, let us suppose there is one, denoted by $\mathfrak{a}=\mathrm{Span}\{X,Y\}$.
 Let us write $X=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+x_{4}e_{4}$ and $Y=y_{1}e_{1}+y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$.
-We compute the brackets $[e_{i},X]$ for all $i\in\{1,2,3,4\}$.
-We get:
+We compute the brackets $e_{i},X]$, for all $i\in\{1,2,3,4\}$.
+We obtain
     $\begin{array}[]{l}[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3},\quad[e_{2},X]=-x_{1}e_{2},\quad[e_{3},X]=-\dfrac{1}{2}x_{1}e_{3}+x_{3}e_{2},\quad[e_{4},X]=0.\end{array}$
@@
   26. (26)
-    The Lie superalgebra $(C^{3}+A)$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{2}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=e_{2}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{1}=2e_{2}^{*}\,|\,e_{3}=e_{4}^{*}$.
-The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
+    The Lie superalgebra $(C^{3}+A)$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{2}|e_{4}\}$ with the bracket $e_{4},e_{4}]_{\mathfrak{h}}=e_{2}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{1}=2e_{2}^{*}\,|\,e_{3}=e_{4}^{*}$.
+The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
     $\nabla_{e_{2}}(e_{2})=0,\quad\nabla_{e_{2}}(e_{4})=0,\quad\nabla_{e_{4}}(e_{2})=0,\quad\nabla_{e_{4}}(e_{4})=\frac{1}{2}e_{2}.$
@@
   28. (28)
-    The Lie superalgebra $D^{1}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
+    The Lie superalgebra $D^{1}$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
   29. (29)
-    The Lie superalgebra $D^{2}_{q},\text{ with }q\not=-1,0$: This Lie superalgebra is not quasi-Frobenuis.
+    The Lie superalgebra $D^{2}_{q},\text{ with }q\not=-1,0$: This Lie superalgebra is indeed not quasi-Frobenuis.
   30. (30)
-    The Lie superalgebra $D^{2}_{-1}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
+    The Lie superalgebra $D^{2}_{-1}$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
   31. (31)
@@
   33. (33)
-    The Lie superalgebra $D^{11}_{pq}\text{ with }(p,q)=(-1,-1)$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
+    The Lie superalgebra $D^{11}_{pq}\text{ with }(p,q)=(-1,-1)$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
   34. (34)
-    The Lie superalgebra $D^{11}_{pq}\text{ with }0<|p|\leq 1,p=-q$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
+    The Lie superalgebra $D^{11}_{pq}\text{ with }0<|p|\leq 1,p=-q$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
   35. (35)
-    The Lie superalgebra $D^{11}_{pq}\text{ with }0<|p|<1,q=-1$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
+    The Lie superalgebra $D^{11}_{pq}\text{ with }0<|p|<1,q=-1$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
   36. (36)
-    The Lie superalgebra $D^{12}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
+    The Lie superalgebra $D^{12}$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
   37. (37)
@@
   38. (38)
-    The Lie superalgebra $D^{13}_{-1}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
+    The Lie superalgebra $D^{13}_{-1}$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
   39. (39)
@@
   40. (40)
-    The Lie superalgebra $D^{15}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
+    The Lie superalgebra $D^{15}$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
   41. (41)
-    The Lie superalgebra $D^{16}$: This Lie superalgebra is in fact not quasi-Frobenuis.
+    The Lie superalgebra $D^{16}$: This Lie superalgebra is not quasi-Frobenuis.
   42. (42)
@@
   44. (44)
-    The Lie superalgebra $(D^{3}_{-1/2,-1/2})$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
+    The Lie superalgebra $(D^{3}_{-1/2,-1/2})$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
   45. (45)
@@
     The Lie superalgebra $(A_{1,1}+3A)^{2}$: This Lie superalgebra is not quasi-Frobenuis.
+Tables 15 below summarize the outcome of the classification.
+Among the Lie superalgebras considered, 22 do not admit a quasi-Frobenius structure, whereas 9 admit a nonhomogeneous structure.
+Of the quasi-Frobenius cases, 4 do not arise as Lagrangian extensions.
+Moreover, 8 are $T^{*}$-extensions and 10 are $\Pi T^{*}$-extensions.
+It is worth mentioning that three Lie superalgebras admit both a closed orthosymplectic and periplectic form, namely $(D_{0}^{10})^{1}$, $(D_{0}^{10})^{2}$, and $D_{pq}^{7}$ with (p,q)=(1,1)(p,q)=(1,-1).
+Lie superalgebras without quasi-Frobenius structures
+$(D_{1/2\;1/2}^{7})^{3}$ $(D^{8}_{1/2})$ $(2A_{1,1}+2A)^{1}$
+$(2A_{1,1}+2A)^{2}$ $D^{7}_{pq}$ $(A_{1,1}+3A)^{1}$
+$(A_{1,1}+3A)^{2}$ $(D^{2}_{-1/2})^{1}$ $(D^{2}_{-1/2})^{2}$
+$(D^{3}_{p,-1/2}),\,p\not=0$ $(A_{3,1}+A)$ $D^{16}$
+$D^{14}_{pq},\,p\not=0,q\geq 0$ $D^{13}_{p},\,p\not=0$ $D^{3}_{pq}$
+$D^{11}_{pq},\,0<|p|\leq|q|\leq 1$ $D^{2}_{q},\,q\not=-1,0$ $D^{9}_{pq},\,q>0$
+$D^{8}_{p},\,p\not=0$ $(C_{0}^{5}+A)$ $(C^{2}_{-1}+A)$
+$(2A_{1,1}+2A)^{4}_{p}$
+Table 1. No quasi-Frobenius structures
+Lie superalgebras with a nonhomogeneous quasi-Frobenius structure
+$(D^{3}_{-1/2\;1/2})$ $D^{15}$ $D^{13}_{-1}$
+$D^{12}$ $D^{11}_{pq},\,0<|p|<1,\,q=-1$ $D^{11}_{pq},\,0<|p|<1,\,p=-q$
+$D^{11}_{pq},\,p=q=-1$ $D^{2}_{-1}$ $D^{1}$.
+Table 2. Nonhomogeneous
+Quasi-Frobenuis Lie superalgebras that are not Lagrangian extensions
+$(C_{1/2}^{1}+A)$ $(D^{9}_{1/2\;p}),\,p>0$ $(D^{7}_{1/2\;1/2})^{1}$
+$D^{6}$
+Table 3. Not Lagrangian extensions
+Quasi-Frobenuis Lie superalgebras that are $T^{*}$-extensions
+$(C^{3}+A)$ $(C_{1}^{1}+A)$ $(D_{0}^{10})^{1},\,(|\omega|=\bar{0})$
+$(D_{0}^{10})^{2},\,(|\omega|=\bar{0})$ $D^{7}_{1-p,p},\,p\leq\frac{1}{2}$ $(D^{7}_{1/2\;1/2})^{2}$
+$D^{7}_{pq},\,p=-q,q\not=0,-1$ $D^{7}_{pq},\,p=-q=1$ $(|\omega|=\bar{0})$
+Table 4. $T^{*}$-extensions
+Quasi-Frobenuis Lie superalgebras that are $\Pi T^{*}$-extensions
+$(2A_{1,1}+2A)^{3}_{p},\,p=\frac{1}{2}$ $(D_{0}^{10})^{2},\,(|\omega|=\bar{1})$ $(D_{0}^{10})^{1},\,(|\omega|=\bar{1})$
+$D^{10}_{q},\,q\not=-1$ $D^{6}$ $D^{7}_{-1,q},\,q\leq 1$
+$D^{7}_{p,q},\,p=-q=1,\,(|\omega|=\bar{1})$ $D^{5}$ $D^{8}_{-1}$
+$D^{7}_{p,q},\,p=-q=1$
+Table 5. $\Pi T^{*}$-extensions
 5. Left-symmetric structures on the 4-dimensional real Lie superalgebras
-In the list of 4-dimensional real Lie superalgebras, we will list left-symmetric structures that are compatible with the Lie structure.
+In the list of indecomposable 4-dimensional real Lie superalgebras, we will list left-symmetric structures that are compatible with the Lie structure.
 With the exception of $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$, all of these structures are Novikov.
@@
   1. (1)
-    The Lie superalgebra $D^{5}$: A Novikov structure is given by:
+    The Lie superalgebra $D^{5}$: A Novikov structure is given by
     $e_{1}\cdot e_{3}=e_{3},\quad e_{1}\cdot e_{4}=e_{4},\quad e_{2}\cdot e_{4}=e_{3}.$
-    Additionally, this structure is also Balinsky-Novikov. As a matter of fact, the non-zero associators are given by:
+    Additionally, this structure is also Balinsky-Novikov. As a matter of fact, the non-zero associators are given by
     $(e_{1},e_{1},e_{4})=-e_{4},\quad(e_{2},e_{1},e_{4})=-e_{3},\quad(e_{1},e_{1},e_{3})=-e_{3},\quad(e_{1},e_{2},e_{4})=-e_{3}.$
   2. (2)
-    The Lie superalgebra $D^{6}$: A Novikov structure is given by:
+    The Lie superalgebra $D^{6}$: A Novikov structure is given by
     $e_{1}\cdot e_{3}=e_{3},\quad e_{1}\cdot e_{4}=e_{4},\quad e_{2}\cdot e_{3}=-e_{4},\quad e_{2}\cdot e_{4}=e_{3}.$
@@
   3. (3)
-    The Lie superalgebra $D^{7}_{pq}$ with ($p\geq q,\;pq\not=0$): A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
+    The Lie superalgebra $D^{7}_{pq}$ with ($p\geq q,\;pq\not=0$): A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $e_{1}\cdot e_{1}=\gamma e_{2},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{1}\cdot e_{3}=pe_{3},\quad e_{1}\cdot e_{4}=qe_{4}.$
@@
   4. (4)
-    The Lie superalgebra $D^{8}_{p}$ with $p\not=0$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
+    The Lie superalgebra $D^{8}_{p}$ with $p\not=0$: A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $e_{1}\cdot e_{1}=\gamma e_{2},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{1}\cdot e_{3}=pe_{3},\quad e_{1}\cdot e_{4}=e_{3}+pe_{4}.$
@@
   5. (5)
-    The Lie superalgebra $D^{9}_{pq}$ with $q>0$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
+    The Lie superalgebra $D^{9}_{pq}$ with $q>0$: A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=(p+\gamma)e_{3}-qe_{4},&e_{1}\cdot e_{4}=qe_{3}+(p+\gamma)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{4}\cdot e_{1}=\gamma e_{4}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R}$):
+    A Balinsky-Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{lllll}e_{1}\cdot e_{1}=\gamma e_{2},&e_{1}\cdot e_{2}=e_{2},&e_{1}\cdot e_{3}=pe_{3}-qe_{4},&e_{1}\cdot e_{4}=qe_{3}+pe_{4}.\end{array}$
   6. (6)
-    The Lie superalgebra $D^{10}_{q}:$ A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
+    The Lie superalgebra $D^{10}_{q}:$ A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=-qe_{1}+\gamma(q-1)e_{2},&e_{1}\cdot e_{2}=(1-q)e_{2},&e_{1}\cdot e_{3}=e_{3},&e_{1}\cdot e_{4}=\gamma e_{3},\\[5.69054pt] e_{2}\cdot e_{1}=-qe_{2},&e_{3}\cdot e_{1}=-qe_{3},&e_{4}\cdot e_{1}=\gamma e_{3}-qe_{4},&e_{4}\cdot e_{2}=-e_{3}.\end{array}$
@@
   7. (7)
-[…]
+    The Lie superalgebra $(D^{7}_{1/2\;1/2})^{1}:$ A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{lllll}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\frac{1}{2}e_{2},&e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},&e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},&e_{3}\cdot e_{3}=\frac{1}{2}e_{2}\\[5.69054pt] e_{3}\cdot e_{4}=\gamma e_{2},&e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},&e_{4}\cdot e_{3}=-\gamma e_{2},&e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3},&e_{1}\cdot e_{4}=\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=(\gamma-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{4}=e_{2}.\end{array}$
   8. (8)
-[…]
+    The Lie superalgebra $(D^{7}_{1/2\;1/2})^{2}:$ A Novikov structure, depending on a parameter $\lambda\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\frac{1}{2}e_{2},&e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},&e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\\[5.69054pt] e_{3}\cdot e_{3}=\frac{1}{2}e_{2},&e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=-\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(1+\gamma)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{4}=-e_{2}.\end{array}$
   9. (9)
-[…]
+    The Lie superalgebra $(D^{7}_{1/2\;1/2})^{3}:$ A Novikov structure, depending on a parameter $\lambda\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\frac{1}{2}e_{2},&e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},&e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\\[5.69054pt] e_{3}\cdot e_{3}=\frac{1}{2}e_{2},&e_{4}\cdot e_{1}=-\frac{1}{2}e_{4}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(1+\gamma)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4}.\end{array}$
   10. (10)
-    The Lie superalgebra $(D^{7}_{1-p,p})$ with $p\leq\frac{1}{2}:$ A Novikov structure is given by (where $\lambda\in\mathbb{R}$):
+    The Lie superalgebra $(D^{7}_{1-p,p})$ with $p\leq\frac{1}{2}:$ A Novikov structure, depending on a parameter $\lambda\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=-pe_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1-p)e_{2},&e_{1}\cdot e_{4}=(1-2p)e_{4},&e_{2}\cdot e_{1}=-pe_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-pe_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=-pe_{4}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(2p+\gamma-1)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(1-2p+\gamma)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{4}\cdot e_{1}=(\gamma-1)e_{4}\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
   11. (11)
-    The Lie superalgebra $(D^{8}_{1/2}):$ A Novikov structure is given by (where $\lambda\in\mathbb{R}$):
+    The Lie superalgebra $(D^{8}_{1/2}):$ A Novikov structure, depending on a parameter $\lambda\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\frac{1}{2}e_{2},&e_{1}\cdot e_{4}=e_{3},&e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},&e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
-    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=e_{3}+\frac{1}{2}(1+\gamma)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{4}\cdot e_{1}=\gamma e_{4},&e_{4}\cdot e_{4}=e_{2}.\\[5.69054pt] \end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
+[…]
   12. (12)
-    The Lie superalgebra $(D^{9}_{1/2,p})$ with $p>0:$ A Novikov structure is given by (where $\lambda\in\mathbb{R}$):
+    The Lie superalgebra $(D^{9}_{1/2,p})$ with $p>0:$ A Novikov structure, depending on a parameter $\lambda\in\mathbb{R}$, is given by
     $\begin{array}[]{lllll}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\frac{1}{2}e_{2},&e_{1}\cdot e_{3}=-pe_{4},&e_{1}\cdot e_{4}=pe_{3},&e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},&e_{3}\cdot e_{3}=\frac{1}{2}e_{2},&e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3}-pe_{4},&e_{1}\cdot e_{4}=pe_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=(\gamma-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{4}=e_{2}.\end{array}$
   13. (13)
-    The Lie superalgebra $(D^{10}_{0})^{1}:$ An LSSA structure is given by (where $\gamma\in\mathbb{R}$):
+    The Lie superalgebra $(D^{10}_{0})^{1}:$ An LSSA structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{lllll}e_{1}\cdot e_{1}=-e_{1},&e_{1}\cdot e_{4}=-e_{4},&e_{2}\cdot e_{1}=-e_{2},&e_{2}\cdot e_{4}=(1+\gamma)e_{3},&e_{3}\cdot e_{1}=-e_{3},\\[5.69054pt] e_{3}\cdot e_{4}=-\frac{1}{2}e_{2},&e_{4}\cdot e_{1}=-e_{4},&e_{4}\cdot e_{2}=\gamma e_{3},&e_{4}\cdot e_{4}=\frac{1}{2}e_{1}.\end{array}$
@@
     Now, $N(e_{3},e_{4},e_{4})=\frac{1}{2}e_{3}$ which is a not zero. Therefore, it cannot be Novikov. ∎
-    However, there is a Balinsky-Novikov structure given by (where $\lambda,\gamma\in\mathbb{R}$):
+    However, there is a Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(2+\gamma)e_{3},&e_{1}\cdot e_{4}=-\lambda e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{2}\cdot e_{4}=-\gamma e_{3},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{4}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{4}\cdot e_{1}=-2\lambda e_{3}+\gamma e_{4},&e_{4}\cdot e_{2}=-2(1+\gamma)e_{3}&e_{4}\cdot e_{3}=-\frac{1}{2}e_{2},&e_{4}\cdot e_{4}=e_{1}.\end{array}$
   14. (14)
-    The Lie superalgebra $(D^{10}_{0})^{2}:$ AAn LSSA structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
+    The Lie superalgebra $(D^{10}_{0})^{2}:$ An LSSA structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{lllll}e_{1}\cdot e_{1}=-e_{1}-2\lambda\gamma e_{2},&e_{1}\cdot e_{4}=\gamma e_{3}-e_{4},&e_{2}\cdot e_{1}=-e_{2},&e_{2}\cdot e_{4}=e_{3},&e_{3}\cdot e_{1}=-e_{3},\\[5.69054pt] e_{3}\cdot e_{4}=\frac{1}{2}(1-2\lambda)e_{2},&e_{4}\cdot e_{1}=\gamma e_{3}-e_{4},&e_{4}\cdot e_{3}=\lambda e_{2},&e_{4}\cdot e_{4}=-\frac{1}{2}e_{1}.\end{array}$
     Claim. There is no Novikov structure on $(D^{10}_{0})^{2}$ compatible with the Lie structure.
-    However, there is a Balinsky-Novikov structure given by (where $\lambda,\gamma\in\mathbb{R}$):
-    $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=-\lambda e_{3}+\frac{1}{2}(\gamma-1)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{2}\cdot e_{4}=(1-\gamma)e_{3},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=\frac{1}{2}e_{2},\\[5.69054pt] e_{4}\cdot e_{1}=-2\lambda e_{3}+(\gamma-1)e_{4},&e_{4}\cdot e_{2}=-2\gamma e_{3}&e_{4}\cdot e_{3}=\frac{1}{2}e_{2},&e_{4}\cdot e_{4}=-e_{1}.\end{array}$
+    However, there is a Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by given by
+    $\begin{array}[]{lll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},\\[5.69054pt] e_{1}\cdot e_{4}=-\lambda e_{3}+\frac{1}{2}(\gamma-1)e_{4},&e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{2}\cdot e_{4}=(1-\gamma)e_{3},\\[5.69054pt] e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=\frac{1}{2}e_{2},&e_{4}\cdot e_{1}=-2\lambda e_{3}+(\gamma-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{2}=-2\gamma e_{3}&e_{4}\cdot e_{3}=\frac{1}{2}e_{2},&e_{4}\cdot e_{4}=-e_{1}.\end{array}$
   15. (15)
-    The Lie superalgebra $(2A_{1,1}+2A)^{1}:$ A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R})$:
+    The Lie superalgebra $(2A_{1,1}+2A)^{1}:$ A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\lambda e_{1}+\gamma e_{2},\quad e_{4}\cdot e_{3}=-\lambda e_{1}-\gamma e_{2},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1},&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3},&e_{2}\cdot e_{2}=2\lambda e_{2},&e_{2}\cdot e_{4}=\lambda e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{1},&e_{4}\cdot e_{2}=2\lambda e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
   16. (16)
-    The Lie superalgebra $(2A_{1,1}+2A)^{2}:$ A Novikov structure is given by (where $\gamma\in{\mathbb{R}})$::
-    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(1-\gamma_{)}e_{1},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
-    $\begin{array}[]{llll}e_{1}\cdot e_{1}=2\gamma e_{1},&e_{1}\cdot e_{2}=2\gamma e_{1},&e_{1}\cdot e_{3}=\gamma e_{3},&e_{1}\cdot e_{4}=\gamma e_{3},\\[5.69054pt] e_{2}\cdot e_{1}=2\gamma e_{1},&e_{2}\cdot e_{2}=-2(\lambda-\gamma)e_{1}+2\lambda e_{2},&e_{2}\cdot e_{3}=\gamma e_{3},&e_{2}\cdot e_{4}=(\gamma-\lambda)e_{3}+\lambda e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=2\gamma e_{3}&e_{3}\cdot e_{2}=2\gamma e_{3},&e_{3}\cdot e_{3}=e_{1}&e_{3}\cdot e_{4}=e_{1},\\[5.69054pt] e_{4}\cdot e_{1}=2\gamma e_{3}&e_{4}\cdot e_{2}=-2(\lambda-\gamma)e_{3}+2\lambda e_{4}&e_{4}\cdot e_{3}=e_{1}&e_{4}\cdot e_{4}=e_{2}.\end{array}$
+    The Lie superalgebra $(2A_{1,1}+2A)^{2}:$ A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(1-\gamma)e_{1},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{llll}e_{1}\cdot e_{1}=2\gamma e_{1},&e_{1}\cdot e_{2}=2\gamma e_{1},&e_{1}\cdot e_{3}=\gamma e_{3},&e_{1}\cdot e_{4}=\gamma e_{3},\\[5.69054pt] e_{2}\cdot e_{1}=2\gamma e_{1},&e_{2}\cdot e_{2}=-2(\lambda-\gamma)e_{1}+2\lambda e_{2},&e_{2}\cdot e_{3}=\gamma e_{3},&e_{2}\cdot e_{4}=(\gamma-\lambda)e_{3}+\lambda e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=2\gamma e_{3},&e_{3}\cdot e_{2}=2\gamma e_{3},&e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=e_{1},\\[5.69054pt] e_{4}\cdot e_{1}=2\gamma e_{3},&e_{4}\cdot e_{2}=-2(\lambda-\gamma)e_{3}+2\lambda e_{4},&e_{4}\cdot e_{3}=e_{1},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
   17. (17)
-    The Lie superalgebra $(2A_{1,1}+2A)^{3}_{p}:$ A Novikov structure is given by (where $\gamma\in{\mathbb{R}})$:
-    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(p-\gamma_{)}e_{1}+pe_{2},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
-    $\begin{array}[]{lll}e_{1}\cdot e_{1}=\frac{2}{p}(p^{2}\gamma+\lambda(2p^{2}-1))e_{1}+2p\lambda e_{2},&e_{1}\cdot e_{2}=-2p\lambda e_{1}+2p\gamma e_{2},\\[5.69054pt] e_{1}\cdot e_{3}=\frac{1}{p}(p^{2}\gamma+\lambda(p^{2}-1))e_{3}+\lambda e_{4},&e_{1}\cdot e_{4}=-\lambda e_{3}+p(\gamma+\lambda)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=-2p\lambda e_{1}+2p\gamma e_{2},&e_{2}\cdot e_{2}=-2p\gamma e_{1}-\frac{2}{p}(p^{2}\lambda+(2p^{2}-1)\gamma)e_{2},\\[5.69054pt] e_{2}\cdot e_{3}=-p(\lambda+\gamma)e_{3}+\gamma e_{4},&e_{2}\cdot e_{4}=-\gamma e_{3}+\frac{1}{p}((1-p^{2})\gamma-p^{2}\lambda)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\frac{2}{p}(p^{2}\gamma+(p^{2}-1)\lambda)e_{3}+2\lambda e_{4},&e_{3}\cdot e_{2}=-2p(\gamma+\lambda)e_{3}+2\gamma e_{4},\\[5.69054pt] e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=pe_{1}+pe_{2},\\[5.69054pt] e_{4}\cdot e_{1}=-2\lambda e_{3}+2p(\gamma+\lambda)e_{4}&e_{4}\cdot e_{2}=-2\gamma e_{3}-\frac{2}{p}((p^{2}-1)\gamma+p^{2}\lambda))e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=pe_{1}+pe_{2},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
+    The Lie superalgebra $(2A_{1,1}+2A)^{3}_{p}:$ A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(p-\gamma)e_{1}+pe_{2},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{lll}e_{1}\cdot e_{1}=\frac{2}{p}(p^{2}\gamma+\lambda(2p^{2}-1))e_{1}+2p\lambda e_{2},&e_{1}\cdot e_{2}=-2p\lambda e_{1}+2p\gamma e_{2},\\[5.69054pt] e_{1}\cdot e_{3}=\frac{1}{p}(p^{2}\gamma+\lambda(p^{2}-1))e_{3}+\lambda e_{4},&e_{1}\cdot e_{4}=-\lambda e_{3}+p(\gamma+\lambda)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=-2p\lambda e_{1}+2p\gamma e_{2},&e_{2}\cdot e_{2}=-2p\gamma e_{1}-\frac{2}{p}(p^{2}\lambda+(2p^{2}-1)\gamma)e_{2},\\[5.69054pt] e_{2}\cdot e_{3}=-p(\lambda+\gamma)e_{3}+\gamma e_{4},&e_{2}\cdot e_{4}=-\gamma e_{3}+\frac{1}{p}((1-p^{2})\gamma-p^{2}\lambda)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\frac{2}{p}(p^{2}\gamma+(p^{2}-1)\lambda)e_{3}+2\lambda e_{4},&e_{3}\cdot e_{2}=-2p(\gamma+\lambda)e_{3}+2\gamma e_{4},\\[5.69054pt] e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=pe_{1}+pe_{2},\\[5.69054pt] e_{4}\cdot e_{1}=-2\lambda e_{3}+2p(\gamma+\lambda)e_{4},&e_{4}\cdot e_{2}=-2\gamma e_{3}-\frac{2}{p}((p^{2}-1)\gamma+p^{2}\lambda))e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=pe_{1}+pe_{2},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
   18. (18)
-    The Lie superalgebra $(2A_{1,1}+2A)^{4}_{p}:$ A Novikov structure is given by (where $\gamma\in{\mathbb{R}})$:
+    The Lie superalgebra $(2A_{1,1}+2A)^{4}_{p}:$ A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(p-\gamma)e_{1}-pe_{2},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
-    $\begin{array}[]{lll}e_{1}\cdot e_{1}=-\frac{2}{p}(p^{2}\gamma-\lambda(2p^{2}+1))e_{1}-2p\lambda e_{2},&e_{1}\cdot e_{2}=2p\lambda e_{1}-2p\gamma e_{2},\\[5.69054pt] e_{1}\cdot e_{3}=\frac{1}{p}(-p^{2}\gamma+\lambda(p^{2}+1))e_{3}+\lambda e_{4},&e_{1}\cdot e_{4}=\lambda e_{3}+p(-\gamma+\lambda)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2p\lambda e_{1}-2p\gamma e_{2},&e_{2}\cdot e_{2}=2p\gamma e_{1}-\frac{2}{p}(-p^{2}\lambda+(2p^{2}+1)\gamma)e_{2},\\[5.69054pt] e_{2}\cdot e_{3}=p(-\gamma+\lambda)e_{3}+\gamma e_{4},&e_{2}\cdot e_{4}=\gamma e_{3}+\frac{1}{p}((-1-p^{2})\gamma+p^{2}\lambda)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\frac{2}{p}(-p^{2}\gamma+(p^{2}+1)\lambda)e_{3}+2\lambda e_{4},&e_{3}\cdot e_{2}=-2p(\gamma-\lambda)e_{3}+2\gamma e_{4},\\[5.69054pt] e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=pe_{1}-pe_{2},\\[5.69054pt] e_{4}\cdot e_{1}=2\lambda e_{3}-2p(\gamma-\lambda)e_{4}&e_{4}\cdot e_{2}=2\gamma e_{3}-\frac{2}{p}((p^{2}+1)\gamma-p^{2}\lambda))e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=pe_{1}-pe_{2},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{lll}e_{1}\cdot e_{1}=-\frac{2}{p}(p^{2}\gamma-\lambda(2p^{2}+1))e_{1}-2p\lambda e_{2},&e_{1}\cdot e_{2}=2p\lambda e_{1}-2p\gamma e_{2},\\[5.69054pt] e_{1}\cdot e_{3}=\frac{1}{p}(-p^{2}\gamma+\lambda(p^{2}+1))e_{3}+\lambda e_{4},&e_{1}\cdot e_{4}=\lambda e_{3}+p(-\gamma+\lambda)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2p\lambda e_{1}-2p\gamma e_{2},&e_{2}\cdot e_{2}=2p\gamma e_{1}-\frac{2}{p}(-p^{2}\lambda+(2p^{2}+1)\gamma)e_{2},\\[5.69054pt] e_{2}\cdot e_{3}=p(-\gamma+\lambda)e_{3}+\gamma e_{4},&e_{2}\cdot e_{4}=\gamma e_{3}+\frac{1}{p}((-1-p^{2})\gamma+p^{2}\lambda)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\frac{2}{p}(-p^{2}\gamma+(p^{2}+1)\lambda)e_{3}+2\lambda e_{4},&e_{3}\cdot e_{2}=-2p(\gamma-\lambda)e_{3}+2\gamma e_{4},\\[5.69054pt] e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=pe_{1}-pe_{2},\\[5.69054pt] e_{4}\cdot e_{1}=2\lambda e_{3}-2p(\gamma-\lambda)e_{4},&e_{4}\cdot e_{2}=2\gamma e_{3}-\frac{2}{p}((p^{2}+1)\gamma-p^{2}\lambda))e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=pe_{1}-pe_{2},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
   19. (19)
-    The Lie superalgebra $(C_{1}^{1}+A):$ A Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R}):$
+    The Lie superalgebra $(C_{1}^{1}+A):$ A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=\lambda e_{1}+\gamma e_{2},&e_{1}\cdot e_{2}=(1+\lambda)e_{2},&e_{1}\cdot e_{3}=(1+\lambda)e_{3},&e_{1}\cdot e_{4}=\lambda e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\lambda e_{2},&e_{3}\cdot e_{1}=\lambda e_{3},&e_{3}\cdot e_{4}=-\lambda e_{2},&e_{4}\cdot e_{1}=\lambda e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=(1+\lambda)e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
-    $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(\gamma-1)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=(\gamma-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(\gamma-1)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=(\gamma-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
   20. (20)
-    The Lie superalgebra $(C_{1/2}^{1}+A):$ A Novikov structure is given by:
+    The Lie superalgebra $(C_{1/2}^{1}+A):$ A Novikov structure is given by
     $\begin{array}[]{l}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\gamma e_{2},\quad e_{1}\cdot e_{2}=\frac{1}{2}e_{2},\quad e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\quad e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\quad e_{3}\cdot e_{3}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
-    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2}&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4}.\end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4}.\end{array}$
   21. (21)
-    The Lie superalgebra $(C_{-1}^{2}+A):$ A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R})$:
+    The Lie superalgebra $(C_{-1}^{2}+A):$ A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{1}\cdot e_{1}=(2\lambda-1)e_{1}+\gamma e_{2},\quad e_{1}\cdot e_{2}=(2\lambda-1)e_{2},\quad e_{1}\cdot e_{3}=2\lambda e_{3},\quad e_{1}\cdot e_{4}=2(\lambda-1)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(2\lambda-1)e_{2},\quad e_{3}\cdot e_{1}=(2\lambda-1)e_{3},\quad e_{3}\cdot e_{4}=(1-\lambda)e_{2},\quad e_{4}\cdot e_{1}=(2\lambda-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=\lambda e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
-    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{1}{2}(2+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(\gamma-2)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(2+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(\gamma-2)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
   22. (22)
-    The Lie superalgebra $(C^{3}+A):$ A Novikov structure is given by:
+    The Lie superalgebra $(C^{3}+A):$ A Novikov structure is given by
     $\begin{array}[]{c}e_{4}\cdot e_{1}=-e_{3},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
-    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3},&e_{1}\cdot e_{4}=e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{4}\cdot e_{1}=\gamma e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3},&e_{1}\cdot e_{4}=e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{4}\cdot e_{1}=\gamma e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
   23. (23)
-    The Lie superalgebra $(C_{0}^{5}+A):$ A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R})$:
+    The Lie superalgebra $(C_{0}^{5}+A):$ A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{1}\cdot e_{1}=2\lambda e_{1}+\gamma e_{2},\quad e_{1}\cdot e_{2}=2\lambda e_{2},\quad e_{1}\cdot e_{3}=2\lambda e_{3}-e_{4},\quad e_{1}\cdot e_{4}=e_{3}+2\lambda e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2\lambda e_{2},\quad e_{3}\cdot e_{1}=2\lambda e_{3},\quad e_{3}\cdot e_{3}=\frac{1}{2}e_{2},\quad e_{3}\cdot e_{4}=-\lambda e_{2},\quad e_{4}\cdot e_{1}=2\lambda e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=\lambda e_{2},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
-    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3}-e_{4},&e_{1}\cdot e_{4}=e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{4}=e_{2}.\end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3}-e_{4},&e_{1}\cdot e_{4}=e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{4}=e_{2}.\end{array}$
   24. (24)
-    The Lie superalgebra $D^{1}$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in\mathbb{R})$:
+    The Lie superalgebra $D^{1}$: A Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{2}\cdot e_{2}=\lambda e_{1}+\gamma e_{3},\quad e_{2}\cdot e_{3}=(1+\mu)e_{1},\quad e_{2}\cdot e_{4}=e_{4},\quad e_{3}\cdot e_{2}=\mu e_{1}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{2}\cdot e_{2}=\gamma e_{1}+\lambda e_{3},&e_{2}\cdot e_{4}=e_{4},&e_{3}\cdot e_{2}=-e_{1}.\end{array}$
   25. (25)
-    The Lie superalgebra $D^{2}_{q}$ with $q\not=0$: A Novikov structure is given by:
-[…]
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
-    $\begin{array}[]{llll}e_{3}\cdot e_{1}=-e_{1},&e_{3}\cdot e_{2}=-e_{1}-e_{2},&e_{3}\cdot e_{3}=\lambda e_{1}+\gamma e_{2}&e_{3}\cdot e_{4}=qe_{4}.\end{array}$
+    The Lie superalgebra $D^{2}_{q}$ with $q\not=0$: A Novikov structure is given by
+    $\begin{array}[]{l}e_{1}\cdot e_{3}=-qe_{1},\quad e_{2}\cdot e_{3}=-qe_{2},\quad e_{3}\cdot e_{1}=-(1+q)e_{1},\quad e_{3}\cdot e_{2}=-e_{1}-(q+1)e_{2},\\[5.69054pt] e_{3}\cdot e_{3}=-qe_{3},\quad e_{4}\cdot e_{3}=-qe_{4}.\end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{llll}e_{3}\cdot e_{1}=-e_{1},&e_{3}\cdot e_{2}=-e_{1}-e_{2},&e_{3}\cdot e_{3}=\lambda e_{1}+\gamma e_{2},&e_{3}\cdot e_{4}=qe_{4}.\end{array}$
   26. (26)
-    The Lie superalgebra $D^{3}_{pq}$ with $pq\not=0$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in{\mathbb{R}}$:
+    The Lie superalgebra $D^{3}_{pq}$ with $pq\not=0$: A Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{1}\cdot e_{3}=(p+\gamma)e_{1},\quad e_{2}\cdot e_{3}=(p+\gamma)e_{2},\quad e_{3}\cdot e_{1}=\gamma e_{1}+e_{2},\quad e_{3}\cdot e_{2}=-e_{1}+\gamma e_{2},\\[5.69054pt] e_{3}\cdot e_{3}=\lambda e_{1}+\mu e_{2}+(p+\gamma)e_{3},\quad e_{3}\cdot e_{4}=(p+q+\gamma)e_{4},\quad e_{4}\cdot e_{3}=(p+\gamma)e_{4}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
-    $\begin{array}[]{llll}e_{3}\cdot e_{1}=-pe_{1}+e_{2},&e_{3}\cdot e_{2}=-e_{1}-pe_{2},&e_{3}\cdot e_{3}=\lambda e_{1}+\gamma e_{2}&e_{3}\cdot e_{4}=qe_{4}.\end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{llll}e_{3}\cdot e_{1}=-pe_{1}+e_{2},&e_{3}\cdot e_{2}=-e_{1}-pe_{2},&e_{3}\cdot e_{3}=\lambda e_{1}+\gamma e_{2},&e_{3}\cdot e_{4}=qe_{4}.\end{array}$
   27. (27)
-    The Lie superalgebra $D^{11}_{pq}$ with $0<|p|\leq|q|\leq 1$: A Novikov structure is given by:
+    The Lie superalgebra $D^{11}_{pq}$ with $0<|p|\leq|q|\leq 1$: A Novikov structure is given by
     $\begin{array}[]{l}e_{1}\cdot e_{2}=e_{2},\quad e_{1}\cdot e_{3}=pe_{3},\quad e_{1}\cdot e_{4}=qe_{4}.\end{array}$
-    A Balinsky-Novikov structure is given by:
+    A Balinsky-Novikov structure is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{2}=e_{2},&e_{1}\cdot e_{3}=pe_{3},&e_{1}\cdot e_{4}=qe_{4}.\end{array}$
   28. (28)
-    The Lie superalgebra $D^{12}$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
+    The Lie superalgebra $D^{12}$: A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{1}\cdot e_{2}=e_{2},\quad e_{1}\cdot e_{4}=(1+\gamma)e_{3},\quad e_{4}\cdot e_{1}=\gamma e_{3}.\end{array}$
-    A Balinsky-Novikov structure is given by:
+    A Balinsky-Novikov structure is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=-2e_{1},&e_{1}\cdot e_{3}=-e_{3},&e_{1}\cdot e_{4}=e_{3}-e_{4},&e_{2}\cdot e_{1}=-2e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-2e_{3},&e_{4}\cdot e_{1}=-2e_{4}.\end{array}$
   29. (29)
-    The Lie superalgebra $D^{13}_{p}$ with $p\not=0$: A Novikov structure is given by:
+    The Lie superalgebra $D^{13}_{p}$ with $p\not=0$: A Novikov structure is given by
     $\begin{array}[]{l}e_{1}\cdot e_{2}=pe_{2},\quad e_{1}\cdot e_{3}=e_{3},\quad e_{1}\cdot e_{4}=e_{3}+e_{4}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
+    A Balinsky-Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=2(\gamma-1)e_{1},&e_{1}\cdot e_{2}=(p-1+\gamma)e_{2},&e_{1}\cdot e_{3}=\gamma e_{3},&e_{1}\cdot e_{4}=e_{3}+\gamma e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2(\gamma-1)e_{2},&e_{3}\cdot e_{1}=2(\gamma-1)e_{3},&e_{4}\cdot e_{1}=2(\gamma-1)e_{4}.\end{array}$
   30. (30)
-    The Lie superalgebra $D^{14}_{pq}$ with $p\not=0,q\geq 0$: A Novikov structure is given by:
+    The Lie superalgebra $D^{14}_{pq}$ with $p\not=0,q\geq 0$: A Novikov structure is given by
     $\begin{array}[]{l}e_{1}\cdot e_{2}=pe_{2},\quad e_{1}\cdot e_{3}=qe_{3}-e_{4},\quad e_{1}\cdot e_{4}=e_{3}+qe_{4}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
+    A Balinsky-Novikov structure is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=-2qe_{1},&e_{1}\cdot e_{2}=(p-q)e_{2},&e_{1}\cdot e_{3}=-e_{4},&e_{1}\cdot e_{4}=e_{3},\\[5.69054pt] e_{2}\cdot e_{1}=-2qe_{2},&e_{3}\cdot e_{1}=-2qe_{3},&e_{4}\cdot e_{1}=-2qe_{4}.\end{array}$
   31. (31)
-    The Lie superalgebra $D^{15}$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
+    The Lie superalgebra $D^{15}$: A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{1}\cdot e_{3}=e_{2},\quad e_{1}\cdot e_{4}=\gamma e_{2}+e_{3},\quad e_{4}\cdot e_{1}=\gamma e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
+    A Balinsky-Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=2\gamma e_{1},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=e_{2}+\gamma e_{3},&e_{1}\cdot e_{4}=e_{3}+\gamma e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2\gamma e_{2},&e_{3}\cdot e_{1}=2\gamma e_{3},&e_{4}\cdot e_{1}=2\gamma e_{4}.\end{array}$
   32. (32)
-    The Lie superalgebra $D^{16}$: A Novikov structure is given by:
+    The Lie superalgebra $D^{16}$: A Novikov structure is given by
     $\begin{array}[]{l}e_{1}\cdot e_{2}=e_{2},\quad e_{1}\cdot e_{3}=e_{2}+e_{3},\quad e_{1}\cdot e_{4}=e_{3}+e_{4}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R})$:
+    A Balinsky-Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{llll}e_{1}\cdot e_{1}=2(\gamma-1)e_{1},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=e_{2}+\gamma e_{3},&e_{1}\cdot e_{4}=e_{3}+\gamma e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2(\gamma-1)e_{2},&e_{3}\cdot e_{1}=2(\gamma-1)e_{3},&e_{4}\cdot e_{1}=2(\gamma-1)e_{4}.\end{array}$
   33. (33)
-    The Lie superalgebra $(A_{3,1}+A)$: A Novikov structure is given by (where $\lambda,\gamma,\mu\in\mathbb{R}$):
+    The Lie superalgebra $(A_{3,1}+A)$: A Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{2}\cdot e_{2}=\gamma e_{1},\quad e_{2}\cdot e_{3}=(1+\lambda)e_{1},\quad e_{3}\cdot e_{2}=\lambda e_{1},\quad e_{3}\cdot e_{3}=\mu e_{1},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{1}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda,\mu\in\mathbb{R})$:
-    $\begin{array}[]{lllll}e_{2}\cdot e_{2}=\gamma e_{1},&e_{2}\cdot e_{3}=\lambda e_{1},&e_{3}\cdot e_{2}=(\lambda-1)e_{1},&e_{3}\cdot e_{3}=\mu e_{1}&e_{4}\cdot e_{4}=e_{1}.\end{array}$
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
+    $\begin{array}[]{lllll}e_{2}\cdot e_{2}=\gamma e_{1},&e_{2}\cdot e_{3}=\lambda e_{1},&e_{3}\cdot e_{2}=(\lambda-1)e_{1},&e_{3}\cdot e_{3}=\mu e_{1},&e_{4}\cdot e_{4}=e_{1}.\end{array}$
   34. (34)
-    The Lie superalgebra $(D^{3}_{p,-1/2})$ with $p\not=0$: A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
+    The Lie superalgebra $(D^{3}_{p,-1/2})$ with $p\not=0$: A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2}+\gamma e_{3},\quad e_{1}\cdot e_{2}=\frac{1}{2}e_{2},\quad e_{1}\cdot e_{3}=\frac{1}{2}(2p-1)e_{3},\quad e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\quad e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{lllll}e_{1}\cdot e_{1}=\gamma e_{2}+\lambda e_{3},&e_{1}\cdot e_{2}=e_{2},&e_{1}\cdot e_{3}=pe_{3},&e_{1}\cdot e_{4}=\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
   35. (35)
-    The Lie superalgebra $(D^{2}_{-1/2})^{1}$: A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
+    The Lie superalgebra $(D^{2}_{-1/2})^{1}$: A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2}+\gamma e_{3},\quad e_{1}\cdot e_{2}=\frac{1}{2}e_{2},\quad e_{1}\cdot e_{3}=e_{2}+\frac{1}{2}e_{3},\quad e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\quad e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{lllll}e_{1}\cdot e_{1}=\gamma e_{2}+\lambda e_{3},&e_{1}\cdot e_{2}=e_{2},&e_{1}\cdot e_{3}=e_{2}+e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
   36. (36)
-    The Lie superalgebra $(D^{2}_{-1/2})^{2}$: A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
+    The Lie superalgebra $(D^{2}_{-1/2})^{2}$: A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{1}\cdot e_{1}=-\frac{1}{2}e_{1}+\lambda e_{2}+\gamma e_{3},\quad e_{1}\cdot e_{2}=\frac{1}{2}e_{2},\quad e_{1}\cdot e_{3}=-e_{2}+\frac{1}{2}e_{3},\quad e_{2}\cdot e_{1}=-\frac{1}{2}e_{2},\\[5.69054pt] e_{3}\cdot e_{1}=-\frac{1}{2}e_{3},\quad e_{4}\cdot e_{1}=-\frac{1}{2}e_{4},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
-    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
+    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{lllll}e_{1}\cdot e_{1}=\gamma e_{2}+\lambda e_{3},&e_{1}\cdot e_{2}=e_{2},&e_{1}\cdot e_{3}=-e_{2}+e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
   37. (37)
-    The Lie superalgebra $(A_{1,1}+3A)^{1}$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in\mathbb{R}$):
+    The Lie superalgebra $(A_{1,1}+3A)^{1}$: A Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{2}\cdot e_{2}=\frac{1}{2}e_{1},\quad e_{2}\cdot e_{3}=\lambda e_{1},\quad e_{2}\cdot e_{4}=\gamma e_{1},\quad e_{3}\cdot e_{2}=-\lambda e_{1},\quad e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\\[5.69054pt] e_{3}\cdot e_{4}=\mu e_{1},\quad e_{4}\cdot e_{2}=-\gamma e_{1},\quad e_{4}\cdot e_{3}=-\mu e_{1},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{1}.\end{array}$
@@
   38. (38)
-    The Lie superalgebra $(A_{1,1}+3A)^{2}$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in\mathbb{R}$):
+    The Lie superalgebra $(A_{1,1}+3A)^{2}$: A Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
     $\begin{array}[]{l}e_{2}\cdot e_{2}=\frac{1}{2}e_{1},\quad e_{2}\cdot e_{3}=\lambda e_{1},\quad e_{2}\cdot e_{4}=\gamma e_{1},\quad e_{3}\cdot e_{2}=-\lambda e_{1},\quad e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\\[5.69054pt] e_{3}\cdot e_{4}=\mu e_{1},\quad e_{4}\cdot e_{2}=-\gamma e_{1},\quad e_{4}\cdot e_{3}=-\mu e_{1},\quad e_{4}\cdot e_{4}=-\frac{1}{2}e_{1}.\end{array}$
-    A Balinsky-Novikov structure is given by:
+    A Balinsky-Novikov structure is given by
     $\begin{array}[]{llll}e_{2}\cdot e_{2}=e_{1},&e_{3}\cdot e_{3}=e_{1},&e_{4}\cdot e_{4}=-e_{1}.\end{array}$
+We conclude by saying that the lists of Lie superalgebras classified by Backhouse are all left-symmetric.
+Moreover, they are all Novikov and Balinsky-Novikov superalgebras, except the two $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$ that are Balinsky-Novikov but not Novikov.
 6. The list of four-dimensional real Lie superalgebras
-The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3},e_{4}$ Symplectic structure $\omega$ $p(\omega)$
-$D^{5}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{4},\\[2.84526pt] [e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
-$D^{6}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{4},\\[2.84526pt] [e_{2},e_{3}]&=&-e_{4},\\[2.84526pt] [e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
-$\begin{array}[]{c}D^{7}_{pq},\\[2.84526pt] pq\not=0,\\[2.84526pt] p\geq q\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ None
-$\begin{array}[]{c}D^{7}_{-1q},\\[2.84526pt] q\leq-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&-e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ $\begin{array}[]{l}e_{4}^{*}\wedge e_{1}^{*}+e_{3}^{*}\wedge e_{2}^{*}\end{array}$ ${\bar{1}}$
-$\begin{array}[]{c}D^{7}_{pp},\\[5.69054pt] p=-q\\[5.69054pt] q\not=0,-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&pe_{3}\\[2.84526pt] [e_{1},e_{4}]&=&-pe_{4}\end{array}$ $\begin{array}[]{l}e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\end{array}$ $\bar{0}$
-$\begin{array}[]{c}D^{7}_{pq},\\[2.84526pt] p=-q,\\[2.84526pt] q=-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&-e_{4}\end{array}$ $\begin{array}[]{l}e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\begin{array}[]{l}\bar{0}\\[2.84526pt] \bar{1}\end{array}$
-$\begin{array}[]{c}D^{8}_{p},\\ p\not=0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+pe_{4}\end{array}$ None
-$D^{8}_{-1}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&-e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}-e_{4}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
-$\begin{array}[]{c}D^{9}_{pq}\\[2.84526pt] q>0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&pe_{3}-qe_{4},\\[2.84526pt] [e_{1},e_{2}]&=&e_{2},\\[5.69054pt] [e_{1},e_{4}]&=&qe_{3}+pe_{4}\end{array}$ None
-$D^{10}_{q}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&(q+1)e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4},\\[2.84526pt] [e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}(1+q)e_{1}^{*}\wedge e_{3}^{*}+e_{2}^{*}\wedge e_{4}^{*},\\[5.69054pt] \text{where }q\not=-1\end{array}$ $\bar{1}$
-Table 1. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=2|2$
-The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3},e_{4}$ Symplectic structure $p(\omega)$
+The tables below differ slightly from those in [BM].
+The symplectic forms given in [BM] are given in terms of parameters.
+For convenience, we specify these parameters here.
+In addition, we correct the statement regarding the Lie superalgebras $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$.
+The two admit both an odd and an even non-degenerate closed form, contrary to what is stated in [BM].
+The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3},e_{4}$ Symplectic structure $\omega$ $|\omega|$
+$D^{5}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&e_{4},\\[2.84526pt] [e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
+$D^{6}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{3},\;[e_{1},e_{4}]&=&e_{4},\\[2.84526pt] [e_{2},e_{3}]&=&-e_{4},\;[e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
+$\begin{array}[]{c}D^{7}_{pq},\\[2.84526pt] pq\not=0,\\[2.84526pt] p\geq q\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ None
+$\begin{array}[]{c}D^{7}_{-1q},\\[2.84526pt] q\leq-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ $\begin{array}[]{l}e_{4}^{*}\wedge e_{1}^{*}+e_{3}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$}
+$\begin{array}[]{c}D^{7}_{pp},\\[5.69054pt] p=-q\\[5.69054pt] q\not=0,-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3}\\[2.84526pt] [e_{1},e_{4}]&=&-pe_{4}\end{array}$ $\begin{array}[]{l}e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\end{array}$ $\bar{0}$
+$\begin{array}[]{c}D^{7}_{pq},\\[2.84526pt] p=-q,\\[2.84526pt] q=-1\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&-e_{4}\end{array}$ $\begin{array}[]{l}e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\begin{array}[]{l}\bar{0}\\[2.84526pt] \bar{1}\end{array}$
+$\begin{array}[]{c}D^{8}_{p},\\ p\not=0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+pe_{4}\end{array}$ None
+$D^{8}_{-1}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}-e_{4}\end{array}$ $\begin{array}[]{l}e_{3}^{*}\wedge e_{1}^{*}+e_{4}^{*}\wedge e_{2}^{*}\end{array}$ $\bar{1}$
+$\begin{array}[]{c}D^{9}_{pq}\\[2.84526pt] q>0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&pe_{3}-qe_{4},&[e_{1},e_{2}]&=&e_{2},\\[5.69054pt] [e_{1},e_{4}]&=&qe_{3}+pe_{4}\end{array}$ None
+$D^{10}_{q}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&(q+1)e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4},&[e_{2},e_{4}]&=&e_{3}\end{array}$ $\begin{array}[]{l}(1+q)e_{1}^{*}\wedge e_{3}^{*}+e_{2}^{*}\wedge e_{4}^{*},\\[5.69054pt] \text{where }q\not=-1\end{array}$ $\bar{1}$
+Table 6. Trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=2|2$
+The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3},e_{4}$ Symplectic structure $|\omega|$
+$(D^{7}_{1/2\;1/2})^{1}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{3},e_{3}]&=&e_{2}\\[2.84526pt] [e_{4},e_{4}]&=&e_{2}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}-\frac{1}{2}e_{3}^{*}\wedge e_{3}^{*}-\frac{1}{2}e_{4}^{*}\wedge e_{4}^{*}$ $\bar{0}$
+$(D^{7}_{1/2\;1/2})^{2}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{3},e_{3}]&=&e_{2}\\[2.84526pt] [e_{4},e_{4}]&=&-e_{2}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}+\frac{1}{2}e_{3}^{*}\wedge e_{3}^{*}-\frac{1}{2}e_{4}^{*}\wedge e_{4}^{*}$ $\bar{0}$
+$(D^{7}_{1/2\;1/2})^{3}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{3},e_{3}]&=&e_{2}\end{array}$ None -
+$\begin{array}[]{c}(D^{7}_{1-p,p})\\[2.84526pt] p\leq\frac{1}{2}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3}\\[2.84526pt] [e_{1},e_{4}]&=&(1-p)e_{4},&[e_{3},e_{4}]&=&e_{2}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}-e_{3}^{*}\wedge e_{4}^{*}$ $\bar{0}$
+$(D^{8}_{1/2})$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None -
+$\begin{array}[]{c}(D^{9}_{1/2,p})\\[2.84526pt] p>0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{3},e_{3}]&=&e_{2}\\[2.84526pt] [e_{1},e_{4}]&=&pe_{3}+\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{1},e_{3}]&=&\frac{1}{2}e_{3}-pe_{4}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}-\frac{1}{2}e_{3}^{*}\wedge e_{3}^{*}-\frac{1}{2}e_{4}^{*}\wedge e_{4}^{*}$ $\bar{0}$}
+$(D^{10}_{0})^{1}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{3}\\[2.84526pt] [e_{2},e_{4}]&=&e_{3},&[e_{4},e_{4}]&=&e_{1}\\[2.84526pt] [e_{3},e_{4}]&=&-\frac{1}{2}e_{2}\end{array}$ $\begin{array}[]{l}2e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] e_{1}^{*}\wedge e_{3}^{*}+e_{2}^{*}\wedge e_{4}^{*}\end{array}$ $\begin{array}[]{l}{\bar{0}}\\[2.84526pt] {\bar{1}}\end{array}$
+$(D^{10}_{0})^{2}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{3}\\[2.84526pt] [e_{2},e_{4}]&=&e_{3},&[e_{4},e_{4}]&=&-e_{1}\\[2.84526pt] [e_{3},e_{4}]&=&\frac{1}{2}e_{2}\end{array}$ $\begin{array}[]{l}2e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] e_{1}^{*}\wedge e_{3}^{*}+e_{2}^{*}\wedge e_{4}^{*}\end{array}$ $\begin{array}[]{l}{\bar{0}}\\[2.84526pt] {\bar{1}}\end{array}$
+$(2A_{1,1}+2A)^{1}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None -
 […]
@@
 […]
-$\begin{array}[]{c}(D^{7}_{1-p,p})\\[2.84526pt] p\leq\frac{1}{2}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3}\\[2.84526pt] [e_{1},e_{4}]&=&(1-p)e_{4},\\[2.84526pt] [e_{3},e_{4}]&=&e_{2}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}-e_{3}^{*}\wedge e_{4}^{*}$ $\bar{0}$
-$(D^{8}_{1/2})$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+\frac{1}{2}e_{4},\\[2.84526pt] [e_{4},e_{4}]&=&e_{2}\end{array}$ None --
-$\begin{array}[]{c}(D^{9}_{1/2,p})\\[2.84526pt] p>0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{3},e_{3}]&=&e_{2}\\[2.84526pt] [e_{1},e_{4}]&=&pe_{3}+\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{1},e_{3}]&=&\frac{1}{2}e_{3}-pe_{4}\end{array}$ $e_{1}^{*}\wedge e_{2}^{*}-\frac{1}{2}e_{3}^{*}\wedge e_{3}^{*}-\frac{1}{2}e_{4}^{*}\wedge e_{4}^{*}$ ${\bar{0}}$
-$(D^{10}_{0})^{1}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{3}\\[2.84526pt] [e_{2},e_{4}]&=&e_{3},&[e_{4},e_{4}]&=&e_{1}\\[2.84526pt] [e_{3},e_{4}]&=&-\frac{1}{2}e_{2}\end{array}$ $\begin{array}[]{l}2e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] e_{1}^{*}\wedge e_{3}^{*}+e_{2}^{*}\wedge e_{4}^{*}\end{array}$ $\begin{array}[]{l}{\bar{0}}\\[2.84526pt] {\bar{1}}\end{array}$
-$(D^{10}_{0})^{2}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{3}\\[2.84526pt] [e_{2},e_{4}]&=&e_{3},&[e_{4},e_{4}]&=&-e_{1}\\[2.84526pt] [e_{3},e_{4}]&=&\frac{1}{2}e_{2}\end{array}$ $\begin{array}[]{l}2e_{2}^{*}\wedge e_{1}^{*}-e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] e_{1}^{*}\wedge e_{3}^{*}+e_{2}^{*}\wedge e_{4}^{*}\end{array}$ $\begin{array}[]{l}{\bar{0}}\\[2.84526pt] {\bar{1}}\end{array}$
-$(2A_{1,1}+2A)^{1}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None --
-$(2A_{1,1}+2A)^{2}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{3},e_{4}]&=&e_{1}\\ \end{array}$ None --
-$\begin{array}[]{c}(2A_{1,1}+2A)^{3}_{p}\\ p>0\end{array}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{3},e_{4}]&=&p(e_{1}+e_{2})\\ \end{array}$ For p12: NoneFor $\begin{array}[]{l}\text{For $p\not=\frac{1}{2}$: None}\\[2.84526pt] \text{For $p=\frac{1}{2}:\,e_{2}^{*}\wedge e_{3}^{*}-e_{1}^{*}\wedge e_{4}^{*}$}\end{array}$ $\ \bar{1}\end{array}$
-$\begin{array}[]{c}(2A_{1,1}+2A)^{4}_{p}\\ p>0\end{array}$ $\begin{array}[]{lcllcl}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] [e_{3},e_{4}]&=&p(e_{1}-e_{2})\\ \end{array}$ None --
-$(C^{1}_{1}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{3}\\[2.84526pt] [e_{3},e_{4}]&=&e_{2}\\ \end{array}$ $\frac{1}{2}e_{2}^{*}\wedge e_{1}^{*}+\frac{1}{2}e_{3}^{*}\wedge e_{4}^{*}+\frac{1}{4}e_{4}^{*}\wedge e_{4}^{*}$ $\bar{0}$
-$(C^{1}_{1/2}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3}\\[2.84526pt] [e_{3},e_{3}]&=&e_{2}\\ \end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}-\frac{\lambda}{2}e_{3}^{*}\wedge e_{3}^{*}-\frac{\mu}{2}e_{4}^{*}\wedge e_{4}^{*},\\ \text{where }\lambda\mu\not=0\end{array}$ $\bar{0}$
-$(C^{2}_{-1}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&-e_{4}\\[2.84526pt] [e_{3},e_{4}]&=&e_{2}\\ \end{array}$ None --
+[…]
+[…]
+[…]
 $(C^{3}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{4}]&=&e_{3},&[e_{4},e_{4}]&=&e_{2}\end{array}$ $2e_{1}^{*}\wedge e_{2}^{*}-e_{3}^{*}\wedge e_{4}^{*}$ $\bar{0}$
-$(C^{5}_{0}+A)$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&-e_{4},&[e_{1},e_{4}]&=&e_{3}\\[2.84526pt] [e_{3},e_{3}]&=&e_{2}&[e_{4},e_{4}]&=&e_{2}\\ \end{array}$ None --
-Table 2. Non-trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=2|2$
-The LSA Relations in the basis: $e_{1},e_{2},e_{3}\,|\,e_{4}$ Symplectic structure $p(\omega)$
+[…]
+Table 7. Non-trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=2|2$
+The LSA Relations in the basis: $e_{1},e_{2},e_{3}\,|\,e_{4}$ Symplectic structure $|\omega|$
 $D^{1}$ $\begin{array}[]{lcrlcl}[e_{2},e_{3}]&=&e_{1},&[e_{2},e_{4}]&=&e_{4}\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{3}^{*}+\mu e_{2}\wedge e_{4}^{*}\\[2.84526pt] +\nu e_{1}^{*}\wedge e_{2}^{*}+\gamma e_{2}\wedge e_{3}^{*}\\[2.84526pt] \text{where }\lambda\mu\not=0\end{array}$ NH
-$\begin{array}[]{l}D^{2}_{q}\\[2.84526pt] q\not=-1,0\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{1},&[e_{2},e_{3}]&=&e_{1}+e_{2},\\[2.84526pt] [e_{3},e_{4}]&=&qe_{4}\end{array}$ None --
+$\begin{array}[]{l}D^{2}_{q}\\[2.84526pt] q\not=-1,0\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{1},&[e_{2},e_{3}]&=&e_{1}+e_{2},\\[2.84526pt] [e_{3},e_{4}]&=&qe_{4}\end{array}$ None -
 $\begin{array}[]{l}D^{2}_{-1}\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{1},&[e_{2},e_{3}]&=&e_{1}+e_{2},\\[2.84526pt] [e_{3},e_{4}]&=&-e_{4}\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{3}^{*}+\mu e_{2}^{*}\wedge e_{3}^{*}+\\[5.69054pt] \nu e_{2}^{*}\wedge e_{4}^{*}+\gamma e_{3}^{*}\wedge e_{4}^{*}\\[5.69054pt] \text{where }\lambda\nu\not=0\end{array}$ NH
-$\begin{array}[]{l}D^{3}_{pq}\\[2.84526pt] pq\not=0\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&pe_{1}-e_{2},&[e_{2},e_{3}]&=&e_{1}+pe_{2},\\[2.84526pt] [e_{3},e_{4}]&=&qe_{4}\end{array}$ None --
-Table 3. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=3|1$
-The LSA Relations in the basis: $e_{1}\,|\,e_{2},e_{3},e_{4}$ Symplectic structure $p(\omega)$
-$\begin{array}[]{c}D^{11}_{pq}\\ 0<|p|\leq|q|\leq 1\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ None --
+$\begin{array}[]{l}D^{3}_{pq}\\[2.84526pt] pq\not=0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{3}]&=&pe_{1}-e_{2},&[e_{2},e_{3}]&=&e_{1}+pe_{2},\\[2.84526pt] [e_{3},e_{4}]&=&qe_{4}\end{array}$ None -
+Table 8. Trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=3|1$
+The LSA Relations in the basis: $e_{1}\,|\,e_{2},e_{3},e_{4}$ Symplectic structure $|\omega|$
+$\begin{array}[]{c}D^{11}_{pq}\\ 0<|p|\leq|q|\leq 1\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ None -
 $\begin{array}[]{c}D^{11}_{pq}\\ (p,q)=(-1,-1)\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}+\nu e_{1}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\gamma e_{2}^{*}\wedge e_{3}^{*}+\delta e_{2}^{*}\wedge e_{4}^{*},\\[2.84526pt] \text{where }\nu\gamma-\mu\delta\not=0\end{array}$ NH
@@
 $\begin{array}[]{c}D^{11}_{pq}\\ 0<|p|<1\\ q=-1\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}+\nu e_{1}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\gamma e_{2}^{*}\wedge e_{4}^{*},\text{ where }\mu\gamma\not=0\end{array}$ NH
-$D^{12}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{4}]&=&e_{3},\end{array}$ None --
-$\begin{array}[]{c}D^{13}_{p}\\[2.84526pt] p\not=0\text{ (generic)}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&pe_{2},&[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ None --
+$D^{12}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{4}]&=&e_{3},\end{array}$ None -
+$\begin{array}[]{c}D^{13}_{p}\\[2.84526pt] p\not=0\text{ (generic)}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&pe_{2},&[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ None -
 $\begin{array}[]{c}D^{13}_{-1}\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&-e_{2},&[e_{1},e_{3}]&=&e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}+\gamma e_{1}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\nu e_{2}^{*}\wedge e_{4}^{*},\text{ where }\mu\nu\not=0\end{array}$ NH
-$\begin{array}[]{c}D^{14}_{pq}\\[2.84526pt] p\not=0,q\geq 0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&pe_{2},\\[2.84526pt] [e_{1},e_{3}]&=&qe_{3}-e_{4},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}+e_{3}\end{array}$ None --
+$\begin{array}[]{c}D^{14}_{pq}\\[2.84526pt] p\not=0,q\geq 0\end{array}$ $\begin{array}[]{lcllcl}[e_{1},e_{2}]&=&pe_{2},\\[2.84526pt] [e_{1},e_{3}]&=&qe_{3}-e_{4},\\[2.84526pt] [e_{1},e_{4}]&=&qe_{4}+e_{3}\end{array}$ None -
 $D^{15}$ $\begin{array}[]{lcrlcl}[e_{1},e_{3}]&=&e_{2},&[e_{1},e_{4}]&=&e_{3},\end{array}$ $\begin{array}[]{l}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}+\nu e_{1}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\delta e_{2}^{*}\wedge e_{4}^{*}-\frac{1}{2}\delta e_{3}^{*}\wedge e_{3}^{*},\\[2.84526pt] \text{where }\delta^{2}(\mu^{2}-2\lambda\nu)\not=0.\end{array}$ NH
-$D^{16}$ $\begin{array}[]{lcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ None --
-Table 4. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=1|3$
-The LSA Relations in the basis: $e_{1},e_{2},e_{3}\,|\,e_{4}$ Symplectic structure $p(\omega)$
-$(A_{3,1}+A)$ $\begin{array}[]{lcrlcl}[e_{2},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{1}\end{array}$ None --
-$\begin{array}[]{l}(D^{3}_{p,-1/2})\\[2.84526pt] p\not=0\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] \end{array}$ None --
-$(D^{3}_{-1/2,-1/2})$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-\frac{1}{2}e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] \end{array}$ $\begin{array}[]{l}\frac{1}{2}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}\\[2.84526pt] +\nu e_{1}^{*}\wedge e_{4}^{*}+\gamma e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\frac{1}{2}\lambda e_{4}^{*}\wedge e_{4}^{*},\text{ where }\gamma\lambda\not=0\end{array}$ NH
-$(D^{2}_{-1/2})^{1}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None --
-$(D^{2}_{-1/2})^{2}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None --
-Table 5. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=3|1$
-The LSA Relations in the basis: $e_{1}\,|\,e_{2},e_{3},e_{4}$ Symplectic structure $p(\omega)$
-$(A_{1,1}+3A)^{1}$ $\begin{array}[]{lcrlcl}[e_{2},e_{2}]&=&e_{1},&[e_{3},e_{3}]&=&e_{1},\\[2.84526pt] [e_{4},e_{4}]&=&e_{1}\end{array}$ None --
-$(A_{1,1}+3A)^{2}$ $\begin{array}[]{lcrlcl}[e_{2},e_{2}]&=&e_{1},&[e_{3},e_{3}]&=&e_{1},\\[2.84526pt] [e_{4},e_{4}]&=&-e_{1}\end{array}$ None --
-Table 6. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{h}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=1|3$
+$D^{16}$ $\begin{array}[]{lcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ None -
+Table 9. Trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=1|3$
+The LSA Relations in the basis: $e_{1},e_{2},e_{3}\,|\,e_{4}$ Symplectic structure $|\omega|$
+$(A_{3,1}+A)$ $\begin{array}[]{lcrlcl}[e_{2},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{1}\end{array}$ None -
+[…]
+[…]
+$(D^{2}_{-1/2})^{1}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None -
+$(D^{2}_{-1/2})^{2}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None -
+Table 10. Trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=3|1$
+The LSA Relations in the basis: $e_{1}\,|\,e_{2},e_{3},e_{4}$ Symplectic structure $|\omega|$
+$(A_{1,1}+3A)^{1}$ $\begin{array}[]{lcrlcl}[e_{2},e_{2}]&=&e_{1},&[e_{3},e_{3}]&=&e_{1},\\[2.84526pt] [e_{4},e_{4}]&=&e_{1}\end{array}$ None -
+$(A_{1,1}+3A)^{2}$ $\begin{array}[]{lcrlcl}[e_{2},e_{2}]&=&e_{1},&[e_{3},e_{3}]&=&e_{1},\\[2.84526pt] [e_{4},e_{4}]&=&-e_{1}\end{array}$ None -
+Table 11. Trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{h}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=1|3$
+7. Appendix: The decomposable case
+We are only considering Lie superalgebras that are not Lie algebras.
+We investigate symplectic structures on the decomposable 4-dimensional real Lie superalgebras below.
+The notation we use is that of Backhouse [B].
+Over $A\oplus B$, the notation $\omega:=\omega_{A}\oplus\omega_{B}$ means that $\omega(a_{1}+b_{1},a_{2}+b_{2})=\omega_{A}(a_{1},a_{2})+\omega_{B}(b_{1},b_{2})$ for all $a_{1},a_{2}\in A$ and $b_{1},b_{2}\in B$.
+7.1. Abelian LSA
+  1. (1)
+    NH: $\mathbb{R}^{1|3},$ $\mathbb{R}^{3|1},$
+  2. (2)
+    Symplectic: $(\mathbb{R}^{2|2},\omega_{\mathbb{R}^{1|1}}\oplus\omega_{\mathbb{R}^{1|1}})$, $(\mathbb{R}^{2|2},\omega_{\mathbb{R}^{2|0}}\oplus\omega_{\mathbb{R}^{0|2}})$, $(\mathbb{R}^{0|4},\omega_{\mathbb{R}^{0|2}}\oplus\omega_{\mathbb{R}^{0|2}}).$
+7.2. LSA of the form $L\oplus M$, where $L,M$ are 2-dimensional
+We refer to Table 12 for the description of the LSA below:
+  1. (1)
+    NH: $B\oplus\mathbb{R}^{2|0},$ $B\oplus\mathbb{R}^{0|2},$ $B\oplus\mathrm{Al},$ $\mathrm{Al}\oplus\mathbb{R}^{1|1},$
+  2. (2)
+    None: $(A_{1,2}+A)\oplus\mathbb{R}^{2|0}$, $(A_{1,2}+A)\oplus\mathbb{R}^{0|2},$ $(A_{1,2}+A)\oplus\mathbb{R}^{1|1}$, $(A_{1,2}+A)\oplus\mathrm{Al}$, $(A_{1,2}+A)\oplus B$, $(A_{1,2}+A)\oplus(A_{1,2}+A)$,
+  3. (3)
+    Symplectic: $(B\oplus B,\omega_{B}\oplus\omega_{B})$, $(B\oplus\mathbb{R}^{1|1},\omega_{B}\oplus\omega_{\mathbb{R}^{1|1}})$, $(\mathrm{Al}\oplus\mathbb{R}^{0|2},\omega_{\mathrm{Al}}\oplus\omega_{\mathbb{R}^{0|2}})$.
+The LA Relations in the basis: $e_{1},e_{2}$ Symplectic structure $|\omega|$
+$\mathrm{Al}$ $\begin{array}[]{lcr}[e_{1},e_{2}]&=&e_{1}\end{array}$ $\omega_{\mathrm{Al}}:=e_{1}^{*}\wedge e_{2}^{*}$ $\overline{0}$
+$\mathbb{R}^{2|0}$ $\omega_{\mathbb{R}^{2|0}}:=e_{1}^{*}\wedge e_{2}^{*}$ $\overline{0}$
+The LSA Relations in the basis: $e_{1}\,|\,e_{2}$ Symplectic structure $|\omega|$
+$B$ $\begin{array}[]{lcr}[e_{1},e_{2}]&=&e_{2}\end{array}$ $\omega_{B}:=e_{1}^{*}\wedge e_{2}^{*}$ $\overline{1}$
+$(A_{1,2}+A)$ $\begin{array}[]{lcr}[e_{2},e_{2}]&=&e_{1}\end{array}$ None -
+$\mathbb{R}^{1|1}$ $\omega_{\mathbb{R}^{1|1}}:=e_{1}^{*}\wedge e_{2}^{*}$ $\overline{1}$
+The LSA Relations in the basis: $|\,e_{1},e_{2}$ Symplectic structure $|\omega|$
+$\mathbb{R}^{0|2}$ $\omega_{\mathbb{R}^{0|2}}:=e_{1}^{*}\wedge e_{2}^{*}$ $\overline{0}$
+Table 12. Quasi-Frobenuis structure
+7.3. LSA of the form $L\oplus\mathbb{R}^{0|1}$ or $L\oplus\mathbb{R}^{1|0}$, where $L$ is 3-dimensional
+Below, $X$ represents the generator of either $\mathbb{R}^{1|0}$ or $\mathbb{R}^{0|1}$. The parity of $X$ should be understood from the context.
+The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3}$ and $X$ Symplectic structure $|\omega|$
+$C^{1}_{p}\oplus\mathbb{R}^{1|0},p\not=0$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3}\end{array}$ None -
+$C^{1}_{p}\oplus\mathbb{R}^{0|1},p\not=0$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3}\end{array}$ None -
+$C^{1}_{-1}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{3}\end{array}$ NH -
+$C^{1}_{-1}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{3}\end{array}$ $e_{1}^{*}\wedge X^{*}+e_{2}^{*}\wedge e_{3}^{*}$ $\bar{1}$
+$C^{1}_{1/2}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3},\\[2.84526pt] [e_{3},e_{3}]&=&e_{2}\end{array}$ None -
+$C^{1}_{1/2}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3},\\[2.84526pt] [e_{3},e_{3}]&=&e_{2}\end{array}$ $e_{3}^{*}\wedge e_{3}^{*}+X^{*}\wedge X^{*}$ $\bar{0}$
+The LSA Relations in the basis: $e_{1}\,|\,e_{3},e_{4}$ and $X$ Symplectic structure $|\omega|$
+$C^{2}_{p}\oplus\mathbb{R}^{1|0},\;0<|p|\leq 1$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&pe_{4}\end{array}$ None -
+$C^{2}_{p}\oplus\mathbb{R}^{0|1},\;0<|p|\leq 1$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&pe_{4}\end{array}$ None -
+$C^{2}_{-1}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&-e_{4}\end{array}$ $e_{1}^{*}\wedge e_{4}^{*}+X^{*}\wedge e_{4}^{*}$ $\bar{1}$
+$C^{2}_{-1}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&-e_{4}\end{array}$ NH -
+$C^{3}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcr}[e_{1},e_{4}]&=&e_{3}\end{array}$ $e_{1}^{*}\wedge e_{3}^{*}+X^{*}\wedge e_{4}^{*}$ $\bar{1}$
+$C^{3}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcr}[e_{1},e_{4}]&=&e_{3}\end{array}$ NH -
+$C^{4}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]=e_{3}+e_{4}\end{array}$ None -
+$C^{4}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]=e_{3}+e_{4}\end{array}$ None -
+$C^{5}_{p}\oplus\mathbb{R}^{1|0},\;p\geq 0$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&pe_{3}-e_{4},&[e_{1},e_{4}]&=&e_{3}+pe_{4}\end{array}$ None -
+$C^{5}_{p}\oplus\mathbb{R}^{0|1},\;p\geq 0$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&pe_{3}-e_{4},&[e_{1},e_{4}]&=&e_{3}+pe_{4}\end{array}$ None -
+$(A_{1,1}+2A)^{1}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]=e_{1}\end{array}$ None -
+$(A_{1,1}+2A)^{1}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]=e_{1}\end{array}$ None -
+$(A_{1,1}+2A)^{2}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]=-e_{1}\end{array}$ None -
+$(A_{1,1}+2A)^{2}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]=-e_{1}\end{array}$ None -
+Table 13. Quasi-Frobenuis structure
 Acknowledgment.
-We would like to acknowledge Chengming Bai for pointing out that Novikov algebras have two superizations.
-Also, we would like to thank Saïd Benayadi, Isabel Cunha, and Quentin Ehret for reading the manuscript and making pertinent comments.
+We would like to acknowledge Chengming Bai for pointing out that Novikov algebras admit two superizations.
+Also, we would like to thank Saïd Benayadi, Isabel Cunha, Quentin Ehret, and the referee for reading the manuscript and making pertinent comments.
+Declarations:
+Funding Statement: SB’s research was supported by GRANT: AD-065.
+Competing Interest Statement: the authors declare no competing interest.
+Authors’ Contribution Statement: the authors contributed equally to this work.